### v1

# Dynamic Conditional Optimal Transport through Simulation-Free Flows

Abstract
We study the geometry of conditional optimal transport (COT) and prove a dynamical formulation which generalizes the Benamou-Brenier Theorem. With these tools, we propose a simulation-free flow-based method for conditional generative modeling. Our method couples an arbitrary source distribution to a specified target distribution through a triangular COT plan. We build on the framework of flow matching to train a conditional generative model by approximating the geodesic path of measures induced by this COT plan. Our theory and methods are applicable in the infinite-dimensional setting, making them well suited for inverse problems. Empirically, we demonstrate our proposed method on two image-to-image translation tasks and an infinite-dimensional Bayesian inverse problem.
1 Introduction
Many fundamental tasks in machine learning and statistics may be posed as estimating or sampling from a distribution, where one only has access to the distribution via a finite set of observations. Such tasks range from generative modeling and density estimation to Bayesian inference. A broad range of methods for sampling from an intractable distribution $p$ may be seen through the lens of measure transport. Intuitively, one fixes a reference distribution $q$ which is easy to sample from, and samples $z\sim q$ are transformed by a learned mapping $T$ such that $T(z)$ is approximately distributed as $p$. The distribution obtained in this way from $q$, denoted $T_{\#}q$, is the pushforward of $q$ along $T$.
Generative models (Ho et al., 2020; Song et al., 2020; Kingma and Welling, 2013; Goodfellow et al., 2014; Papamakarios et al., 2021) typically find such a mapping $T$ by minimizing a divergence between the target and generated distribution. In the standard maximum-likelihood based setting, the mapping $T$ is learned by minimizing (an upper bound on) the KL divergence between $p$ and the pushforward $T_{\#}q$. However, a natural alternative approach, based on optimal transport (Villani et al., 2009; Santambrogio, 2015), is to search for a map $T$ which satisfies the constraint $T_{\#}q=p$ and has minimal cost (in some sense).
In many scenarios, we are interested in conditional sampling. That is, we have access to some observation $y\in Y$, and from this information we would like to infer information regarding an unknown quantity $u\in U$. For instance, in Bayesian inference, $y$ denotes the data and $u$ the parameters of a model. A joint distribution $p(y,u)=p(y\mid u)p(u)$ is specified via a likelihood $p(y\mid u)$ and a prior $p(u)$, and one would like to draw samples from the posterior $p(u\mid y)$. However, in applications such as Bayesian inverse problems (Dashti and Stuart, 2013), one is interested in approximating $p(u\mid y)$ for many different observations $y$. Moreover, in applications such as simulation-based inference (Cranmer et al., 2020), the likelihood function may be intractable while we are still able to sample from $p(y,u)$. This motivates an amortized and likelihood-free (Amos et al., 2023) approach, where a single model may be trained and applied across different observations.
In the framework of measure transport, we first specify a source distribution $q(u)$. For any given, fixed $y\in Y$, we seek to find a mapping $T(y)$, depending on $y$, such that $T(y)_{\#}q(u)\approx p(u\mid y)$. Finding such a mapping for each individual $y$ via optimal transport, though, is typically intractable, as we frequently do not have samples from the posterior $p(u\mid y)$. As such, we must leverage information from other observations. One way to achieve this is through triangular mappings (Baptista et al., 2020), where a joint source distribution $q(y,u)$ is transformed by a mapping of the form $T:(y,u)\mapsto(T_{Y}(y),T_{U}(y,u))$. It can be shown (see Section 3.2) that if $T_{\#}q=p$, then $T_{U}(y,{-})$ almost surely couples $q(u\mid y)$ and $p(u\mid y)$. Thus, if we are able to find such a triangular $T$ using samples from the joint distribution $p(y,u)$, then we may obtain conditional samplers.
In this work, we develop a dynamical framework for conditional optimal transport (COT). In particular, we study absolutely curves of measures in a conditional Wasserstein space over $Y\times U$. Our main results (Theorem 14 and Theorem 15) show that such curves are, informally stated, generated by the flow induced by a triangular vector field. Intuitively, a triangular vector field on $Y\times U$ has no velocity in the $Y$ component and thus induces a flow for each conditional measure. As a consequence, in Theorem 16 we obtain a conditional version of the well-known Benamou-Brenier theorem (Benamou and Brenier, 2000). Importantly, our framework is applicable in infinite-dimensional spaces, making it well suited for function space inference problems.
We apply our results on dynamical COT to design conditional generative models, based on the flow matching framework (Lipman et al., 2022; Albergo et al., 2023a) . In particular, we generalize the methods of Pooladian et al. (2023a) and Tong et al. (2023), which utilize optimal transport to pair source and target samples, to the conditional setting. Our method works by solving a relaxed optimal transport problem (Hosseini et al., 2023) during training. In this way, we amortize solutions to the dynamic COT problem across arbitrary $y$. Our method may use arbitrary source distributions, and is applicable across a wide range of conditional generative modeling tasks.
Specifically, the main contributions of our work are as follows.
  1. 1.
    In Section 4, we study the conditional Wasserstein space $\mathbb{P}_{p}^{\mu}(Y\times U)$ equipped with the conditional Wasserstein distance $W_{p}^{\mu}(Y\times U)$. We show that this space is a metric space, and moreover, admits constant speed geodesics between any two points.
  2. 2.
    In Section 5, we characterize the absolutely continuous curves of measures in $\mathbb{P}_{p}^{\mu}(Y\times U)$ via the continuity equation and triangular vector fields. As a corollary, we obtain a conditional version of the Benamou-Brenier theorem (Benamou and Brenier, 2000).
  3. 3.
    In Section 6 we propose a method for learning dynamic COT maps based on flow matching. In Section 7, we demonstrate our method on two image-to-image translation tasks, as well as an infinite-dimensional Bayesian inverse problem.
2 Related Work
Optimal Transport.
Machine learning based approaches to unconditional optimal transport plans typically focus on the static problem (Makkuva et al., 2020; Korotin et al., 2022; Taghvaei and Jalali, 2019). Approaches to learning optimal transport maps in a dynamic fashion often require restricted architectures (Bunne et al., 2022b) or use likelihood-based approaches (Tong et al., 2020; Finlay et al., 2020; Onken et al., 2021; Huguet et al., 2022), which require simulating from the model during training. While most work focuses on the squared norm cost, the recent approaches of Pooladian et al. (2023b) and Neklyudov et al. (2023) study dynamical optimal transport with costs specified by a Lagrangian. These frameworks, however, typically have no clear interpretation as a static problem.
Conditional optimal transport (COT), on the other hand, is relatively unexplored. Bunne et al. (2022a) and Wang et al. (2023) learn supervised COT maps in a static setting via convex networks (Makkuva et al., 2020). Hosseini et al. (2023) recently develop a framework for conditional optimal transport that is applicable in infinite-dimensional spaces. This builds on the work of Carlier et al. (2016) which applies COT for vector quantile regression. A common approach to COT is to employ triangular plans (Baptista et al., 2020). Our work develops a dynamic formulation of the static COT problem (Hosseini et al., 2023; Carlier et al., 2016) using a notion of triangular vector fields. Additionally, Wang et al. (2023) propose a likelihood-based approach using continuous normalizing flows, but do not develop the theory of dynamical COT as we do here.
Simulation-Free Continuous Normalizing Flows.
Flow matching (Lipman et al., 2022) and the equivalent stochastic interpolant framework (Albergo and Vanden-Eijnden, 2022; Albergo et al., 2023a) are a class of methods for building continuous-time normalizing flows in a simulation-free manner. These methods transform an arbitrary source distribution into a target distribution via a specified path of measures, which is modeled through a learned vector field. These methods have been extended to setting of Riemannian manifolds (Chen and Lipman, 2023) and infinite-dimensional spaces (Kerrigan et al., 2023).
Notably, these works do not approximate an optimal transport between the source and target measures and instead use an independent coupling, although stochastic interpolants (Albergo and Vanden-Eijnden, 2022; Albergo et al., 2023b) may be rectified (Liu et al., 2022; Lee et al., 2023) through an iterative optimization procedure. Importantly, the paths in flow matching (Lipman et al., 2022) maybe taken to be conditionally optimal, but marginally they are not. Pooladian et al. (2023a) and Tong et al. (2023) propose instead to couple the source and target distributions via optimal transport.
While some works (Lipman et al., 2022; Davtyan et al., 2023; Gebhard et al., 2023) have studied flow matching approaches for conditional distributions, these approaches simply provide additional conditioning information to the model and do not study COT. Concurrent work (Isobe et al., 2024) proposes extended flow matching, which learns a matrix field to perform conditional generation. However, the results of Isobe et al. (2024) are limited to the finite-dimensional setting and do not develop the theory of COT.
At the time of our submission, concurrent work by Barboni et al. (2024) and Chemseddine et al. (2024) appeared with similar results. Our results were developed independently from these works.
3 Background and Notation
Let $X,X^{\prime},X^{\prime\prime}$ represent arbitrary separable Hilbert spaces, equipped with the Borel $\sigma$-algebra. We use $\mathbb{P}(X)$ to represent the space of Borel probability measures on $X$, and $\mathbb{P}_{p}(X)\subseteq\mathbb{P}(X)$ to represent the subspace of measures having finite $p$th moment. If $\eta\in\mathbb{P}(X)$ is a probability measure on $X$ and $T:X\to X^{\prime}$ is measurable, then the pushforward measure $T_{\#}\eta({-})=\eta(T^{-1}({-}))$ is a probability measure on $X^{\prime}$. Maps of the form $\pi^{X}:X\times X^{\prime}\to X$ represent the canonical projection onto the respective space. We will also use maps such as $\pi^{1}:X\times X^{\prime}\times X^{\prime\prime}\to X$ and $\pi^{1,2}:X\times X^{\prime}\times X^{\prime\prime}\to X\times X^{\prime}$ to denote projection onto specific coordinates. The set $C_{b}(X)$ represents the space of continuous and bounded functions and $C^{\infty}_{c}(X)$ is the space of smooth and compactly supported functions.
In our specific case, we assume that we have two separable Hilbert spaces of interest. The first, $Y$, is a space of observations, and the second, $U$, is a space of unknowns. These spaces may be of infinite dimensions, but a special case of practical interest is when $Y$ and $U$ are finite dimensional Euclidean spaces. We will consider the product space $Y\times U$, equipped with the canonical inner product obtained via the sum of the inner products on $Y$ and $U$, under which the space $Y\times U$ is also a separable Hilbert space.
Let $\eta\in\mathbb{P}(Y\times U)$ be a joint measure. The measures $\pi^{Y}_{\#}\eta\in\mathbb{P}(Y)$ and $\pi^{U}_{\#}\eta\in\mathbb{P}(U)$ obtained via projection are the marginals of $\eta$. We use $\eta^{y}\in\mathbb{P}(U)$ to represent the measure obtained by conditioning $\eta$ on the value $y\in Y$. By the disintegration theorem, such conditional measures exist and are essentially unique, in the sense that there exists a Borel set $E\subseteq Y$ with $\pi^{Y}_{\#}\eta(E)=0$, and $\eta^{y}$ are unique for $y\notin E$.
3.1 Unconditional Optimal Transport
We provide here a brief overview of optimal transport in the standard unconditional setting. For more details, we refer to the standard references of Villani et al. (2009); Santambrogio (2015); Ambrosio et al. (2005). Fix a cost function $c:X\times X\to\mathbb{R}\cup\{+\infty\}$. Suppose we have two Borel measures $\eta,\nu\in\mathbb{P}(X)$. The Monge Problem seeks to find a measurable transport map $T:X\to X$ minimizing the expected cost of transport, i.e. corresponding to the optimization problem
$\inf_{T}\left\{\int_{X}c(x,T(x))\,\mathrm{d}\eta(x)\mid T_{\#}\eta=\nu\right\}.$ (1)
This optimization problem is challenging, though, as it involves a nonlinear constraint and the set of feasible maps may be empty. In contrast, the Kantorovich problem is a relaxation which seeks to find an optimal coupling $\gamma\in\Pi(\eta,\nu)$, i.e. a probability distribution over $X\times X$ with marginals $\eta,\nu$, which solves
$\inf_{\gamma}\left\{\int_{X\times X}c(x_{0},x_{1})\,\mathrm{d}\gamma(x_{0},x_{1})\mid\gamma\in\Pi(\eta,\nu)\right\}.$ (2)
Under fairly weak conditions (e.g., the cost is lower semicontinuous and bounded from below (Ambrosio et al., 2013, Theorem 2.5)), minimizers to the Kantorovich problem are guaranteed to exist. If the cost function is $c(x_{0},x_{1})=|x_{0}-x_{1}|^{p}$ for some $1<p<\infty$, under sufficient regularity conditions on $\eta$ a solution $T^{\star}$ to the Monge problem is guaranteed to exist and, moreover, the coupling $\gamma^{\star}=(I,T^{\star})_{\#}\eta$ is optimal for the Kantorovich problem. See Ambrosio et al. (2013, Chapter 2) and Ambrosio et al. (2005, Theorem 6.2.10).
Wasserstein Space.
In the special case that $c(x_{0},x_{1})=|x_{0}-x_{1}|^{p}$ for $1\leq p<\infty$, and $\eta,\nu\in\mathbb{P}_{p}(X)$, the Kantorovich problem admits a finite-cost solution. The cost of such an optimal coupling is the $p$-Wasserstein distance
$W_{p}^{p}(\eta,\nu)=\min_{\gamma}\left\{\int_{X\times X}|x_{0}-x_{1}|^{p}\,\mathrm{d}\gamma(x_{0},x_{1})\mid\gamma\in\Pi(\eta,\nu)\right\}$ (3)
which, as the name suggests, is a metric on the space $\mathbb{P}_{p}(X)$ (Ambrosio et al., 2005, Section 7.1) (Santambrogio, 2015, Section 5.1). The Wasserstein distance admits a dynamical formulation via the Benamou-Brenier theorem (Benamou and Brenier, 2000). Namely, the $p$-Wasserstein distance can be obtained by finding a time-dependent vector field transforming $\eta$ to $\nu$ across time $t\in[0,1]$ with minimal energy:
$W_{p}(\eta,\nu)=\min_{(\gamma_{t},v_{t})}\left\{\int_{0}^{1}\int_{X}|v_{t}(x)|^{p}\,\mathrm{d}\gamma_{t}(x)\,\mathrm{d}t\mid\gamma_{0}=\eta,\gamma_{1}=\nu,\partial_{t}\gamma_{t}+\text{div}(v_{t}\gamma_{t})=0\right\}.$ (4)
Here, we constrain our minimization problem over the set of measures and vector fields $(\gamma_{t},v_{t})$ interpolating between $\eta$ and $\nu$, satisfying a continuity equation (see Section 5.1). In Section 4, we study a generalization of the Wasserstein distances for conditional optimal transport problems. In particular, Theorem 16 provides a generalization of the Benamou-Brenier theorem to the conditional setting which recovers a conditional Wasserstein distance.
3.2 Static Conditional Optimal Transport
In the unconditional setting, one seeks an optimal coupling between two given measures. In contrast, in the conditional setting, we would like to find a family of couplings that can transport a given source measure into any conditional measure of the target distribution. In other words, given a target measure $\nu\in\mathbb{P}(Y\times U)$ and some source measure $\eta\in\mathbb{P}(U)$, one seeks a transport map $T:Y\times U\to U$ such that
$T_{\#}(y,{-})_{\#}\eta=\nu^{y}\qquad\forall y\in Y.$ (5)
If such a map were available, by drawing samples $u_{0}\sim\eta$ and transforming them, one would obtain samples $T(y,u)\sim\nu^{y}$. Solving this transport problem for each fixed $y$ is expensive at best, or impossible when only has a single (or no) samples $(y,u)\sim\nu$ for any given $y$.
Thus, one must leverage information across different observations $y$. To that end, recent work has focused on the notion of triangular mappings $T:Y\times U\to Y\times U$ (Hosseini et al., 2023; Baptista et al., 2020) (Bogachev and Ruas, 2007, Section 10.10) of the form
$T(y,u)=\left(T_{Y}(y),T_{U}(T_{Y}(y),u)\right)$ (6)
for some $T_{Y}:Y\to Y$ and $T_{U}:Y\times U\to U$. Triangular mappings are of interest as they allow us to obtain conditional couplings from joint couplings.
Proposition 1 (Theorem 2.4 (Baptista et al., 2020), Prop. 2.3 (Hosseini et al., 2023)).
Suppose $\eta,\nu\in\mathbb{P}(Y\times U)$ and $T:Y\times U\to Y\times U$ is triangular. If $T_{\#}\eta=\nu$, then $T_{U}(T_{Y}(y),{-})_{\#}\eta^{y}=\nu^{T_{Y}(y)}$ for $\pi^{Y}_{\#}\eta$-almost every $y$.
In many scenarios of practical interest, the source measure $\eta$ and the target measure $\nu$ have the same $Y$-marginals. We will henceforth make this assumption, and use $\mu=\pi^{Y}_{\#}\eta=\pi^{Y}_{\#}\nu$ to represent this marginal. In this case, we may take $T_{Y}$ to be the identity mapping, so that the conclusion of Proposition 1 simplifies to $T_{U}(y,{-})_{\#}\eta^{y}=\nu^{y}$ for $\mu$-almost every $y$. We note that in situations where such an assumption does not hold, one may simply preprocess the source measure $\eta$ via an invertible mapping $T_{Y}$ satisfying $[T_{Y}]_{\#}[\pi^{Y}_{\#}\eta]=\pi^{Y}_{\#}\nu$ (Hosseini et al., 2023, Prop 3.2).
Definition 2 (Conditional Wasserstein Space).
Suppose $\mu\in\mathbb{P}(Y)$ is a given probability measure over $Y$ and $1\leq p<\infty$. The set
$\mathbb{P}^{\mu}(Y\times U)=\left\{\gamma\in\mathbb{P}_{p}(Y\times U)\mid\pi^{Y}_{\#}\gamma=\mu\right\}$ (7)
is the space of joint measures on $Y\times U$ having fixed $Y$-marginals $\mu$. The set $\mathbb{P}_{p}^{\mu}(Y\times U):=\mathbb{P}^{\mu}(Y\times U)\cap\mathbb{P}_{p}(Y\times U)$ is the conditional $p$-Wasserstein space, consisting of joint measures with finite $p$th moments and $Y$-marginals $\mu$.
Given a source and target measures $\eta,\nu\in\mathbb{P}^{\mu}(Y\times U)$, the conditional Monge problem seeks to find a triangular mapping solving
$\inf_{T}\left\{\int_{Y\times U}c(y,u,T(y,u))\,\mathrm{d}\eta(y,u)\mid T_{\#}\eta=\nu,T:(y,u)\mapsto(y,T_{U}(y,u))\right\}.$ (8)
The conditional Monge problem also admits a relaxation under which one only considers couplings whose $Y$-components are almost surely equal. To that end, we consider the subset $\mathscr{C}\subset(Y\times U)^{2}$ whose $Y$ components are identical, i.e.,
$\mathscr{C}:=\left\{(y_{0},u_{0},y_{1},u_{1})\in(Y\times U)^{2}\mid y_{0}=y_{1}\right\}$ (9)
and we define the set of $(Y)$-restricted probability measures $\mathcal{R}_{Y}\subset\mathbb{P}\left((Y\times U)^{2}\right)$ such that every $\gamma\in\mathcal{R}_{Y}$ is concentrated on $\mathscr{C}$. In other words, if $\gamma\in\mathcal{R}_{Y}$, then samples $(y_{0},u_{0},y_{1},u_{1})\sim\gamma$ have $y_{0}=y_{1}$ almost surely. In addition, for any $\eta,\nu\in\mathbb{P}(Y\times U)$, we define the set of $(Y)$-restricted couplings $\Pi_{Y}(\eta,\nu)$ to be the probability measures in $\mathcal{R}_{Y}$ whose marginals are $\eta$ and $\nu$, i.e.
$\Pi_{Y}(\eta,\nu)=\left\{\gamma\in\mathcal{R}_{Y}\mid\pi^{1,2}_{\#}\gamma=\eta,\pi^{3,4}_{\#}\gamma=\nu\right\}.$ (10)
The conditional Kantorovich problem seeks a triangular coupling $\gamma^{\star}$ solving
$\inf_{\gamma}\left\{\int_{(Y\times U)^{2}}c(y_{0},u_{0},y_{1},u_{1})\,\mathrm{d}\gamma(y_{0},u_{0},y_{1},u_{1})\mid\gamma\in\Pi_{Y}(\eta,\nu)\right\}.$ (11)
Hosseini et al. (2023) prove the existence of minimizers to the conditional Kantorovich and Monge problems under very general assumptions. Moreover, optimal couplings to the conditional Kantorovich problem induce optimal couplings for $\mu$-almost every conditional measure. Assuming sufficient regularity assumptions on the conditional measures, unique solutions to the conditional Monge problem exist.
Proposition 3 (Prop 3.3 (Hosseini et al., 2023)).
Fix $\eta,\nu\in\mathbb{P}^{\mu}(Y\times U)$. Suppose the cost function $c$ is continuous, $\inf c>-\infty$, and there exists a finite cost coupling $\gamma\in\Pi_{Y}(\eta,\nu)$. Then, the conditional Kantorovich problem admits a minimizer $\gamma^{\star}$. Moreover, $\gamma^{\star,y_{0}}(y_{1},u_{0},u_{1})=\hat{\gamma}^{\star,y_{0}}(u_{0},u_{1})\delta(y_{1}-y_{0})$ where for $\mu$-almost every $y$ the measure $\gamma^{\star,y}$ is an optimal coupling for $\eta^{y},\nu^{y}$ under the cost $c^{y}(u_{0},u_{1})=c(y,u_{0},y,u_{1})$
Proposition 4 (Prop 3.8 (Hosseini et al., 2023)).
Fix $1<p<\infty$ and $\eta,\nu\in\mathbb{P}_{p}^{\mu}(Y\times U)$. Suppose $c(y_{0},u_{0},y_{1},u_{1})=|u_{0}-u_{1}|^{p}$. If $\eta^{y}$ assign zero measure to Gaussian null sets for $\mu$-almost every $y$, then there is a unique solution $T^{\star}$ to the conditional Monge problem, and $\gamma^{\star}=(I,T^{\star})_{\#}\eta$ is the unique solution to the conditional Kantorovich problem. If $\nu^{y}$ also assign zero measure to Gaussian null sets for $\mu$-almost every $y$, then $T^{\star}$ is injective $\eta$-almost everywhere.
4 Conditional Wasserstein Space
We now introduce a metric on the space $\mathbb{P}_{p}^{\mu}(Y\times U)$. Intuitively, the conditional Wasserstein distance measures the usual Wasserstein distance between all of the conditional distributions in expectation under the fixed $Y$-marginal $\mu$.
Definition 5 (Conditional $p$-Wasserstein Distance).
Suppose $\eta,\nu\in\mathbb{P}_{p}^{\mu}(Y\times U)$ and $1\leq p<\infty$. The function
$W_{p}^{\mu}(\eta,\nu)=\left(\mathbb{E}_{y\sim\mu}\left[W_{p}^{p}(\eta^{y},\nu^{y})\right]\right)^{1/p}=\left(\int_{Y}W_{p}^{p}(\eta^{y},\nu^{y})\,\mathrm{d}\mu(y)\right)^{1/p}$ (12)
is the conditional $p$-Wasserstein distance. Here, $W_{p}^{p}:\mathbb{P}_{p}(U)\times\mathbb{P}_{p}(U)\to\mathbb{R}$ is the usual $p$-Wasserstein distance for measures on $U$.
By an application of Jensen’s inequality we observe that $W_{p}^{\mu}(\eta,\nu)\geq\mathbb{E}_{y\sim\mu}\left[W_{p}(\eta^{y},\nu^{y})\right]$. For $p>1$, this inequality is strict unless $W_{p}(\eta^{y},\nu^{y})$ is $\mu$-almost surely constant. We emphasize that Definition 5 is purely formal at this stage. We do not know if this integral is well-defined, or if it actually defines a metric. Before proceeding, we first note that $W_{p}^{\mu}(\eta,\nu)$ may be viewed as the minimal value of the constrained Kantorovich problem in Equation (11) when one takes the cost to be the metric on the space $Y\times U$. Although Hosseini et al. (2023) studied the existence of minimizers to this problem, they did not study the metric properties of the resulting space.
Proposition 6 (Equivalent Formulation of the $p$-Wasserstein Distance).
Fix $\eta,\nu\in\mathbb{P}_{p}^{\mu}(Y\times U)$ and $1\leq p<\infty$. Then, $W_{p}^{\mu}(\eta,\nu)$ is well-defined, finite, and
$W_{p}^{\mu,p}(\eta,\nu)=\min_{\gamma}\left\{\int_{(Y\times U)^{2}}d^{p}(y_{0},u_{0},y_{1},u_{1})\,\mathrm{d}\gamma\mid\gamma\in\Pi_{Y}(\eta,\nu)\right\}$ (13)
where $W_{p}^{\mu,p}(\eta,\nu)$ represents the $p$-th power of the conditional $p$-Wasserstein distance.
Proof.
The cost function $d^{p}$ is clearly continuous and non-negative, and hence by Proposition 3 it suffices to exhibit a finite-cost coupling $\gamma\in\Pi_{Y}(\eta,\nu)$ between $\eta$ and $\nu$. Indeed, take the conditionally independent coupling
$\gamma(y_{0},u_{0},y_{1},u_{1})=\eta(u_{0}\mid y_{1})\nu(u_{1}\mid y_{1})\delta(y_{1}-y_{0})\mu(y_{1})$ (14)
which is clearly in $\Pi_{Y}(\eta,\nu)$. We then have that
$\int_{(Y\times U)^{2}}d^{p}(y_{0},u_{0},y_{1},u_{1})\,\mathrm{d}\gamma(y_{0},u_{0},y_{1},u_{1})=\int_{(Y\times U)^{2}}\left\lVert(y_{0},u_{0})-(y_{1},u_{1})\right\rVert_{Y\times U}^{p}\,\mathrm{d}\gamma(y_{0},u_{0},y_{1},u_{1})$
$\leq 2^{p}\int_{(Y\times U)^{2}}\left(\left\lVert(y_{0},u_{0})\right\rVert_{Y\times U}^{p}+\left\lVert(y_{1},u_{1})\right\rVert_{Y\times U}^{p}\right)\,\mathrm{d}\gamma(y_{0},u_{0},y_{1},u_{1})$
$=2^{p}\left(\int_{Y\times U}\left\lVert(y_{0},u_{0})\right\rVert_{Y\times U}^{p}\,\mathrm{d}\eta(y_{0},u_{0})+\int_{Y\times U}\left\lVert(y_{1},u_{1})\right\rVert_{Y\times U}^{p}\,\mathrm{d}\nu(y_{1},u_{1})\right)<+\infty.$
Hence, Equation (13) admits a minimizer $\gamma^{\star}\in\Pi_{Y}(\eta,\nu)$. By Proposition 3, this minimizer may be taken to have the form $\gamma^{\star}=\gamma^{\star,{y_{1}}}(u_{0},u_{1})\delta(y_{1}-y_{0})\mu(y_{1})$ where $\gamma^{\star,{y_{1}}}(u_{0},u_{1})$ is $\mu(y_{1})$-almost surely an optimal coupling between $\eta^{y_{1}},\nu^{{y_{1}}}$ for the cost $|u_{1}-u_{0}|^{p}$. Thus,
$\int_{(Y\times U)^{2}}d^{p}\,\mathrm{d}\gamma^{\star}$ $=\int_{Y}\int_{U^{2}}|u_{1}-u_{0}|^{p}\,\mathrm{d}\gamma^{\star,y}(u_{0},u_{1})\,\mathrm{d}\mu(y)$ (15)
$=\int_{Y}W_{p}^{p}(\eta^{y},\nu^{y})\,\mathrm{d}\mu(y)=W_{p}^{p,\mu}(\eta,\nu).$ (16)
Here, we emphasize that the $\mu$-almost sure uniqueness of the disintegrations of $\eta,\nu$ along $Y$ result in a well-defined expression.
Moreover, if $\eta\in\mathbb{P}_{p}^{\mu}(Y\times U)$ it follows that $\eta^{y}\in\mathbb{P}_{p}(U)$ for $\mu$-a.e. $y$, because
$\int_{Y}\int_{U}|u|^{p}\,\mathrm{d}\eta^{y}(u)\,\mathrm{d}\mu(y)$ $\leq\int_{Y}\int_{U}|(y,u)|^{p}\,\mathrm{d}\eta^{y}(u)\,\mathrm{d}\mu(y)$ (17)
$=\int_{Y\times U}|(y,u)|^{p}\,\mathrm{d}\eta(y,u)<+\infty.$ (18)
Thus all considered $p$-Wasserstein distances on $U$ are finite. ∎
We now show that $W_{p}^{\mu}$ is a metric on $\mathbb{P}_{p}^{\mu}(Y\times U)$.
Proposition 7 ($W_{p}^{\mu}$ is a Metric).
For $1\leq p<\infty$, the conditional $p$-Wasserstein distance $W_{p}^{\mu}$ is a metric on $\mathbb{P}_{p}^{\mu}(Y\times U)$.
Proof.
Fix $\eta,\nu,\rho\in\mathbb{P}_{p}^{\mu}(Y\times U)$. Since $W_{p}$ is a metric on $\mathbb{P}_{p}(U)$, we immediately obtain the symmetry of $W_{p}^{\mu}$. Moreover, we have that $W_{p}^{\mu}(\eta,\nu)=0$ if and only if $\eta^{y}=\nu^{y}$ for $\mu$-almost every $y$. Thus, if $W_{p}^{\mu}(\eta,\nu)=0$ and $E\subseteq Y\times U$ is Borel measurable,
$\eta(E)=\int_{Y}\eta^{y}(E^{y})\,\mathrm{d}\mu(y)=\int_{Y}\nu^{y}(E^{y})\,\mathrm{d}\mu(y)=\nu(E).$ (19)
which shows that $\eta=\nu$. Here, $E^{y}=\{u\mid(y,u)\in E\}$ is the $y$-slice of $E$. Conversely, if $\eta=\nu$, then $\eta^{y}=\nu^{y}$ up to a $\mu$-null set by the essential uniqueness of disintegrations. Thus, $W_{p}^{\mu}(\eta,\nu)=0$ if and only if $\eta=\nu$.
By Minkowski’s inequality and the triangle inequality for $W_{p}$ on $\mathbb{P}_{p}(U)$, we see
$W_{p}^{\mu}(\eta,\nu)$ $\leq\left(\mathbb{E}_{y\sim\mu}\left[(W_{p}(\eta^{y},\rho^{y})+W_{p}(\rho^{y},\nu^{y}))^{p}\right]\right)^{1/p}$ (20)
$\leq\mathbb{E}_{y\sim\mu}[W_{p}^{p}(\eta^{y},\rho^{y})]^{1/p}+\mathbb{E}_{y\sim\mu}[W_{p}^{p}(\rho^{y},\nu^{y})]^{1/p}$ (21)
$=W_{p}^{\mu}(\eta,\rho)+W_{p}^{\mu}(\rho,\nu).$ (22)
4.1 Metric Properties of Conditional Wasserstein Distances
We now briefly explore some properties of the space $\mathbb{P}_{p}^{\mu}(Y\times U)$ equipped with the conditional Wasserstein distance $W_{p}^{\mu}$. In particular, we observe that the topology on $\mathbb{P}_{p}^{\mu}(Y\times U)$ generated by $W_{p}^{\mu}$ is stronger than that generated by $W_{p}$. Note that we easily see $W_{p}^{\mu}(\eta,\nu)\geq W_{p}(\eta,\nu)$ for every $\eta,\nu\in\mathbb{P}_{p}^{\mu}(Y\times U)$ by the characterization in Proposition 6. The converse is, however, false, and the metrics are inequivalent.
Proposition 8 ($W_{p}^{\mu}$ is stronger than $W_{p}$).
There does not in general exist a constant $C>0$ such that $W_{p}^{\mu}(\eta,\nu)\leq CW_{p}(\eta,\nu)$ for all $\eta,\nu\in\mathbb{P}_{p}^{\mu}(Y\times U)$.
Proof.
We provide a counterexample. Fix any $u_{0}\neq 0\in U$ and $y_{0},y_{1}\in Y$ such that $y_{0}\neq y_{1}$. Define $\mu=\frac{1}{2}\left(\delta_{y_{0}}+\delta_{y_{1}}\right)$. Set $u_{k}=(k+1)u_{0}$ for $k=1,2,\dots$ and for each $k$, define two measures on $Y\times U$ by
$\eta_{k}=\frac{1}{2}\left(\delta_{y_{0}u_{0}}+\delta_{y_{1}u_{k}}\right)\qquad\nu_{k}=\frac{1}{2}\left(\delta_{y_{1}u_{0}}+\delta_{u_{k}y_{0}}\right).$ (23)
It is clear that
$W_{p}^{\mu,p}(\eta_{k},\nu_{k})=k^{p}|u_{0}|^{p}\qquad W_{p}^{p}(\eta_{k},\nu_{k})=\min\{k^{p}|u_{0}|^{p},|y_{1}-y_{0}|^{p}\}.$ (24)
Moreover, as $k\to\infty$ we have $W_{p}^{\mu}(\mu_{k},\nu_{k})\to\infty$ but $W_{p}^{p}(\nu_{k},\eta_{k})$ remains bounded. ∎
In addition, we can show that the conditional Wasserstein distance dominates the Wasserstein distance on the $U$ marginals.
Proposition 9 (Conditional Wasserstein Dominates $U$-Marginal Distance).
For $1\leq p<\infty$ and $\eta,\nu\in\mathbb{P}_{p}^{\mu}(Y\times U)$, we have
$W_{p}\left(\pi^{U}_{\#}\eta,\pi^{U}_{\#}\nu\right)\leq W_{p}^{\mu}(\eta,\nu).$ (25)
Proof.
Let $\gamma^{\star}(y_{0},u_{0},y_{1},u_{1})=\gamma^{\star,y_{1}}(u_{0},u_{1})\delta(y_{1}-y_{0})\mu(y_{1})$ be an optimal $\gamma^{\star}\in\Pi_{Y}(\eta,\nu)$. We claim that $\gamma(u_{0},u_{1}):=\int_{Y}\gamma^{\star,y}(u_{0},u_{1})\,\mathrm{d}\mu(y)$ couples $\pi^{U}_{\#}\eta$ and $\pi^{U}_{\#}\nu$. Let $\pi^{0}:(u_{0},u_{1})\mapsto u_{0}$ be the projection onto the first coordinate of $U\times U$. Observe that for $\mu$-almost every $y$, we have that $\gamma^{\star,y}\in\Pi(\eta^{y},\nu^{y})$ is optimal, and, in particular, $\pi^{0}_{\#}\gamma^{\star,y}=\eta^{y}.$ Fix an arbitrary $\varphi\in C_{b}(U)$. We then have
$\int_{U}\varphi(u_{0})\,\mathrm{d}\pi^{0}_{\#}\gamma(u_{0})=\int_{U^{2}}(\varphi\circ\pi^{0})\,\mathrm{d}\gamma(u_{0},u_{1})$ (26)
$=\int_{Y}\int_{U^{2}}(\varphi\circ\pi^{0})\,\mathrm{d}\gamma^{\star,y}(u_{0},u_{1})\,\mathrm{d}\mu(y)=\int_{Y}\int_{U}\varphi(u_{0})\,\mathrm{d}\pi^{0}_{\#}\gamma^{\star,y}(u_{0})\,\mathrm{d}\mu(y)$ (27)
$=\int_{Y}\int_{U}\varphi(u_{0})\,\mathrm{d}\eta^{y}(u_{0})\,\mathrm{d}\mu(y){=}\int_{Y\times U}\varphi(u_{0})\,\mathrm{d}\eta(u_{0},y)$ (28)
$=\int_{Y\times U}(\varphi\circ\pi^{U})\,\mathrm{d}\eta(u_{0},y)=\int_{U}\varphi\,\mathrm{d}\pi^{U}_{\#}\eta(u_{0}).$ (29)
Thus $\pi^{U}_{\#}\gamma=\pi^{U}_{\#}\eta$. A similar argument shows that for the map $\pi^{1}:(u_{0},u_{1})\mapsto u_{1}$ we have $\pi^{1}_{\#}\gamma=\pi^{U}_{\#}\nu$, so that $\gamma\in\Pi(\pi^{U}_{\#}\eta,\pi^{U}_{\#}\nu)$.
Now, as $\gamma^{\star,y_{1}}(u_{0},u_{1})\in\Pi(\eta^{y_{1}},\nu^{y_{1}})$ is $\mu$-almost surely optimal in the usual Wasserstein sense,
$W_{p}^{p,\mu}(\eta,\nu)$ $=\int_{Y}\int_{U^{2}}|u_{0}-u_{1}|^{p}\,\mathrm{d}\gamma^{\star,y}(u_{0},u_{1})\,\mathrm{d}\mu(y)$ (30)
$=\int_{U^{2}}|u_{0}-u_{1}|^{p}\,\mathrm{d}\gamma(u_{0},u_{1})$ (31)
$\geq W_{p}^{p}(\pi^{U}_{\#}\eta,\pi^{U}_{\#}\nu)$ (32)
since $\gamma\in\Pi(\pi^{U}_{\#}\eta,\pi^{U}_{\#}\nu)$ is a coupling but potentially sub-optimal. ∎
Example: Gaussian Measures.
We conclude with an example where the conditional $2$-Wasserstein distance may be explicitly computed. In particular, this example shows that the inequality in Proposition 9 may be strict. Suppose $Y$ and $U$ are Euclidean spaces (of possibly different dimensions), and that $\eta,\nu\in\mathbb{P}_{p}^{\mu}(Y\times U)$ are Gaussians of the form
$\eta=\mathcal{N}\left(\begin{bmatrix}m\\ m_{u}^{\eta}\end{bmatrix},\begin{bmatrix}\Sigma&\Sigma^{\eta}_{12}\\ \Sigma^{\eta}_{21}&\Sigma^{\eta}_{22}\end{bmatrix}\right)\qquad\nu=\mathcal{N}\left(\begin{bmatrix}m\\ m_{u}^{\nu}\end{bmatrix},\begin{bmatrix}\Sigma&\Sigma^{\nu}_{12}\\ \Sigma^{\nu}_{21}&\Sigma^{\nu}_{22}\end{bmatrix}\right).$ (33)
It follows that $\mu=\pi^{Y}_{\#}\eta=\pi^{Y}_{\#}\nu=\mathcal{N}(m,\Sigma)$. Because we may obtain $\eta^{y},\nu^{y}$ in closed form, we can compute $W_{p}^{\mu}(\eta,\nu)$ in closed form. Let
$Q^{\eta}=\Sigma_{22}^{\eta}-\Sigma_{21}^{\eta}\Sigma^{-1}\Sigma_{12}^{\eta}\qquad Q^{\nu}=\Sigma_{22}^{\nu}-\Sigma_{21}^{\nu}\Sigma^{-1}\Sigma_{12}^{\nu}\qquad R=(\Sigma_{21}^{\eta}-\Sigma_{21}^{\nu})\Sigma^{-1}$ (34)
Then, we obtain by a brute-force calculation that
$W_{2}^{\mu,2}(\eta,\nu)=|m_{u}^{\eta}-m_{u}^{\nu}|^{2}+\text{Tr}\left(Q^{\eta}+Q^{\nu}-2\left((Q^{\eta})^{1/2}Q^{\nu}(Q^{\eta})^{1/2}\right)^{1/2}+R\Sigma R^{T}\right).$ (35)
Note that when $\eta,\nu$ have uncorrelated $Y,U$ components, we precisely recover $W_{2}^{2}(\pi^{U}_{\#}\eta,\pi^{U}_{\#}\nu)$ as one may expect. As a special case of interest, if $Y=U=\mathbb{R}$ and
$\eta=\mathcal{N}(0,I)\qquad\nu=\mathcal{N}\left(0,\begin{bmatrix}1&\rho\\ \rho&1\end{bmatrix}\right)\quad|\rho|<1$ (36)
then we obtain as a special case of Equation (35) that $W_{2}^{\mu,2}(\eta,\nu)=2(1-\sqrt{1-\rho^{2}})$. This is zero if and only if $\rho=0$, i.e. $\eta=\nu$. However, $\pi^{U}_{\#}\eta=\pi^{U}_{\#}\nu=\mathcal{N}(0,1)$.
4.2 Conditional Wasserstein Space as a Geodesic Space
In this section, we show that $\mathbb{P}_{p}^{\mu}(Y\times U)$ is a geodesic space. In fact, we show the stronger statement that there exists a constant speed geodesic between any two measures in $\mathbb{P}_{p}^{\mu}(Y\times U)$, which is a generalization of known results (Santambrogio, 2015, Theorem 5.27). First, we introduce some preliminary notions. A curve is a continuous function $\gamma_{\bullet}:I\to\mathbb{P}_{p}^{\mu}(Y\times U)$ where $I=(a,b)\subseteq\mathbb{R}$ is any open interval of finite length. A curve is said to be absolutely continuous if there exists $m\in L^{1}((a,b))$ such that
$W_{p}^{\mu}(\gamma_{s},\gamma_{t})\leq\int_{s}^{t}m(\tau)\,\mathrm{d}\tau\qquad\forall a<s\leq t<b.$ (37)
If $(\gamma_{t})$ is an absolutely continuous curve, then its metric derivative
$|\gamma^{\prime}|(t)=\lim_{s\to t}\frac{W_{p}^{\mu}(\gamma_{s},\gamma_{t})}{|s-t|}$ (38)
exists for almost every $t\in(a,b)$, and, moreover, we almost surely have $|\gamma^{\prime}|(t)\leq m(t)$ pointwise for any $m$ satisfying Equation (37) (Ambrosio et al., 2005, Theorem 1.1.2). A curve $(\gamma_{t})$ is called a constant speed geodesic if for all $a<s\leq t<b$, we have
$W_{p}^{\mu}(\gamma_{s},\gamma_{t})=|t-s|W_{p}^{\mu}(\gamma_{a},\gamma_{b}).$ (39)
It is straightforward to show that every constant speed geodesic is absolutely continuous.
Theorem 10 ($\mathbb{P}_{p}^{\mu}(Y\times U$) is a Geodesic Space).
$\mathbb{P}_{p}^{\mu}(Y\times U)$ admits constant-speed geodesics. That is, for any $\eta,\nu\in\mathbb{P}_{p}^{\mu}(Y\times U)$, there exists a constant speed geodesic, contained within $\mathbb{P}_{p}^{\mu}(Y\times U)$, between $\eta$ and $\nu$.
Proof.
Write $\lambda_{t}:(Y\times U)^{2}\to Y\times U$ for the linear interpolant
$\lambda_{t}(y_{0},u_{0},y_{1},u_{1})=(ty_{0}+(1-t)y_{1},tu_{0}+(1-t)u_{1})\qquad 0\leq t\leq 1.$ (40)
Let $\gamma^{\star}\in\Pi_{Y}(\eta,\nu)$ be an optimal restricted coupling, and consider the path of measures in $\mathbb{P}_{p}(Y\times U)$ given by
$\gamma_{t}=[\lambda_{t}]_{\#}\gamma^{\star}\qquad 0\leq t\leq 1.$ (41)
Step one: We check that for each $0\leq t\leq 1$, we have $\gamma_{t}\in\mathbb{P}_{p}^{\mu}(Y\times U)$. That is, we need to check that for all Borel $A\subseteq Y$, we have $\gamma_{t}(A\times U)=\mu(A)$. Indeed, recall that restricted measures are concentrated on the set $\mathscr{C}$ (see Equation (9)). Thus,
$\gamma_{t}(A\times U)$ $=\gamma^{\star}\left\{\lambda_{t}^{-1}(A\times U)\right\}$
$=\gamma^{\star}\left\{(y,u_{0},y,u_{1})\mid y\in A\right\}$
$=\pi^{1}_{\#}\gamma^{\star}(A)=(\pi^{1}\circ\pi^{1,2})_{\#}\gamma^{\star}(A)$
$=\pi^{1}_{\#}\eta(A)=\mu(A)$
i.e. $\gamma_{t}(A\times Y)=\mu(A)$ as claimed.
Step two: We show that $W_{p}^{\mu}(\gamma_{t},\gamma_{s})=|t-s|W_{p}^{\mu}(\eta,\nu)$. Set $\gamma_{t}^{s}:=(\lambda_{t},\lambda_{s})_{\#}\gamma^{\star}$ for $0\leq s<t\leq 1$. We claim $\gamma_{t}^{s}\in\Pi_{Y}(\gamma_{t},\gamma_{s})$. Indeed, we have $\pi^{1,2}_{\#}\gamma_{t}^{s}=\gamma_{t}$ because for all Borel $A\subseteq Y\times U$,
$(\lambda_{t},\lambda_{s})_{\#}\gamma^{\star}\left(A\times Y\times U\right)=\gamma^{\star}\left(\lambda_{t}^{-1}(A)\right)=(\lambda_{t})_{\#}\gamma^{*}(A).$ (42)
An analogous calculation shows that $\pi_{\#}^{3,4}\gamma_{t}^{s}=\gamma_{s}$, so that $\gamma_{t}^{s}\in\Pi(\gamma_{t},\gamma_{s})$. We now check that $\gamma_{t}^{s}\in\mathcal{R}_{Y}(Y\times U)$. Indeed, suppose $E\subseteq Y\times U$ is a Borel set such that $E\cap\mathscr{C}=\varnothing$. In other words, for every $(y_{0},u_{0},y_{1},u_{1})\in E$ we have $y_{0}\neq y_{1}$. Set $D:=(\lambda_{t},\lambda_{s})^{-1}(E)$. We claim $D\cap\mathscr{C}=\varnothing$, so that
$\gamma_{t}^{s}(E)=(\lambda_{t},\lambda_{s})_{\#}\gamma^{\star}(E)$ $=\gamma^{\star}((\lambda_{t},\lambda_{s})^{-1}(E))$ (43)
$=\gamma^{\star}(D\cap\mathscr{C})=0.$ (44)
Indeed, if $c=(y,u_{0},y,u_{1})\in\mathscr{C}$, then
$(\lambda_{t},\lambda_{s})(c)=(y,tu_{0}+(1-t)u_{1},y,su_{0}+(1-s)u_{1})\notin E$ (45)
$\implies c\notin(\pi_{t},\pi_{s})^{-1}(E).$ (46)
Thus $\gamma_{t}^{s}\in\Pi_{Y}(\eta,\nu)$ as claimed. Now, we have
$W_{p}^{\mu,p}(\gamma_{t},\gamma_{s})$ $\leq\int_{(Y\times U)^{2}}d^{p}\left(y_{0},u_{0},y_{1},u_{1}\right)\,\mathrm{d}\lambda_{t}^{s}(y_{0},u_{0},y_{1},u_{1})$
$=\int_{(Y\times U)^{2}}d^{p}\left(\lambda_{t}(y_{0},u_{0},y_{1},u_{1}),\lambda_{s}(y_{0},u_{0},y_{1},u_{1})\right)\,\mathrm{d}\gamma^{\star}(y_{0},u_{0},y_{1},u_{1})$
$=\int_{(Y\times U)^{2}}\left(|(t-s)(y_{0}-y_{1})|^{2}+|(t-s)(u_{0}-u_{1})|^{2}\right)^{p/2}\,\mathrm{d}\gamma^{\star}(y_{0},u_{0},y_{1},u_{1})$
$=|t-s|^{p}\int_{(Y\times U)^{2}}d^{p}(y_{0},u_{0},y_{1},u_{1})\,\mathrm{d}\gamma^{\star}(y_{0},u_{0},y_{1},u_{1})$
$=|t-s|^{p}W_{p}^{\mu,p}(\eta,\nu).$
Conversely, an application of the previous inequality and the triangle inequality show that for $0\leq s\leq t\leq 1$,
$W_{p}^{\mu}(\eta,\nu)$ $\leq W_{p}^{\mu}(\eta,\gamma_{s})+W_{p}^{\mu}(\gamma_{s},\gamma_{t})+W_{p}^{\mu}(\gamma_{t},\nu)$ (47)
$\leq sW_{p}^{\mu}(\eta,\nu)+W_{p}^{\mu}(\gamma_{s},\gamma_{t})+(1-t)W_{p}^{\mu}(\eta,\nu).$ (48)
Rearranging the previous inequality implies $|t-s|W_{p}^{\mu}(\eta,\nu)\leq W_{p}^{\mu}(\gamma_{s},\gamma_{t})$ for all $s,t\in[0,1]$, and hence ${W_{p}^{\mu}(\gamma_{t},\gamma_{s})=|t-s|^{p}W_{p}^{\mu}(\eta,\nu)}$. ∎
When an optimal restricted coupling $\gamma^{\star}\in\Pi_{Y}(\eta,\nu)$ is induced by an injective triangular map $T^{\star}$, we may recover a constant speed geodesic in $\mathbb{P}_{p}^{\mu}(Y\times U)$, generalizing the McCann interpolant (McCann, 1997) to the conditional setting. We refer to Proposition 4 for sufficient conditions on $\eta,\nu$ under which such a $T^{\star}$ exists. Informally, samples from $(y_{0},u_{0})\sim\eta$ flow in a straight path at a constant speed to their destination $T^{\star}(y_{0},u_{0})$. We note that this theorem, while primarily about geodesics, relies on notions discussed in Section 5. In fact, this may be viewed as a special case of the results in Section 5, where we may explicitly identify the vector field.
Theorem 11 (Conditional McCann Interpolants).
Fix $\eta,\nu\in\mathbb{P}_{p}^{\mu}(Y\times U)$. Suppose $T^{\star}(y,u)=(y,T^{\star}_{\mathcal{U}}(y,u))$ is a block triangular map solving the conditional Monge problem (8). Define the maps $T_{t}:Y\times U\to Y\times U$ for $0\leq t\leq 1$ via
$T_{t}=(1-t)I+tT^{\star}\qquad(y,u)\mapsto(y,(1-t)u+tT^{\star}_{\mathcal{U}}(y,u))$ (49)
and define the curve of measures $\gamma_{t}=[T_{t}]_{\#}\eta\in\mathbb{P}_{p}^{\gamma}(Y\times U)$. Then,
  1. 1.
    $(\gamma_{t})$ is absolutely continuous and a constant speed geodesic between $\eta,\nu$
  2. 2.
    The vector field $v_{t}(T_{t}^{\star}(y,u))=(0,T_{U}^{\star}(y,u)-u)$ generates the path $\gamma_{t}$, in the sense that $(\gamma_{t},v_{t})$ solve the continuity equation (61).
Proof.
Consider the function $w_{t}:Y\times U\to U$ given by
$w_{t}(y,u)=\left(0,T^{\star}_{U}(y,u)-u\right)=(0,w_{t,U}(y,u))$ (50)
and note this is precisely $w_{t}(y,u)=\partial_{t}T^{\star}_{t}(y,u)$. Define the vector field
$v_{t}(y,u)=\left(w_{t}\circ T_{t}^{\star,-1}\right)(y,u)=\left(0,(w_{t,\mathcal{U}}\circ T_{t,\mathcal{U}}^{\star,-1})(y,u)\right).$ (51)
For any $\varphi\in\text{Cyl}(Y\times U)$, we have
$\frac{\,\mathrm{d}}{\,\mathrm{d}t}\int_{Y\times U}\varphi(y,u)\,\mathrm{d}\gamma_{t}(y,u)$ $=\frac{\,\mathrm{d}}{\,\mathrm{d}t}\int_{Y\times U}\varphi(y,u)\,\mathrm{d}[T_{t}]_{\#}\eta(y,u)$ (52)
$=\frac{\,\mathrm{d}}{\,\mathrm{d}t}\int_{Y\times U}\varphi(y,T_{t,U}^{\star}(y,u))\,\mathrm{d}\eta(y,u)$ (53)
$=\int_{Y\times U}\langle\nabla\varphi(y,T_{t,U}^{\star}(y,u),w_{t}(y,u))\rangle\,\mathrm{d}\eta(y,u)$ (54)
$=\int_{Y\times U}\langle\nabla\varphi(y,u),v_{t}(y,u)\rangle\,\mathrm{d}\gamma_{t}(y,u)$ (55)
which shows that $(\gamma_{t},v_{t})$ solve the continuity equation.
Now, note that for $0\leq a\leq b\leq 1$, we have
$\int_{a}^{b}\left\lVert v_{t}\right\rVert_{L^{p}(\gamma_{t},Y\times U)}\,\mathrm{d}t$ $=\int_{a}^{b}\left(\int_{Y\times U}\left|w_{t}\circ T_{t}^{\star,-1}\right|^{p}(y,u)\,\mathrm{d}\gamma_{t}(y,u)\right)^{1/p}\,\mathrm{d}t$ (56)
$=\int_{a}^{b}\left(\int_{Y\times U}|w_{t}|^{p}(y,u)\,\mathrm{d}\eta(y,u)\right)^{1/p}\,\mathrm{d}t$ (57)
$=\int_{a}^{b}\left(\int_{Y\times U}|u-T_{U}^{\star}(y,u)|^{p}(y,u)\,\mathrm{d}\eta(y,u)\right)^{1/p}\,\mathrm{d}t$ (58)
$=(b-a)W_{p}^{\mu}(\eta,\nu).$ (59)
In particular, $\int_{0}^{1}\left\lVert v_{t}\right\rVert_{L^{p}(\gamma_{t},Y\times U)}\,\mathrm{d}t<\infty$ and so by Theorem 15 $(\gamma_{t})$ is absolutely continuous. A similar calculation shows that $(b-a)W_{p}^{\mu}(\eta,\nu)=W_{p}^{\mu}(\gamma_{b},\gamma_{a})=\int_{a}^{b}|\gamma^{\prime}(t)|$, where the last line follows from the absolute continuity of $\gamma_{t}$. Thus, $\left\lVert v_{t}\right\rVert_{L^{p}(\gamma_{t},Y\times U)}=|\gamma^{\prime}|(t)$ for almost every $t\in[0,1]$ by Lebesgue differentiation.
5 Conditional Benamou-Brenier Theorem
In this section, we prove a characterization of all absolutely continuous curves in the space $\mathbb{P}_{p}^{\mu}(Y\times U)$. Here we assume $1<p<\infty$. First, we introduce the notion of a triangular vector field on $Y\times U$, which informally has no velocity in the $Y$ component. Triangular vector fields serve as the dynamic analog of triangular maps (6).
Definition 12 (Triangular Vector Fields).
Let $I\subset\mathbb{R}$ be an open interval. A time-dependent Borel vector field $v:I\times Y\times U\to Y\times U$ is said to be triangular if there exists a Borel vector field $v^{U}:I\times Y\times U\to U$ such that
$v_{t}(y,u)=\left(0,v_{t}^{U}(y,u)\right).$ (60)
In other words, $v_{t}=\iota\left(v_{t}^{U}\right)$ where $\iota:U\to Y\times U$ is the inclusion map.
5.1 Continuity Equation
Let $I\subset\mathbb{R}$ be an open interval. The continuity equation
$\partial_{t}\gamma_{t}+\text{div}(v_{t}\gamma_{t})=0\qquad\text{on }Y\times U\times I$ (61)
describes the evolution of a measure $\gamma_{t}$ which flows along a given vector field $v_{t}$. This equation must be understood distributionally, i.e. for every $\varphi$ in an appropriate space of test functions,
$\int_{I}\int_{Y\times U}\left(\partial_{t}\varphi(y,u,t)+\langle v_{t}(y,u),\nabla_{y,u}\varphi(y,u,t)\rangle\right)\,\mathrm{d}\gamma_{t}(y,u)\,\mathrm{d}t=0.$ (62)
We consider cylindrical test functions $\varphi\in\text{Cyl}(Y\times U\times I)$, i.e. of the form $\varphi(y,u,t)=\psi(\pi^{d}(y,u),t)$ where $\pi^{d}:Y\times U\to\mathbb{R}^{d}$ maps $(y,u)\to(\langle(y,u),e_{1}\rangle,\dots,\langle(y,u),e_{d}\rangle)$ where $\{e_{1},e_{2},\dots,e_{d}\}$ is any orthonormal family in $Y\times U$. In the finite dimensional setting, one may take $\varphi\in C_{c}^{\infty}(Y\times U)$ to be smooth and compactly supported (Ambrosio et al., 2005, Remark 8.1.1).
We now prove a lemma that will be critical to our later results. Informally, a solution to the continuity equation with a triangular vector field will result in the conditional measures almost surely satisfying the continuity equaiton as well.
Lemma 13 (Triangular Vector Fields Preserve Conditionals).
Suppose $v_{t}(y,u)=(0,v_{t}^{U}(y,u))$ is triangular and that $(\gamma_{t})\subset\mathbb{P}_{p}^{\mu}(Y\times U)$ is a path of measures such that $(v_{t},\gamma_{t})$ satisfy the continuity equation (61) in the distributional sense. Then, it follows that for $\mu$-almost every $y\in Y$, we have
$\partial_{t}\gamma_{t}^{y}+\nabla\cdot(v_{t}^{U}(y,{-})\gamma_{t}^{y})=0.$ (63)
Proof.
Fix any $\varphi\in\text{Cyl}(U\times I)$. Suppose $\psi\in\text{Cyl}(Y)$ is given, and note that $\psi(y)\varphi(u,t)\in\text{Cyl}(Y\times U\times I)$. As $(v_{t},\gamma_{t})$ solve the continuity equation, it follows from the triangular structure of $v_{t}$ that upon testing against $\psi\varphi$ we have
$\int_{I}\int_{Y}\psi(y)\int_{U}\left(\partial_{t}\varphi(u,t)+\langle v_{t}^{U}(y,u),\nabla_{u}\varphi(u,t)\right)\,\mathrm{d}\gamma_{t}^{y}(u)\,\mathrm{d}\mu(y)\,\mathrm{d}t=0.$ (64)
Because $\psi(y)\in\text{Cyl}(Y)$, it is of the form $\rho(\pi(y))$ where $\pi:Y\to\mathbb{R}^{k}$ for some $k\geq 1$ and $\rho\in C_{c}^{\infty}(\mathbb{R}^{k})$. Taking $\rho$ to be a sequence of smooth approximations to the indicator function of an arbitrary rectangle $E=E_{1}\times E_{2}\times\dots\times E_{k}\subseteq\mathbb{R}^{k}$, we see
$\int_{\pi^{-1}(E)}\int_{I}\int_{U}\left(\partial_{t}\varphi(u,t)+\langle v_{t}^{U}(y,u),\nabla_{u}\varphi(u,t)\right)\,\mathrm{d}\gamma_{t}^{y}(u)\,\mathrm{d}t\,\mathrm{d}\mu(y)=0.$ (65)
As $Y$ is separable, the Borel $\sigma$-algebra on $Y$ is generated by the cylinder sets, i.e. those which are precisely of the form $\pi^{-1}(E)$ for some finite-dimensional rectangle $E$. We have thus shown that for an arbitrary Borel measurable set $E\subseteq Y$,
$\int_{E}\int_{I}\int_{U}\left(\partial_{t}\varphi(u,t)+\langle v_{t}^{U}(y,u),\nabla_{u}\varphi(u,t)\right)\,\mathrm{d}\gamma_{t}^{y}(u)\,\mathrm{d}t\,\mathrm{d}\mu(y)=0.$ (66)
From this, it follows that
$\int_{I}\int_{U}\left(\partial_{t}\varphi(u,t)+\langle v_{t}^{U}(y,u),\nabla_{u}\varphi(u,t)\right)\,\mathrm{d}\gamma_{t}^{y}(u)\,\mathrm{d}\mu(y)\,\mathrm{d}t=0\qquad\mu\text{-almost every }y.$ (67)
We note that having $v_{t}$ be triangular is sufficient, but certainly not necessary, for the conditional continuity equation to almost surely hold. For instance, the vector field in $\mathbb{R}^{d}$ that rotates $\mathcal{N}(0,I)$ about the origin preserves all conditional distributions.
5.2 Absolutely Continuous Curves
We will show in Theorem 14 that every absolutely continuous curve $(\gamma_{t})\subseteq\mathbb{P}_{p}^{\mu}(Y\times U)$ is generated (in the sense of satisfying the continuity equation) by a triangular vector field such that ${\left\lVert v_{t}\right\rVert_{L^{p}(\gamma_{t},Y\times U)}\leq|\gamma^{\prime}|(t)}$ for almost every $t\in I$. This result is obtained as a generalization of the technique in Ambrosio et al. (2005, Theorem 8.3.1), which is an application of variational methods for PDEs.
Conversely, we show in Theorem 15 that if the pair $(\gamma_{t},v_{t})$ solve the continuity equation and $v_{t}$ is triangular, then the curve $(\gamma_{t})$ is absolutely continuous and $|\gamma^{\prime}(t)|\leq\left\lVert v_{t}\right\rVert_{L^{p}(\gamma_{t},Y\times U)}$. The main technique of this result is to study the collection of conditional continuity equations (which is feasible by Lemma 13) and to apply the converse of Ambrosio et al. (2005, Theorem 8.3.1). In this setting, the infinite-dimensional result is obtained via a finite-dimensional approximation argument.
We define the map $j_{q}:L^{q}(\gamma,Y\times U)\to L^{p}(\gamma,Y\times U)$ for $11/p+1/q=$ 1 via
$j_{q}(w)=\begin{cases}|w|^{q-2}w&w\neq 0\\ 0&w=0\end{cases}$ (68)
which is the Fréchet differential of the convex functional $\frac{1}{q}\left\lVert w\right\rVert_{L^{q}(\gamma,Y\times U)}^{q}$. A straightforward calculation shows that this map satisfies
$\left\lVert j_{q}(w)\right\rVert^{p}_{L^{p}(\gamma,Y\times U)}=\left\lVert w\right\rVert_{L^{q}(\gamma,Y\times U}^{q}=\int_{Y\times U}\langle j_{q}(w),w\rangle\,\mathrm{d}\gamma(y,u).$ (69)
Theorem 14 (AC Curves in $\mathbb{P}_{p}^{\mu}(Y\times U)$).
Let $I\subset\mathbb{R}$ be an open interval, and suppose $\gamma_{t}:I\to\mathbb{P}_{p}^{\mu}(Y\times U)$ is an absolutely continuous in the $W_{p}^{\mu}$ metric with $|\gamma^{\prime}|(t)\in L^{1}(I)$. Then, there exists a Borel vector field $v_{t}(x)$ such that
  1. 1.
    $v_{t}$ is triangular, i.e. $v_{t}(y,u)=\iota(v_{t}^{U}(y,u))$ for some Borel $v_{t}^{U}:Y\times U\to U$
  2. 2.
    $v_{t}\in L^{p}(\gamma_{t},Y\times U)$ for a.e.-$t$
  3. 3.
    $\left\lVert v_{t}\right\rVert_{L^{p}(\gamma_{t},Y\times U)}\leq|\gamma^{\prime}|(t)$ for a.e.-$t$
  4. 4.
    $(v_{t},\gamma_{t})$ solve the continuity equation distributionally.
Proof.
Assume without loss of generality that $|\gamma^{\prime}|(t)\in L^{\infty}(I)$ and that $I=(0,1)$ (Ambrosio et al., 2005, Lemma 1.1.4, Lemma 8.1.3). Fix any $\varphi\in\text{Cyl}(Y\times U)$. For $s,t\in I$ there exists an optimal triangular coupling $\gamma_{st}\in\Pi_{Y}(\gamma_{s},\gamma_{t})$. By Hölder’s inequality,
$|\gamma_{t}(\varphi)-\gamma_{s}(\varphi)|\leq\text{Lip}(\varphi)W_{p}^{\mu}(\gamma_{s},\gamma_{t}).$ (70)
It follows that $t\mapsto\gamma_{t}(\varphi)$ is absolutely continuous. We can introduce the upper semicontinuous and bounded map
$H(y_{0},u_{0},y_{1},u_{1})=\begin{cases}|\nabla\varphi(y_{0},u_{0})|&(y_{0},u_{0})=(y_{1},u_{1})\\ \frac{|\varphi(y_{0},u_{0})-\varphi(y_{1},u_{1})|}{|(y_{0},u_{0})-(y_{1},u_{1})|}&(y_{0},u_{0})\neq(y_{1},u_{1})\end{cases}.$ (71)
For $|h|$ sufficiently small, choose any optimal coupling $\gamma_{(s+h)h}\in\Pi_{Y}(\gamma_{s+h},\gamma_{s})$ and note that
$\frac{|\gamma_{s+h}(\varphi)-\gamma_{s}(\varphi)|}{|h|}$ $\leq\frac{1}{|h|}\int_{(Y\times U)^{2}}|(y_{0},u_{0})-(y_{1},u_{1})|H(y_{0},u_{0},y_{1},u_{1})\,\mathrm{d}\gamma_{(s+h)s}$ (72)
$\leq\frac{W_{p}^{\mu}(\gamma_{s+h},\gamma_{s})}{|h|}\left(\int_{(Y\times U)^{2}}H^{q}(y_{0},u_{0},y_{1},u_{1})\,\mathrm{d}\gamma_{(s+h)h,s}\right)^{1/q}.$ (73)
If $t$ is a point of metric differentiability for $t\mapsto\gamma_{t}$, note that $\gamma_{(t+h)t}\to(I,I)_{\#}\gamma_{t}$ narrowly, where $I$ is the identity map on $Y\times U$. Moreover, since $\gamma_{t}\in\mathbb{P}_{p}^{\mu}(Y\times U)$, it follows that on the diagonal we have that almost surely $H(y_{0},u_{0},y_{0},u_{1})=\iota(|\nabla_{u}\varphi(y_{0},u_{0})|$. Thus,
$\limsup_{h\to 0}\frac{|\gamma_{t+h}(\varphi)-\gamma_{t}(\varphi)|}{|h|}$ $\leq|\gamma^{\prime}|(t)\left(\int_{Y\times U}|H|^{q}(y_{0},u_{0},y_{0},u_{0})\,\mathrm{d}\gamma_{t}(y_{0},u_{0})\right)^{1/q}$ (74)
$=|\gamma^{\prime}|(t)\left\lVert\iota(\nabla_{u}\varphi)\right\rVert_{L^{q}(\gamma_{t},Y\times U)}=|\gamma^{\prime}|(t)\left\lVert\nabla_{u}\varphi\right\rVert_{L^{q}(\gamma_{t},U)}.$ (75)
Taking $Q=Y\times U\times I$ and $\gamma=\int\gamma_{t}\,\mathrm{d}t$, fix any $\varphi\in\text{Cyl}(Q)$. We have that
$\int_{Q}\partial_{s}\varphi(y,u,s)\,\mathrm{d}\gamma(y,u,s)$ (76)
$=\lim_{h\downarrow 0}\int_{I}\frac{1}{h}\left(\int_{Y\times U}\varphi(y,u,s)\,\mathrm{d}\gamma_{s}(y,u)-\int_{(Y\times U)}\varphi(y,u,s)\,\mathrm{d}\gamma_{s+h}(y,u)\right)\,\mathrm{d}s.$
An application of Fatou’s Lemma, Equation (74), and Hölder’s inequality gives us
$\left|\int_{Q}\partial_{s}\varphi(y,u,s)\,\mathrm{d}\gamma(y,u,s)\right|\leq\left(\int_{J}|\gamma^{\prime}|(s)\,\mathrm{d}s\right)^{1/p}\left(\int_{Q}|\nabla_{u}\varphi(y,u,s)|^{q}\,\mathrm{d}\mu(y,u,s)\right)^{1/q}$ (77)
for any interval $J\subset I$ with supp $\text{supp }\varphi\subset J\times Y\times U$.
Fix the subspace
$V=\left\{\iota(\nabla_{u}\varphi(y,u,s)):\varphi\in\text{Cyl}(Q)\right\}\subseteq Y\times U$ (78)
and denote by $\overline{V}$ its $L^{q}(\gamma,Y\times U\times I)$ closure. Define the linear functional $L:V\to\mathbb{R}$ via
$L(\nabla_{u}\varphi)=-\int_{Q}\partial_{s}\varphi(y,u,s)\,\mathrm{d}\gamma(y,u,s)$ (79)
and note that Equation (77) implies that $L$ is a bounded linear functional on $V$. Thus (by Hahn-Banach and the fact that $V\subseteq\overline{V}$ is dense) we may uniquely extend $L$ to $\overline{V}$. We thus have a convex minimization problem
$\min_{w\in\overline{V}}\frac{1}{q}\int_{Q}|w(y,u,s)|^{q}\,\mathrm{d}\gamma(y,u,s)-L(w)$ (80)
which admits the unique solution $w$ such that $j_{q}(w)-L=0$. In particular, the estimate (77) shows that the above functional is coercive and hence admits a minimizer which we may obtain via its differential as a consequence of convexity. Thus, we obtain a triangular vector field $v=j_{q}(w)$ such that for all $\varphi\in\text{Cyl}(Q)$,
$\langle v,\nabla\varphi\rangle=\int_{Q}\langle v(y,u,s),\nabla\varphi(y,u,s)\rangle\,\mathrm{d}\gamma(y,u,s)=\langle L,\nabla\varphi\rangle=-\int_{Q}\partial_{s}\varphi(y,u,s)\,\mathrm{d}\gamma(y,u,s).$ (81)
This precisely shows that $(v_{t},\gamma_{t})$ is a triangular distributional solution to the continuity equation.
Now, choose any interval $J\subset I$ and choose a sequence $\eta^{k}\in C^{\infty}_{c}(J)$, with $0\leq\eta^{k}\leq 1$ and $\eta_{k}\to\mathbbm{1}_{J}$ as $k\to\infty$. Moreover choose a sequence $(\nabla_{u}\varphi_{n})\subset V$ converging to $w=j_{p}(v)$ in $L^{q}(\gamma,Q)$. Our previous calculations give
$\int_{Q}\eta^{k}(s)|v(y,u,s)|^{p}\,\mathrm{d}\gamma(y,u,s)=\int_{Q}\eta^{k}(s)\langle v,w\rangle\,\mathrm{d}\gamma=\lim_{n\to\infty}\int_{Q}\eta^{k}\langle v,\nabla_{u}\varphi_{n}\rangle\,\mathrm{d}\gamma$ (82)
$=\lim_{n\to\infty}\langle L,\nabla_{u}(\eta^{k}\varphi_{n})\rangle\leq\left(\int_{J}|\gamma^{\prime}|^{p}(s)\,\mathrm{d}s\right)^{1/p}\left(\int_{J\times Y\times U}|v|^{p}\,\mathrm{d}\gamma\right)^{1/p}.$ (83)
Taking $k\to\infty$ we see that
$\int_{J}\int_{Y\times U}|v_{t}(y,u)|^{p}\,\mathrm{d}\gamma_{t}(y,u)\,\mathrm{d}t\leq\int_{J}|\gamma^{\prime}|^{p}(s)\,\mathrm{d}s$ (84)
and since $J\subset I$ was arbitrary, we conclude
$\left\lVert v_{t}\right\rVert_{L^{p}(\gamma_{t},Y\times U)}\leq|\gamma^{\prime}|(t)\qquad\text{a.e.-}t.$ (85)
Theorem 15 (Continuous Curves Generated by triangular Vector Fields).
Suppose that $\gamma_{t}:I\to\mathbb{P}_{p}^{\mu}(Y\times U)$ is narrowly continuous and $(v_{t})$ is a triangular vector field such that $(\gamma_{t},v_{t})$ solve the continuity equation with $\left\lVert v_{t}\right\rVert_{L^{p}(\gamma_{t},Y\times U)}\in L^{1}(I)$. Then, $\gamma_{t}:I\to\mathbb{P}_{p}^{\mu}(Y\times U)$ is absolutely continuous in the $W_{p}^{\mu}$ metric and $|\gamma^{\prime}|(t)\leq\left\lVert v_{t}\right\rVert_{L^{p}(\mu_{,}Y\times U)}$ for almost every $t$.
Proof.
We first assume that $U$ is finite dimensional. Our strategy is to check the hypotheses necessary for Ambrosio et al. (2005, Theorem 8.3.1) to hold for $\mu$-almost every $y$, followed by an application of this theorem. By Lemma 13, for $\mu$-almost every $y$ we have that $(\gamma_{t}^{y},v_{t}^{U}(y,{-}))$ solve the continuity equation distributionally on $I\times U$.
By Jensen’s inequality (and the assumption $p\geq 1$) we see
$\int_{I}\left\lVert v_{t}\right\rVert_{L^{p}(\gamma_{t},Y\times U)}\,\mathrm{d}t$ $=\int_{I}\mathbb{E}_{y\sim\mu}\left[\left\lVert v_{t}^{U}(y,{-})\right\rVert_{L^{p}(\gamma_{t}^{y},U)}^{p}\right]^{1/p}\,\mathrm{d}t$ (86)
$\geq\int_{I}\mathbb{E}_{y\sim\mu}\left[\left\lVert v_{t}^{U}(y,{-})\right\rVert_{L^{p}(\gamma_{t}^{y},U)}\right]\,\mathrm{d}t$ (87)
$=\mathbb{E}_{y\sim\mu}\left[\int_{I}\left\lVert v_{t}^{U}(y,{-})\right\rVert_{L^{p}(\gamma_{t}^{y},U)}\,\mathrm{d}t\right].$ (88)
Since the first term is finite, it follows that
$\left\lVert v_{t}^{U}(y,{-})\right\rVert_{L^{p}(\gamma_{t}^{y},U)}\in L^{1}(I)\qquad\mu\text{-almost every }y.$ (89)
Now Ambrosio et al. (2005, Lemma 8.1.2) shows that for $\mu$-almost every $y$ we have that $(\gamma_{t}^{y})$ admits a narrowly continuous representative $(\tilde{\gamma}_{t}^{y})$ with $\tilde{\gamma}^{y}_{t}=\gamma^{y}_{t}$ for almost every $t$. It follows from Ambrosio et al. (2005, Theorem 8.3.1) that for any $t_{1}\leq t_{2}$ in $I$, we have
$W_{p}^{p}(\tilde{\gamma}^{y}_{t_{1}},\tilde{\gamma}^{y}_{t_{2}})$ $\leq(t_{2}-t_{1})^{p-1}\int_{t_{1}}^{t_{2}}|v_{t}^{U}(y,u)|^{p}\,\mathrm{d}\tilde{\gamma}_{t}^{y}(u)\,\mathrm{d}t$ (90)
$=(t_{2}-t_{1})^{p-1}\int_{t_{1}}^{t_{2}}|v_{t}^{U}(y,u)|^{p}\,\mathrm{d}{\gamma}_{t}^{y}(u)\,\mathrm{d}t$ (91)
where the second line follows as $\tilde{\gamma}^{y}_{t}=\gamma_{t}^{y}$ for almost every $t$.
Let $\tilde{\gamma}_{t}=\int_{Y}\tilde{\gamma}^{y}_{t}\,\mathrm{d}\mu(y)$ be the measure obtained via marginalizing over the $Y$-variables. Taking an expectation over $y\sim\mu$, the previous inequality shows us that
$\frac{W_{p}^{\mu,p}(\tilde{\gamma}_{t_{1}},\tilde{\gamma}_{t_{2}})}{(t_{2}-t_{1})^{p}}\leq\frac{1}{t_{2}-t_{1}}\int_{t_{1}}^{t_{2}}\left\lVert v_{t}\right\rVert_{L^{p}(\gamma_{t},Y\times U)}^{p}\,\mathrm{d}t.$ (92)
Now, note that $t_{1}$ is almost surely a Lebesgue point of the right-hand side and $\tilde{\gamma}_{t_{1}}=\gamma_{t_{1}}$. Taking $t_{2}\to t_{1}$ along a sequence where $\tilde{\gamma}_{t_{2}}=\gamma_{t_{2}}$ shows us that
$|\gamma^{\prime}|(t)\leq\left\lVert v_{t}\right\rVert_{L^{p}(\gamma_{t}),Y\times U}$ (93)
for almost every $t\in I$.
In the case that $U$ is infinite dimensional, fix any $y\in Y$ such that Lemma 13 holds (which is of full measure) and fix a countable orthonormal basis $(e_{k})$ for $U$. Set $\pi^{d}:U\to\mathbb{R}^{d}$ to be the projection operator for this basis, i.e. $u\mapsto(\langle u,e_{1}\rangle,\dots,\langle u,e_{d}\rangle)$. We consider the collection of finite dimensional conditional measures $\gamma_{t}^{d,y}=\pi^{d}_{\#}\gamma_{t}^{y}$. By the same argument in Ambrosio et al. (2005, Theorem 8.3.1), there exists a vector field $v_{t}^{d,y}$ on $\mathbb{R}^{d}$ such that $(\gamma_{t}^{d,y},v_{t}^{d,y})$ solve the continuity equation and
$\left\lVert v_{t}^{d,y}\right\rVert_{L^{p}(\gamma_{t}^{d,y},\mathbb{R}^{d})}\leq\left\lVert v_{t}^{U}(y,{-})\right\rVert_{L^{p}(\gamma_{t}^{y},U)}.$ (94)
It follows from the finite-dimensional case above that for almost every $t_{1}\leq t_{2}$, we have
$W_{p}^{p}(\gamma_{t_{1}}^{d,y},\gamma_{t_{2}}^{d,y})\leq(t_{2}-t_{1})^{p-1}\int_{t_{1}}^{t_{2}}\left\lVert v_{t}^{U}(y,{-})\right\rVert_{L^{p}(\gamma_{t}^{y},U)}^{p}\,\mathrm{d}t.$ (95)
Let $\hat{\gamma}_{t}^{y,d}=(\pi^{d})_{\#}^{\star}\gamma_{t}^{y,d}$ where $(\pi^{d})^{\star}:\mathbb{R}^{d}\to U$ maps $z\mapsto\sum_{k=1}^{d}z_{k}e_{k}$. As $d\to\infty$ we have $\hat{\gamma}_{t}^{d,y}\to\gamma_{t}^{y}$ narrowly for all $t\in I$. Since $(\pi^{d})^{\star}$ is an isometry, Ambrosio et al. (2005, Lemma 7.1.4) shows that
$W_{p}^{p}(\gamma_{t_{1}}^{y},\gamma_{t_{2}}^{y})\leq\liminf_{d\to\infty}W_{p}^{p}(\gamma_{t_{1}}^{d,y},\gamma_{t_{2}}^{d,y})\leq(t_{2}-t_{1})^{p-1}\int_{t_{1}}^{t_{2}}\left\lVert v_{t}^{U}(y,{-})\right\rVert_{L^{p}(\gamma_{t}^{y},U)}^{p}\,\mathrm{d}t.$ (96)
Now, integration with respect to $\,\mathrm{d}\mu(y)$ yields
$W_{p}^{p,\mu}(\gamma_{t_{1}},\gamma_{t_{2}})\leq(t_{2}-t_{1})^{p-1}\int_{t_{1}}^{t_{2}}\left\lVert v_{t}\right\rVert^{p}_{L^{p}(\gamma_{t},Y\times U)}\,\mathrm{d}t.$ (97)
Taking $t_{2}\to t_{1}$ shows that for almost every $t$ we have
$|\gamma^{\prime}|(t)\leq\left\lVert v_{t}\right\rVert_{L^{p}(\gamma_{t},Y\times U)}.$ (98)
As a corollary of Theorem 14 and Theorem 15, we obtain a conditional version of the Benamou-Brenier theorem (Benamou and Brenier, 2000). We note that the proof of Theorem 16 largely follows the unconditional case (see e.g. Ambrosio et al. (2005, Chapter 8)), but we include it for the sake of completeness.
Theorem 16 (Conditional Benamou-Brenier).
Suppose $Y,U$ are separable Hilbert spaces, $1<p<\infty$, and $\mu\in\mathbb{P}(Y)$. Then, for any $\eta,\nu\in\mathbb{P}_{p}^{\mu}(Y\times U)$, we have
[…]
Proof.
Write $M$ for the infimum on the right-hand side.
First, suppose that $(v_{t},\mu_{t})$ are admissible and $\int_{0}^{1}\left\lVert v_{t}\right\rVert_{L^{p}(\mu_{t})}<\infty$. It follows from Theorem 15 that $(\gamma_{t})$ is an absolutely continuous curve in $\mathbb{P}_{p}^{\mu}(Y\times U)$ and $\left\lVert v_{t}\right\rVert_{L^{p}(\mu_{t},Y\times U)}\geq|\gamma^{\prime}|(t)$. Thus,
$W_{p}^{\mu,p}(\eta,\nu)\leq\left(\int_{0}^{1}|\gamma^{\prime}|(t)\,\mathrm{d}t\right)^{p}\leq\int_{0}^{1}\left\lVert v_{t}\right\rVert_{L^{p}(\mu_{t},Y\times U)}^{p}\,\mathrm{d}t\leq M.$ (100)
Conversely, by Theorem 10 there exists a constant speed geodesic $(\gamma_{t})\subset\mathbb{P}_{p}^{\mu}(Y\times U)$ connecting $\eta$ and $\nu$. Recall that constant speed geodesics are absolutely continuous. By Theorem 14, there exists a Borel triangular vector field $v_{t}$ such that $(v_{t},\gamma_{t})$ solve the continuity equation, and moreover ${\left\lVert v_{t}\right\rVert_{L^{p}(\mu_{t},Y\times U)}\leq|\gamma^{\prime}|(t)}$. In fact, because $(v_{t},\gamma_{t})$ solve the continuity equation, Theorem 15 yields that $\left\lVert v_{t}\right\rVert_{L^{p}(\mu_{t},Y\times U)}=|\gamma^{\prime}|(t)$.
Since $\gamma_{t}$ is a constant speed geodesic in $\mathbb{P}_{p}^{\mu}(Y\times U)$, it follows that $|\mu^{\prime}|(t)=W_{p}^{\mu}(\eta,\nu)$ for almost every $t\in(0,1)$. Hence,
$W_{p}^{\mu,p}(\eta,\nu)=\int_{0}^{1}|\gamma^{\prime}|(t)^{p}\,\mathrm{d}t=\int_{0}^{1}\left\lVert v_{t}\right\rVert^{p}_{L^{p}(\gamma_{t},Y\times U)}\geq M.$ (101)
Thus, $W_{p}^{\mu,p}(\eta,\nu)=M$ as desired. ∎
6 COT Flow Matching
Figure 1: Results for the Darcy flow inverse problem. We compare the posterior mean and variance of our method (COT-FM) to pCN (Cotter et al., 2013) and WaMGAN (Hosseini et al., 2023).
We have thus far seen that the COT problem (11) admits a dynamical formulation (99), where one may take the underlying vector fields to be triangular. We use these results to design a principled model for conditional generation based on flow matching (Lipman et al., 2022; Albergo et al., 2023b). More specifically, recent work (Tong et al., 2023; Pooladian et al., 2023a) has proposed to couple the source and target measures in flow matching via optimal transport. We extend these techniques to the conditional setting.
In this section we make several simplifying assumptions. First, we use the squared-distance cost $|(y_{0},u_{0})-(y_{1},u_{1})|^{2}$ throughout (i.e. $p=2$). Second, we assume that $Y$ and $U$ are finite dimensional Euclidean spaces, but we emphasize that our techniques may be applied to infinite dimensional spaces under additional regularity assumptions (Kerrigan et al., 2023, 2022; Lim et al., 2023; Baldassari et al., 2024). Third, we assume that our source and target measures admit densities with respect to the Lebesgue measure on $Y\times U$. This is sufficient (Prop. 4) for the existence of a unique COT plan which is induced by a triangular map. We will henceforth, in an abuse of notation, identify measures with their densities.
Flow Matching.
We assume that we have access to samples $z_{0}=(y_{0},u_{0})\sim\eta(y_{0},u_{0})\in\mathbb{P}_{p}^{\mu}(Y\times U)$ from a source measure, and samples $z_{1}=(y_{1},u_{1})\sim\nu(y_{1},u_{1})\in\mathbb{P}_{p}^{\mu}(Y\times U)$ from a target measure. Let $z=(z_{0},z_{1})\sim\rho(z_{0},z_{1})\in\Pi(\eta,\nu)$ be any coupling of the source and target measure. Following Tong et al. (2023), we specify a collection of measures and vector fields via
$\gamma_{t}(y,u\mid z)=\mathcal{N}\left(y,u\mid tz_{1}+(1-t)z_{0},C\right)$ (102)
$u_{t}(y,u\mid z)=z_{1}-z_{0}$ (103)
where $C$ is any trace-class covariance operator (Da Prato and Zabczyk, 2014), e.g. $C=\sigma^{2}I$.
As is standard in flow matching (Lipman et al., 2022; Tong et al., 2023; Kerrigan et al., 2023), we obtain from Equations (102) and (103) a marginal measure $\gamma_{t}(y,u)$ and vector field $u_{t}(y,u)$ satisfying the continuity equation via
$\gamma_{t}(y,u)=\int_{(Y\times U)^{2}}\gamma_{t}(y,u\mid z)\,\mathrm{d}\rho(z)$ (104)
$u_{t}(y,u)=\int_{(Y\times U)^{2}}u_{t}(y,u\mid z)\frac{\,\mathrm{d}\gamma_{t}(y,u\mid z)}{\,\mathrm{d}\gamma_{t}(y,u)}\,\mathrm{d}\rho(z).$ (105)
To model the intractable $u_{t}(y,u)$, we minimize the loss [^1] This is often called the conditional flow matching loss (Tong et al., 2023), which is not to be confused with the notion of conditioning that we focus on in this work.
$\mathcal{L}(\theta)=\mathbb{E}_{t,\rho(z),\gamma_{t}(y,u\mid z)}\left\lVert u^{\theta}(t,y,u)-u_{t}(y,u\mid z)\right\rVert^{2}$ (106)
which has the same gradient (with respect to $\theta$ as the MSE loss to the true vector field of interest $u_{t}(y,u)$ (Tong et al., 2023, Theorem 3.2).
COT Flow Matching.
Figure 2: Further visualizations of our results on the Darcy flow inverse problem. The first row shows the ground-truth posterior. Row two shows samples obtained from out COT-FM method, and row three shows the same for WaMGAN (Hosseini et al., 2023). The prior and conditioning information are shown in the final two rows.
In the preceding section, $\rho(z)$ may be an arbitrary coupling between $\eta$ and $\nu$. Motivated by Propositions 1 and 4, we will choose $\rho$ to be a conditional optimal coupling, so that $(y_{0},u_{0},y_{1},u_{1})\sim\rho$ has the form $\left(y,u,y,T_{U}(y,u)\right)$. By Theorem 16 and Theorem 11, under such a $\rho$ the vector fields in Equation (103) and Equation (105) will be triangular. Thus, we parametrize our model $u^{\theta}$ to also be triangular. If the covariance operator $C$ in Equation (102) is taken such that $\text{tr}(C)\to 0$, then we recover the optimal path of measures given by the conditional Benamou-Brenier Theorem via a pointwise application of (Tong et al., 2023, Proposition 3.4).
Given a collection of samples $\{z_{0}^{i},z_{1}^{i}\}_{i=1}^{n}$ drawn from $\eta$ and $\nu$, we approximate a conditional optimal coupling $\rho$ using standard numerical techniques (e.g., those available in the POT package (Flamary et al., 2021)) with the cost function
$c_{\epsilon}(y_{0},u_{0},y_{1},u_{1})=|y_{1}-y_{0}|^{2}+\epsilon|u_{1}-u_{0}|^{2}$ (107)
for some $0<\epsilon\ll 1$. Intuitively, such a cost penalizes mass transfer along the $Y$ dimension, which is precisely the constraint sought in the COT problem (11). Hosseini et al. (2023, Prop. 3.11) show that as $\epsilon\downarrow 0$, we recover the true optimal triangular map. When our datasets are not too large, we precompute this COT coupling prior to training our flow matching model. Otherwise, we compute this COT coupling for each minibatch drawn at training time.
After training, we obtain a learned triangular vector field $u^{\theta}_{t}(y,u)$. Given an arbitrary fixed $y\in Y$, we may approximately sample from the target $\nu(u\mid y)$ by sampling $u_{0}\sim\eta(u_{0}\mid y)$ and numerically solving the corresponding differential equation. For some tasks, like domain translation, note that we are given a fixed $u_{0}$ of interest rather than sampling $u_{0}$ from the source. In these cases, we are effectively sampling from $\nu(u\mid y,u_{0})$.
Source Measure.
Our framework is agnostic to the choice of source measure $\eta$, allowing for great flexibility in the modeling process. The main requirement is that the $Y$-marginal of $\eta$ must match the $Y$-marginal of the target $\nu$. In some scenarios, this is trivially satisfied. For instance, if one is interested in using a source distribution which is simply random noise, one may take $\eta(y_{0},u_{0})=\pi^{Y}_{\#}\nu(y_{0})\eta_{U}(u_{0})$ to be the product of two independent distributions where $\eta_{U}(u)$ is arbitrary, e.g. Gaussian noise. Sampling from such an $\eta$ can be achieved by sampling noise from $\eta_{U}(u)$ and an independent $y$ from the target measure. We explore more complex choices for $\eta$ in Section 7.
7 Experiments
The following section is devoted to illustrate our methodology on a variety of different problems related to conditional simulation. First, we explore a likelihood-free inference problem in function space on a Darcy flow inverse problem. The dataset used in this case mirrors those in Hosseini et al. (2023), as we aim to directly compare with their Conditional Wasserstein Monge GAN (WaMGAN) method. In addition, we demonstrate our method on two unpaired image-to-image translation tasks (Nazarpour and Chen, 2017; Choi et al., 2020)
7.1 Bayesian Inverse Problems
In this experiment, we consider simulation-based posterior inference on the parameters characterizing a stochastic forward model, whose likelihood evaluation is intractable.
Darcy Flow
In this experiment the forward model corresponding to the 2D Darcy flow, an ellipitic PDE on a smooth domain $\Omega\subset\mathbb{R}^{d}$ characterized by a permeability field $\exp(u)$, a pressure field $\rho$, and a source term $f$:
$-\text{div}\exp(u)\nabla\rho=f\quad\text{in}\quad\Omega,\qquad\rho=0\quad\text{on}\quad\partial\Omega.$ (108)
For simplicity we denote the deterministic part of the forward model as $\mathfrak{F}:U\rightarrow Y$, and the stochasticity arises from only having access to noisy measurements $y=\mathfrak{F}(u)+\epsilon\in Y,$ where $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$. Although the general setting of the experiment corresponds to the one in Hosseini et al. (2023), we consider the more general task of performing posterior inference for $u$ when both $u$ and $y$ are observed at arbitrary resolutions, effectively posing the problem in function space. Hence, we consider $Y\subseteq\mathcal{Y}$, $U\subseteq\mathcal{U}$, where $\mathcal{Y},\mathcal{U}$ are infinite-dimensional Hilbert spaces. We remark that, although the measurements $Y$ are observed on a grid their resolution does not need to be fixed, allowing for training on observations at different resolutions. It follows that $\epsilon$ follows a Gaussian measure on $Y$ with a white noise kernel. Notice that although the conditional probability measure $\nu^{u}:=\pi^{U}_{\#}\nu$ is Gaussian, evaluating it requires simulating from Equation (108).
Figure 3: Zero-shot super-resolution on the Darcy flow inverse problem. Our model (COT-FM) is trained using a Fourier Neural Operator (Li et al., 2020), allowing us to generate samples at higher resolutions than available at training time. We compare to WaMGAN (Hosseini et al., 2023) trained with the same architechture.
We assign to $u_{1}$ a prior corresponding to the Gaussian measure $\mathcal{N}(0,C_{1})$, where $C_{1}$ corresponds to a symmetric, non-negative and trace-class covariance operator (Da Prato and Zabczyk, 2014). In order to obtain a dataset $\mathcal{D}_{1}=\{y_{1,i},u_{1,i}\}_{i=1}^{n}$ of samples on the target domain, we simulate $y_{1}$ from Equation (108) parameterizing the forward model with samples $u_{1}\sim\mathcal{N}(0,C_{1})$. We denote the target measure as $\nu$. We construct the source measure $\eta$ as the product measure $\mathcal{N}(0,C_{0})\times\mu$, where $C_{0}$ is symmetric, non-negative, trace-class and $\mu=\pi^{Y}_{\#}\eta$ corresponds to the marginal data distribution. The choice of picking $C_{0}\neq C_{1}$, while complicating the problem, reflects a scenario where we want to perform inference with a weakly informative prior that does not reflect the true data generating process. Notice that sampling from a conditional coupling $\gamma\in\Pi_{Y}(\eta,\nu)$ would still recover samples from the true posterior. Here, we assume that $C_{0},C_{1}$ are sufficiently regular such that the path of measures in Equation (104) is well-defined.
[…]
Figure 4: Samples from our model on a supervised image-to-image translation task, where translation is performed from Chinese MNIST to MNIST (a) and from MNIST to Chinese MNIST (b). This demonstrates that our method is able to use non-trivial source distributions and is able to successfully draw samples from the corresponding conditional measure.
The resulting amortized sampler, denoted for simplification by the mapping $(y,u_{0})\mapsto u_{1}=\tilde{T}_{U}(y,u_{0})$, will parameterize an approximate posterior measure. Notice that, in contrast to variational inference techniques, no distributional assumptions are made about the approximate posterior. In turn, integrals are obtained numerically by Monte Carlo sampling $K-$many samples from the prior, resulting in the approximation
[…]
In Figure 1, we show an example of how our method compares to pCN and WaMGAN on recovering the posterior mean and variance on a given example. For a fair comparison and to ensure resolution-invariance, we used the same architecture for both neural models, with the discriminator having an additional layer projecting to a scalar value. However, the WaMGAN method used twice as many parameters as it requires two neural networks. Each statistic is computed on $2\times 10^{3}$ samples from each method. We showcase the ability of our method to perform zero-shot super-resolution in Figure 3. A random selection of samples from our method and from WaMGAN, as well as the prior sample and the different conditions used for generation, are displayed in Figure 2. Qualitatively, our samples appear to better match the true posterior compared to WaMGAN (Hosseini et al., 2023), while being cheaper to sample than pCN (Cotter et al., 2013).
7.2 Unpaired Image-to-Image Translation
Figure 5: Uncurated sample from the test set of the Animal Faces HQ dataset (Choi et al., 2020) from the dogs class, translated to cats. These results demonstrate our method on an unsupervised image-to-image task, where we utilize a pre-trained content extractor.
We now demonstrate our method on an image-to-image translation task. In particular, in this section our source measures are distributions over images. The objective of domain translation is to convert information from one domain to another by adjusting domain-specific attributes, denoted style, while preserving the domain-invariant ones, denoted content (Isola et al., 2017). Semantic attributes can be associated with human-assigned labels (supervised) or inferred by a learning algorithm that learns domain-invariant representations of the data (unsupervised). Pseudocode is available in Appendix A.
Historically, the problem of domain translation was tackled through methods that heavily relied on parallel datasets—pairs of corresponding samples from the source and target domains. However, obtaining such paired datasets is often impractical or impossible in many real-world applications, prompting the shift towards methods that do not require paired data (unpaired domain translation) (Pang et al., 2021). Domain translation is possible for a learning algorithm if it satisfies cycle consistency, positing that translation of a source sample $u_{s}$ into a target sample $\hat{u}_{t}$ and then back to the source domain should recover $\hat{u}_{s}=u_{s}$, and the same shall hold by starting from a target sample. Notice that this property is achieved naturally by our model, and hence does not require additional terms in the loss as is the case for GAN- or VAE-based methods (Zhu et al., 2017; Isola et al., 2017; Liu et al., 2017).
Here, the space $U$ corresponds to the space of images. For the space $Y$, we explore two choices. First, if supervised information (i.e., a class label) is available, we take $Y$ to be the space of labels. In this setting, if the class distribution in both the source and target distribution is balanced, then the $Y$-marginals of the source and target measures match. When such information is not available, we extract semantic information from our images using a pre-trained encoder, mapping images to a latent space $Y$. In this case, we assume the encoding distribution matches across the source and target distributions. Exploring techniques for balancing this distribution is an interesting avenue for future work.
Supervised.
We showcase in Figure 4 the ability of our method to address the task of class-conditional image translation, using the MNIST and the Chinese MNIST (Nazarpour and Chen, 2017) datasets. The second dataset was preprocessed in order to center the images around their content, and both are reshaped to a resolution of $32\times 32$. The size of the training set was small enough for us to compute the empirical COT plan on the entire dataset, and pairs were resampled at each epoch. We parameterize the neural network in COT-FM with a UNet (Ronneberger et al., 2015), and we compute the $Y-$distance in the COT plan by taking the $L_{2}$ distance between one-hot encoded vectors for the labels. The UNet has hidden dimensionality of 128, and we use three downsampling/upsampling blocks with multipliers $(1,2,2)$. The optimizer is Adam (Kingma and Ba, 2014) with a learning rate of $10^{-4}$, at a batch size of 128 for $3\times 10^{5}$ training steps.
Unsupervised.
Samples are mapped into a shared representation space by a pre-trained feature extractor $\mathfrak{F}:\mathbb{R}^{d}\rightarrow[0,1]^{e}$, obtaining content variables $y=\mathfrak{F}(u)$. These effectively act as pseudo-labels when building the empirical COT matrix. In our example, we perform translation from the dogs class to the cats class in the AnimalFacesHQ dataset (Choi et al., 2020) and we use CLIP as our feature extractor (Radford et al., 2021) as found in the official Hugging Face repository in its version clip-vit-large-patch14. The training set is comprised of 5067 images of cats and 4680 images of dogs, enabling the computation of the COT plan in advance. As with the previous experiment, samples are repaired by sampling from the COT plan at each epoch. Images are observed at a native resolution of $512\times 512$, but to ease the memory requirements we employ another pre-trained encoder $\mathfrak{E}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{z}$ for dimensionality reduction, effectively performing flow matching in latent space (Dao et al., 2023). The procedure is detailed in Algorithm 1. We take the encoder $\mathfrak{E}$ to be the variational autoencoder (Kingma and Welling, 2013) from Stable Diffusion v2.1, enabling training of the COT-FM network at a dimensionality of $4\times 64\times 64$. We use the Adam optimizer (Kingma and Ba, 2014) with a learning rate of $2\times 10^{-4}$, with a batch size of 32 for $4\times 10^{5}$ training steps. We show several uncurated conditional samples from our model in Figure 5.
8 Conclusion
We analyze conditional optimal transport from a geometric and dynamical point of view. Our analysis culminates in the characterization of absolutely continuous curves of measures in a conditional Wasserstein space, resulting in a conditional analog of the Benamou-Brenier Theorem. We use these result to build on the framework of triangular transport and flow matching to develop simulation-free methods for conditional generative models. Our methods are applicable across a wide class of problems, and we demonstrate our methodology on an infinite-dimensional Bayesian inverse problem as well as image translation tasks.
Acknowledgments
This research was supported by the Hasso Plattner Institute (HPI) Research Center in Machine Learning and Data Science at the University of California, Irvine, by the National Science Foundation under award 1900644, and by the National Institutes of Health under award R01-LM013344.
Appendix A Pseudocode
1:  Input: Data $\{u_{0,i},u_{1,i}\}_{i=1}^{B}\sim\gamma^{\star}_{COT}$, pretrained content extractor $\mathfrak{F}$, interpolant network $v_{\theta}$, pretrained encoder $\mathfrak{E}$ (optional).
2:  Initialize: $\mathcal{L}=0$
3:  $\{y_{0,i},y_{1,i}\}_{i=1}^{B}\leftarrow\{\mathfrak{F}(u_{0,i}),\mathfrak{F}(u_{1,i})\}_{i=1}^{B}$
4:  if  $\mathfrak{E}$ is given  then
5:     $\{u_{0,i},u_{1,i}\}_{i=1}^{B}\leftarrow\{\mathfrak{E}(u_{0,i}),\mathfrak{E}(u_{1,i})\}_{i=1}^{B}$
6:  end if
7:  for $i=1,\dots,B$ do
8:     Sample $t_{i}\sim\mathcal{U}([0,1])$
9:     Interpolate: $u_{t,i}=(1-t_{i})u_{0,i}+t_{i}u_{1,i}$
10:     Evaluate interpolant network:
$\hat{v}_{t_{i}}^{0}=v_{\theta}(u_{t,i},t_{i},y_{0,i}),\quad\hat{v}_{t_{i}}^{1}=v_{\theta}(u_{t,i},t_{i},y_{1,i})$
[…]
12:  end for
13:  $\theta_{\text{new}}\leftarrow$ OptimizerStep$(\nabla_{\theta}\mathcal{L})$
Algorithm 1 DCOT Flow for Unsupervised Image-to-Image Translation, Training Step

### v2

@@
-Dynamic Conditional Optimal Transport through Simulation-Free Flows
+Dynamical Conditional Optimal Transport through Simulation-Free Flows
 Abstract
 We study the geometry of conditional optimal transport (COT) and prove a dynamical formulation which generalizes the Benamou-Brenier Theorem.
-With these tools, we propose a simulation-free flow-based method for conditional generative modeling.
-Our method couples an arbitrary source distribution to a specified target distribution through a triangular COT plan.
-We build on the framework of flow matching to train a conditional generative model by approximating the geodesic path of measures induced by this COT plan.
-Our theory and methods are applicable in the infinite-dimensional setting, making them well suited for inverse problems.
-Empirically, we demonstrate our proposed method on two image-to-image translation tasks and an infinite-dimensional Bayesian inverse problem.
+Equipped with these tools, we propose a simulation-free flow-based method for conditional generative modeling.
+Our method couples an arbitrary source distribution to a specified target distribution through a triangular COT plan, and a conditional generative model is obtained by approximating the geodesic path of measures induced by this COT plan.
+Our theory and methods are applicable in infinite-dimensional settings, making them well suited for a wide class of Bayesian inverse problems.
+Empirically, we demonstrate that our method is competitive on several challenging conditional generation tasks, including an infinite-dimensional inverse problem.
 1 Introduction
-Many fundamental tasks in machine learning and statistics may be posed as estimating or sampling from a distribution, where one only has access to the distribution via a finite set of observations.
-Such tasks range from generative modeling and density estimation to Bayesian inference.
-A broad range of methods for sampling from an intractable distribution $p$ may be seen through the lens of measure transport.
-Intuitively, one fixes a reference distribution $q$ which is easy to sample from, and samples $z\sim q$ are transformed by a learned mapping $T$ such that $T(z)$ is approximately distributed as $p$.
-The distribution obtained in this way from $q$, denoted $T_{\#}q$, is the pushforward of $q$ along $T$.
-Generative models (Ho et al., 2020; Song et al., 2020; Kingma and Welling, 2013; Goodfellow et al., 2014; Papamakarios et al., 2021) typically find such a mapping $T$ by minimizing a divergence between the target and generated distribution.
-In the standard maximum-likelihood based setting, the mapping $T$ is learned by minimizing (an upper bound on) the KL divergence between $p$ and the pushforward $T_{\#}q$.
-However, a natural alternative approach, based on optimal transport (Villani et al., 2009; Santambrogio, 2015), is to search for a map $T$ which satisfies the constraint $T_{\#}q=p$ and has minimal cost (in some sense).
-In many scenarios, we are interested in conditional sampling.
-That is, we have access to some observation $y\in Y$, and from this information we would like to infer information regarding an unknown quantity $u\in U$.
-For instance, in Bayesian inference, $y$ denotes the data and $u$ the parameters of a model.
-A joint distribution $p(y,u)=p(y\mid u)p(u)$ is specified via a likelihood $p(y\mid u)$ and a prior $p(u)$, and one would like to draw samples from the posterior $p(u\mid y)$.
-However, in applications such as Bayesian inverse problems (Dashti and Stuart, 2013), one is interested in approximating $p(u\mid y)$ for many different observations $y$.
-Moreover, in applications such as simulation-based inference (Cranmer et al., 2020), the likelihood function may be intractable while we are still able to sample from $p(y,u)$.
-This motivates an amortized and likelihood-free (Amos et al., 2023) approach, where a single model may be trained and applied across different observations.
-In the framework of measure transport, we first specify a source distribution $q(u)$.
-For any given, fixed $y\in Y$, we seek to find a mapping $T(y)$, depending on $y$, such that $T(y)_{\#}q(u)\approx p(u\mid y)$.
-Finding such a mapping for each individual $y$ via optimal transport, though, is typically intractable, as we frequently do not have samples from the posterior $p(u\mid y)$.
-As such, we must leverage information from other observations.
-One way to achieve this is through triangular mappings (Baptista et al., 2020), where a joint source distribution $q(y,u)$ is transformed by a mapping of the form $T:(y,u)\mapsto(T_{Y}(y),T_{U}(y,u))$.
-It can be shown (see Section 3.2) that if $T_{\#}q=p$, then $T_{U}(y,{-})$ almost surely couples $q(u\mid y)$ and $p(u\mid y)$.
-Thus, if we are able to find such a triangular $T$ using samples from the joint distribution $p(y,u)$, then we may obtain conditional samplers.
-In this work, we develop a dynamical framework for conditional optimal transport (COT).
-In particular, we study absolutely curves of measures in a conditional Wasserstein space over $Y\times U$.
-Our main results (Theorem 14 and Theorem 15) show that such curves are, informally stated, generated by the flow induced by a triangular vector field.
-Intuitively, a triangular vector field on $Y\times U$ has no velocity in the $Y$ component and thus induces a flow for each conditional measure.
-As a consequence, in Theorem 16 we obtain a conditional version of the well-known Benamou-Brenier theorem (Benamou and Brenier, 2000).
-Importantly, our framework is applicable in infinite-dimensional spaces, making it well suited for function space inference problems.
-We apply our results on dynamical COT to design conditional generative models, based on the flow matching framework (Lipman et al., 2022; Albergo et al., 2023a) .
-In particular, we generalize the methods of Pooladian et al.
-(2023a) and Tong et al.
-(2023), which utilize optimal transport to pair source and target samples, to the conditional setting.
-Our method works by solving a relaxed optimal transport problem (Hosseini et al., 2023) during training.
-In this way, we amortize solutions to the dynamic COT problem across arbitrary $y$.
-Our method may use arbitrary source distributions, and is applicable across a wide range of conditional generative modeling tasks.
-Specifically, the main contributions of our work are as follows.
+Many fundamental tasks in machine learning and statistics may be posed as modeling a conditional distribution $\nu(u\mid y)$, but where obtaining an analytical representation of $\nu(u\mid y)$ is often impractical.
+While sampling-based approaches such as Markov Chain Monte Carlo (MCMC) methods are useful, they suffer from several limitations.
+First, MCMC requires numerous likelihood evaluations, rendering it prohibitively expensive in scientific and engineering applications where the likelihood is determined by an expensive numerical simulator.
+Second, MCMC must be run anew for every observation $y$, which is impractical in applications such as Bayesian inverse problems (Dashti and Stuart, 2013) and generative modeling (Mirza and Osindero, 2014).
+These limitations motivate the need for a likelihood-free (Cranmer et al., 2020) and amortized (Amos et al., 2023) approach.
+While methods like ABC (Beaumont, 2010) and variational inference (Blei et al., 2017) address these challenges, they are difficult to scale to high dimensions or have limited flexibility.
+Recently, generative models such as normalizing flows (Papamakarios et al., 2019, 2021), GANs (Ramesh et al., 2022), and diffusion models (Sharrock et al., 2022) have shown promise in amortized and likelihood-free inference.
+These models may be viewed in the framework of measure transport (Baptista et al., 2020), where samples $u\sim\eta(u)$ from a tractable source distribution are transformed by a mapping $T(y,u)$ such that the transformed samples are approximately distributed as $\nu(u\mid y)$.
+One way to achieve this is through triangular mappings (Baptista et al., 2020; Spantini et al., 2022), where a joint source distribution $\eta(y,u)$ is transformed by a mapping of the form $T:(y,u)\mapsto(T_{Y}(y),T_{U}(y,u))$.
+Under suitable assumptions, if $T$ transforms the source $\eta(y,u)$ into the target $\nu(y,u)$, then $T_{U}(y,{-})$ couples the conditionals $\eta(u\mid y)$ and $\nu(u\mid y)$.
+Typically, such a map $T$ is not unique (Wang et al., 2023), and a natural idea is thus to regularize the transport and search for an admissible mapping that is in some sense optimal.
+In other words, learning a conditional sampler may be phrased as finding a conditional optimal transport (COT) map.
+While there exists some work on learning COT maps, these approaches often rely on a difficult adversarial optimization problem (Baptista et al., 2020; Hosseini et al., 2023; Bunne et al., 2022; Ray et al., 2022) or simulating from the model during training (Baptista et al., 2023; Wang et al., 2023).
+In this work, we propose a conditional generative model for likelihood-free inference based on a dynamic formulation of conditional optimal transport.
+Specifically, our contributions are as follows:
   1. 1.
-    In Section 4, we study the conditional Wasserstein space $\mathbb{P}_{p}^{\mu}(Y\times U)$ equipped with the conditional Wasserstein distance $W_{p}^{\mu}(Y\times U)$.
-We show that this space is a metric space, and moreover, admits constant speed geodesics between any two points.
+    We develop a general theoretical framework for dynamic conditional optimal transport in separable Hilbert spaces.
+Our framework is applicable in infinite-dimensional spaces, enabling applications in function space Bayesian inference.
+In Section 4, we study the conditional Wasserstein space $\mathbb{P}_{p}^{\mu}(Y\times U)$ and show that this space admits constant speed geodesics between any two measures.
+In Section 5, we characterize the absolutely continuous curves of measures in $\mathbb{P}_{p}^{\mu}(Y\times U)$ via the continuity equation and triangular vector fields.
+As a consequence, we obtain conditional generalizations of the McCann interpolants (McCann, 1997) and the Benamou-Brenier Theorem (Benamou and Brenier, 2000).
   2. 2.
-    In Section 5, we characterize the absolutely continuous curves of measures in $\mathbb{P}_{p}^{\mu}(Y\times U)$ via the continuity equation and triangular vector fields.
-As a corollary, we obtain a conditional version of the Benamou-Brenier theorem (Benamou and Brenier, 2000).
+    In Section 6, we propose COT flow matching (COT-FM), a simulation-free flow-based model for conditional generation.
+This model directly leverages our theoretical framework, where we learn to model a path of measures interpolating between an arbitrary source and target distribution via a geodesic in the conditional Wasserstein space.
   3. 3.
-    In Section 6 we propose a method for learning dynamic COT maps based on flow matching.
-In Section 7, we demonstrate our method on two image-to-image translation tasks, as well as an infinite-dimensional Bayesian inverse problem.
+    In Section 7, we demonstrate our method on several challenging conditional generation tasks.
+We apply our method to two Bayesian inverse problems – one arising from the Lotka-Volterra dynamical system, and an infinite-dimensional problem arising from the Darcy Flow PDE.
+Our method shows competitive performance against recent COT methods.
 2 Related Work
-Optimal Transport.
-Machine learning based approaches to unconditional optimal transport plans typically focus on the static problem (Makkuva et al., 2020; Korotin et al., 2022; Taghvaei and Jalali, 2019).
-Approaches to learning optimal transport maps in a dynamic fashion often require restricted architectures (Bunne et al., 2022b) or use likelihood-based approaches (Tong et al., 2020; Finlay et al., 2020; Onken et al., 2021; Huguet et al., 2022), which require simulating from the model during training.
-While most work focuses on the squared norm cost, the recent approaches of Pooladian et al.
-(2023b) and Neklyudov et al.
-(2023) study dynamical optimal transport with costs specified by a Lagrangian.
-These frameworks, however, typically have no clear interpretation as a static problem.
-Conditional optimal transport (COT), on the other hand, is relatively unexplored.
-Bunne et al.
-(2022a) and Wang et al.
-(2023) learn supervised COT maps in a static setting via convex networks (Makkuva et al., 2020).
-Hosseini et al.
-(2023) recently develop a framework for conditional optimal transport that is applicable in infinite-dimensional spaces.
-This builds on the work of Carlier et al.
-(2016) which applies COT for vector quantile regression.
-A common approach to COT is to employ triangular plans (Baptista et al., 2020).
-Our work develops a dynamic formulation of the static COT problem (Hosseini et al., 2023; Carlier et al., 2016) using a notion of triangular vector fields.
-Additionally, Wang et al.
-(2023) propose a likelihood-based approach using continuous normalizing flows, but do not develop the theory of dynamical COT as we do here.
+Conditional Optimal Transport.
+Conditional Optimal Transport (COT) remains relatively under-explored in both machine learning and related fields.
+Recent approaches learn static COT maps via input convex networks (Bunne et al., 2022; Wang et al., 2023) or normalizing flows (Wang et al., 2023).
+In addition, there have been a number of heuristic approaches to conditional simulation through W-GANs (Sajjadi et al., 2017; Adler and Öktem, 2018; Kim et al., 2022, 2023), for which Chemseddine et al.
+(2023) provide a rigorous basis.
+Closely related to our method are those which employ triangular plans (Carlier et al., 2016; Trigila and Tabak, 2016), which have been modeled through GANs in Euclidean spaces (Baptista et al., 2020) and function spaces (Hosseini et al., 2023).
+In contrast, our work uses a novel dynamic formulation of COT, which we model through a generalization of flow matching (Lipman et al., 2022; Albergo et al., 2023b).
+This allows us to use flexible architectures while avoiding the difficulties of training GANs (Arora et al., 2018).
 Simulation-Free Continuous Normalizing Flows.
-Flow matching (Lipman et al., 2022) and the equivalent stochastic interpolant framework (Albergo and Vanden-Eijnden, 2022; Albergo et al., 2023a) are a class of methods for building continuous-time normalizing flows in a simulation-free manner.
-These methods transform an arbitrary source distribution into a target distribution via a specified path of measures, which is modeled through a learned vector field.
-These methods have been extended to setting of Riemannian manifolds (Chen and Lipman, 2023) and infinite-dimensional spaces (Kerrigan et al., 2023).
-Notably, these works do not approximate an optimal transport between the source and target measures and instead use an independent coupling, although stochastic interpolants (Albergo and Vanden-Eijnden, 2022; Albergo et al., 2023b) may be rectified (Liu et al., 2022; Lee et al., 2023) through an iterative optimization procedure.
-Importantly, the paths in flow matching (Lipman et al., 2022) maybe taken to be conditionally optimal, but marginally they are not.
+Flow matching (Lipman et al., 2022) and stochastic interpolants (Albergo et al., 2023a) are a class of methods for building continuous-time normalizing flows in a simulation-free manner.
+Notably, these works do not approximate an optimal transport between the source and target measures.
 Pooladian et al.
-(2023a) and Tong et al.
-(2023) propose instead to couple the source and target distributions via optimal transport.
-While some works (Lipman et al., 2022; Davtyan et al., 2023; Gebhard et al., 2023) have studied flow matching approaches for conditional distributions, these approaches simply provide additional conditioning information to the model and do not study COT.
-Concurrent work (Isobe et al., 2024) proposes extended flow matching, which learns a matrix field to perform conditional generation.
-However, the results of Isobe et al.
-(2024) are limited to the finite-dimensional setting and do not develop the theory of COT.
-At the time of our submission, concurrent work by Barboni et al. (2024) and Chemseddine et al. (2024) appeared with similar results. Our results were developed independently from these works.
+(2023) and Tong et al.
+(2023) propose instead to couple the source and target distributions via optimal transport, leading to marginally optimal paths.
+In this work, we study an extension of these techniques for conditional generation.
+While some works (Davtyan et al., 2023; Gebhard et al., 2023; Isobe et al., 2024; Wildberger et al., 2024) have applied flow matching for conditional generation, these approaches do not employ COT.
+Notably, the aforementioned approaches are limited to the finite-dimensional setting, whereas our method adds to the growing literature on function-space generative models (Hosseini et al., 2023; Kerrigan et al., 2023, 2024; Lim et al., 2023; Franzese et al., 2024).
+At the time of our submission, concurrent work by Barboni et al.
+(2024) and Chemseddine et al.
+(2024) appeared with similar results.
+Our results were developed independently from these works.
 3 Background and Notation
-Let $X,X^{\prime},X^{\prime\prime}$ represent arbitrary separable Hilbert spaces, equipped with the Borel $\sigma$-algebra.
+Let $X,X^{\prime}$ represent arbitrary separable Hilbert spaces, equipped with the Borel $\sigma$-algebra.
 We use $\mathbb{P}(X)$ to represent the space of Borel probability measures on $X$, and $\mathbb{P}_{p}(X)\subseteq\mathbb{P}(X)$ to represent the subspace of measures having finite $p$th moment.
 If $\eta\in\mathbb{P}(X)$ is a probability measure on $X$ and $T:X\to X^{\prime}$ is measurable, then the pushforward measure $T_{\#}\eta({-})=\eta(T^{-1}({-}))$ is a probability measure on $X^{\prime}$.
-Maps of the form $\pi^{X}:X\times X^{\prime}\to X$ represent the canonical projection onto the respective space.
-We will also use maps such as $\pi^{1}:X\times X^{\prime}\times X^{\prime\prime}\to X$ and $\pi^{1,2}:X\times X^{\prime}\times X^{\prime\prime}\to X\times X^{\prime}$ to denote projection onto specific coordinates.
-The set $C_{b}(X)$ represents the space of continuous and bounded functions and $C^{\infty}_{c}(X)$ is the space of smooth and compactly supported functions.
-In our specific case, we assume that we have two separable Hilbert spaces of interest.
+Maps of the form e.g.
+$\pi^{X}:X\times X^{\prime}\to X$ represent the canonical projection.
+We assume that we have two separable Hilbert spaces of interest.
 The first, $Y$, is a space of observations, and the second, $U$, is a space of unknowns.
-These spaces may be of infinite dimensions, but a special case of practical interest is when $Y$ and $U$ are finite dimensional Euclidean spaces.
+These spaces may be of infinite dimensions, but a case of practical interest is when $Y$ and $U$ are finite dimensional Euclidean spaces.
 We will consider the product space $Y\times U$, equipped with the canonical inner product obtained via the sum of the inner products on $Y$ and $U$, under which the space $Y\times U$ is also a separable Hilbert space.
-Let $\eta\in\mathbb{P}(Y\times U)$ be a joint measure.
+Let $\eta\in\mathbb{P}(Y\times U)$ be a joint probability measure.
 The measures $\pi^{Y}_{\#}\eta\in\mathbb{P}(Y)$ and $\pi^{U}_{\#}\eta\in\mathbb{P}(U)$ obtained via projection are the marginals of $\eta$.
 We use $\eta^{y}\in\mathbb{P}(U)$ to represent the measure obtained by conditioning $\eta$ on the value $y\in Y$.
-By the disintegration theorem, such conditional measures exist and are essentially unique, in the sense that there exists a Borel set $E\subseteq Y$ with $\pi^{Y}_{\#}\eta(E)=0$, and $\eta^{y}$ are unique for $y\notin E$.
-3.1 Unconditional Optimal Transport
-We provide here a brief overview of optimal transport in the standard unconditional setting.
+By the disintegration theorem (Bogachev and Ruas, 2007, Chapter 10), such conditional measures exist and are essentially unique, in the sense that there exists a Borel set $E\subseteq Y$ with $\pi^{Y}_{\#}\eta(E)=0$, and the $\eta^{y}$ are unique for $y\notin E$.
+3.1 Static Conditional Optimal Transport
+In conditional optimal transport, we are given a target measure $\nu\in\mathbb{P}(Y\times U)$ and some source measure $\eta\in\mathbb{P}(U)$, and we seek a transport map $T:Y\times U\to U$ such that $T_{\#}(y,{-})_{\#}\eta=\nu^{y}$ for all $y\in Y$.
+If such a map were available, by drawing samples $u_{0}\sim\eta$ and transforming them, one would obtain samples $T(y,u)\sim\nu^{y}$.
+Solving this transport problem for each fixed $y$ is expensive at best, or impossible when only has a single (or no) samples $(y,u)\sim\nu$ for any given $y$.
+Thus, one must leverage information across different observations $y$.
+To that end, recent work has focused on the notion of triangular mappings $T:Y\times U\to Y\times U$ (Hosseini et al., 2023; Baptista et al., 2020) of the form $T(y,u)=\left(T_{Y}(y),T_{U}(T_{Y}(y),u)\right)$ for some $T_{Y}:Y\to Y$ and $T_{U}:Y\times U\to U$.
+Triangular mappings are of interest as they allow us to obtain conditional couplings from joint couplings.
+Proposition 1 (Theorem 2.4 (Baptista et al., 2020), Prop. 2.3 (Hosseini et al., 2023))
+Suppose $\eta,\nu\in\mathbb{P}(Y\times U)$ and $T:Y\times U\to Y\times U$ is triangular.
+If $T_{\#}\eta=\nu$, then $T_{U}(T_{Y}(y),{-})_{\#}\eta^{y}=\nu^{T_{Y}(y)}$ for $\pi^{Y}_{\#}\eta$-almost every $y$.
+In many scenarios of practical interest, the source measure $\eta$ and the target measure $\nu$ have the same $Y$-marginals.
+We will henceforth make this assumption, and use $\mu=\pi^{Y}_{\#}\eta=\pi^{Y}_{\#}\nu$ to represent this marginal.
+In this case, we may take $T_{Y}$ to be the identity mapping, so that the conclusion of Proposition 1 simplifies to $T_{U}(y,{-})_{\#}\eta^{y}=\nu^{y}$ for $\mu$-almost every $y$.
+We note that in situations where such an assumption does not hold, one may simply preprocess the source measure $\eta$ via an invertible mapping $T_{Y}$ satisfying $[T_{Y}]_{\#}[\pi^{Y}_{\#}\eta]=\pi^{Y}_{\#}\nu$ (Hosseini et al., 2023, Prop 3.2).
+Given a source and target measures $\eta,\nu\in\mathbb{P}^{\mu}(Y\times U)$ and a cost function $c:(Y\times U)^{2}\to\mathbb{R}$, the conditional Monge problem seeks to find a triangular mapping solving
+$\inf_{T}\left\{\int_{Y\times U}c(y,u,T(y,u))\,\mathrm{d}\eta(y,u)\mid T_{\#}\eta=\nu,T:(y,u)\mapsto(y,T_{U}(y,u))\right\}.$ (1)
+The conditional Monge problem also admits a relaxation under which one only considers couplings whose $Y$-components are almost surely equal.
+To that end, for $\eta,\nu\in\mathbb{P}_{p}^{\mu}(Y\times U)$ we define the set of triangular couplings $\Pi_{Y}(\eta,\nu)$ to be the couplings of $\eta$ and $\nu$ that almost surely fix the $Y$-components,
+$\Pi_{Y}(\eta,\nu)=\left\{\gamma\in\mathbb{P}\left(\left(Y\times U\right)^{2}\right)\mid\pi^{1,2}_{\#}\gamma=\eta,\pi^{3,4}_{\#}\gamma=\nu,\pi_{\#}^{1,3}=(I,I)_{\#}\mu\right\}.$ (2)
+In other words, a triangular coupling $\gamma\in\Pi_{Y}(\eta,\nu)$ has samples $(y_{0},u_{0},y_{1},u_{1})\sim\gamma$ such that $y_{0}=y_{1}$ almost surely.
+The conditional Kantorovich problem seeks a triangular coupling solving
+$\inf_{\gamma}\left\{\int_{(Y\times U)^{2}}c(y_{0},u_{0},y_{1},u_{1})\,\mathrm{d}\gamma(y_{0},u_{0},y_{1},u_{1})\mid\gamma\in\Pi_{Y}(\eta,\nu)\right\}.$ (3)
+Hosseini et al.
+(2023) prove the existence of minimizers to the conditional Kantorovich and Monge problems under very general assumptions.
+Moreover, optimal couplings to the conditional Kantorovich problem induce optimal couplings for $\mu$-almost every conditional measure.
+We refer to Appendix B and Hosseini et al.
+(2023) for further details.
+4 Conditional Wasserstein Space
+Motivated by our discussion on triangular transport maps, we introduce the conditional Wasserstein spaces, consisting of joint measures with finite $p$th moments and having fixed $Y$-marginals $\mu$.
+Interestingly, Gigli (2008, Chapter 4) studies the same space for the purposes of constructing geometric tangent spaces in the usual Wasserstein space.
+Definition 1 (Conditional Wasserstein Space)
+Suppose $\mu\in\mathbb{P}(Y)$ is given and $1\leq p<\infty$. The conditional $p$-Wasserstein space is
+$\mathbb{P}_{p}^{\mu}(Y\times U)=\left\{\gamma\in\mathbb{P}_{p}(Y\times U)\mid\pi^{Y}_{\#}\gamma=\mu\right\}.$ (4)
+We now equip $\mathbb{P}_{p}^{\mu}(Y\times U)$ with a metric $W_{p}^{\mu}$, the conditional Wasserstein distance.
+Intuitively, the conditional Wasserstein distance measures the usual Wasserstein distance between all of the conditional distributions in expectation under the fixed $Y$-marginal $\mu$.
+Definition 2 (Conditional $p$-Wasserstein Distance)
+Suppose $\eta,\nu\in\mathbb{P}_{p}^{\mu}(Y\times U)$ and $1\leq p<\infty$. The function $W_{p}^{\mu}:\mathbb{P}_{p}^{\mu}(Y\times U)\times\mathbb{P}_{p}^{\mu}(Y\times U)\to\mathbb{R}$,
+$W_{p}^{\mu}(\eta,\nu)=\left(\mathbb{E}_{y\sim\mu}\left[W_{p}^{p}(\eta^{y},\nu^{y})\right]\right)^{1/p}=\left(\int_{Y}W_{p}^{p}(\eta^{y},\nu^{y})\,\mathrm{d}\mu(y)\right)^{1/p}$ (5)
+is the conditional $p$-Wasserstein distance. $W_{p}$ is the usual Wasserstein distance for measures on $U$.
+By Jensen’s inequality we have $W_{p}^{\mu}(\eta,\nu)\geq\mathbb{E}_{y\sim\mu}\left[W_{p}(\eta^{y},\nu^{y})\right]$.
+In the following, we show that the conditional Wasserstein distance is a well-defined metric as well as a few other metric properties.
+Proposition 2 (Some Properties of $W_{p}^{\mu}$)
+Let $1\leq p<\infty$.
+  1. (a)
+    $W_{p}^{\mu}$ is well-defined, finite, and equals the minimal conditional Kantorovich cost.
+  2. (b)
+    $W_{p}^{\mu}$ is a metric on the space $\mathbb{P}_{p}^{\mu}(Y\times U$).
+  3. (c)
+    There does not exist $C>0$ such that $W_{p}^{\mu}(\eta,\nu)\leq CW_{p}(\eta,\nu)$ for all $\eta,\nu\in\mathbb{P}_{p}^{\mu}(Y\times U)$.
+  4. (d)
+    For all $\eta,\nu\in\mathbb{P}_{p}^{\mu}(Y\times U)$, $W_{p}\left(\pi^{U}_{\#}\eta,\pi^{U}_{\#}\nu\right)\leq W_{p}^{\mu}(\eta,\nu)$ and $W_{p}(\eta,\nu)\leq W_{p}^{\mu}(\eta,\nu)$.
+Proposition 2(c, d) together shows that one should expect the topology generated by $W_{p}^{\mu}$ to be stronger than the unconditional distance $W_{p}$.
+Here, we note that Gigli (2008) and Chemseddine et al.
+(2023) previously showed that $W_{p}^{\mu}$ is a metric through an equivalence with restricted couplings.
+Our approach builds on the results of Hosseini et al.
+(2023) and is somewhat more direct, and hence our proofs may be of independent interest.
+We include here an example where the conditional $2$-Wasserstein distance may be explicitly computed.
+Example: Gaussian Measures.
+Suppose $Y=U=\mathbb{R}$ and that $\eta,\nu\in\mathbb{P}_{2}^{\mu}(Y\times U)$ are Gaussians of the form
+$\eta=\mathcal{N}(0,I)\qquad\nu=\mathcal{N}\left(0,\begin{bmatrix}1&\rho\\ \rho&1\end{bmatrix}\right)\qquad|\rho|<1.$ (6)
+It follows that $\mu=\pi^{Y}_{\#}\eta=\pi^{Y}_{\#}\nu=\mathcal{N}(0,1)$.
+As $\eta^{y},\nu^{y}$ are Gaussians, their $W_{2}$ distance admits a closed form and we can directly compute $W_{2}^{\mu,2}(\eta,\nu)=2(1-\sqrt{1-\rho^{2}})$.
+This is zero if and only if $\rho=0$, i.e.
+$\eta=\nu$.
+However, $\pi^{U}_{\#}\eta=\pi^{U}_{\#}\nu=\mathcal{N}(0,1)$ and $W_{2}(\pi^{U}_{\#}\eta,\pi^{U}_{\#}\nu)=0$ regardless of $\rho$.
+Moreover, the unconditional distance is $W_{2}^{2}(\eta,\nu)=2\left(2-\sqrt{1-\rho}-\sqrt{1+\rho}\right)$, from which it is easy to verify that $W_{2}(\eta,\nu)\leq W_{2}^{\mu}(\eta,\nu)$.
+See Appendix C for arbitrary Gaussians.
+Conditional Wasserstein Space as a Geodesic Space.
+We now turn our attention to the geodesics in $\mathbb{P}_{p}^{\mu}(Y\times U)$.
+In particular, we show that there exists a constant speed geodesic between any two measures in $\mathbb{P}_{p}^{\mu}(Y\times U)$, generalizing a similar result in the unconditional setting (Santambrogio, 2015, Theorem 5.27).
+Moreover, we show that under suitable regularity assumptions, solutions to the conditional Monge problem (1) induce constant speed geodesics.
+Our motivation for studying geodesics in $\mathbb{P}_{p}^{\mu}(Y\times U)$ is practical – in Section 6, we show how one can model geodesics in $\mathbb{P}_{p}^{\mu}(Y\times U)$ in order to obtain a conditional flow-based model whose paths are easy to integrate.
+A curve is a continuous function $\gamma_{\bullet}:I\to\mathbb{P}_{p}^{\mu}(Y\times U)$ where $I=(a,b)\subseteq\mathbb{R}$ is any open interval of finite length.
+If $(\gamma_{t})$ is an absolutely continuous curve, then its metric derivative $|\gamma^{\prime}|(t)$ (Ambrosio et al., 2005, Chapter 1) exists for almost every $t\in(a,b)$.
+A curve $(\gamma_{t})$ is called a constant speed geodesic if for all $a<s\leq t<b$, we have $W_{p}^{\mu}(\gamma_{s},\gamma_{t})=|t-s|W_{p}^{\mu}(\gamma_{a},\gamma_{b})$.
+It is straightforward to show that every constant speed geodesic is absolutely continuous.
+Theorem 1 ($\mathbb{P}_{p}^{\mu}(Y\times U$) is a Geodesic Space)
+For any $\eta,\nu\in\mathbb{P}_{p}^{\mu}(Y\times U)$, there exists a constant speed geodesic between $\eta$ and $\nu$.
+When an optimal triangular coupling $\gamma^{\star}\in\Pi_{Y}(\eta,\nu)$ is induced by an injective triangular map $T^{\star}$, we may recover a constant speed geodesic in $\mathbb{P}_{p}^{\mu}(Y\times U)$, generalizing the McCann interpolant (McCann, 1997) to the conditional setting.
+We refer to Proposition 4 for sufficient conditions on $\eta,\nu$ under which such a $T^{\star}$ exists.
+Informally, samples from $(y_{0},u_{0})\sim\eta$ flow in a straight path at a constant speed to their destination $T^{\star}(y_{0},u_{0})$.
+Theorem 2 (Conditional McCann Interpolants)
+Fix $\eta,\nu\in\mathbb{P}_{p}^{\mu}(Y\times U)$.
+Suppose $T^{\star}(y,u)=(y,T^{\star}_{\mathcal{U}}(y,u))$ is an injective triangular map solving the conditional Monge problem (1).
+Define the maps $T_{t}:Y\times U\to Y\times U$ for $0\leq t\leq 1$ via $T_{t}=(1-t)I+tT^{\star}$, and define the curve of measures $\gamma_{t}=[T_{t}]_{\#}\eta\in\mathbb{P}_{p}^{\gamma}(Y\times U)$.
+Then,
+  1. (a)
+    $(\gamma_{t})$ is absolutely continuous and a constant speed geodesic between $\eta,\nu$
+  2. (b)
+    The vector field $v_{t}(T_{t}^{\star}(y,u))=(0,T_{U}^{\star}(y,u)-u)$ generates the path $\gamma_{t}$, in the sense that $(\gamma_{t},v_{t})$ solve the continuity equation (7).
+5 Conditional Benamou-Brenier Theorem
+In this section, we prove a characterization of the absolutely continuous curves in $\mathbb{P}_{p}^{\mu}(Y\times U)$.
+As a corollary, we obtain a conditional generalization of the Benamou-Brenier Theorem (Benamou and Brenier, 2000), giving us a dynamical characterization of the conditional Wasserstein distance.
+Roughly speaking, all such curves are generated by a vector field on $Y\times U$ which has zero velocity in the $Y$ component.
+This is natural, as all measures in $\mathbb{P}_{p}^{\mu}(Y\times U)$ have a fixed $Y$-marginal $\mu$.
+Such a vector field can be informally seen as tangent to a curve of measures, and is the dynamic analogue of the triangular maps discussed in Section 3.
+More formally, given an open interval $I\subseteq\mathbb{R}$, a time-dependent Borel vector field $v:I\times Y\times U\to Y\times U$ is said to be triangular if there exists a Borel vector field $v^{U}:I\times Y\times U\to U$ such that $v_{t}(y,u)=\left(0,v_{t}^{U}(y,u)\right)$.
+Continuity Equation.
+We introduce some necessary background which allows us to link vector fields to curves of measures.
+The continuity equation $\partial_{t}\gamma_{t}+\text{div}(v_{t}\gamma_{t})=0$ describes the evolution of a measure $\gamma_{t}$ which flows along a given vector field $v_{t}$ (Ambrosio et al., 2005, Chapter 8).
+This equation must be understood distributionally, i.e.
+for every $\varphi$ in an appropriate space of test functions,
+$\int_{I}\int_{Y\times U}\left(\partial_{t}\varphi(y,u,t)+\langle v_{t}(y,u),\nabla_{y,u}\varphi(y,u,t)\rangle\right)\,\mathrm{d}\gamma_{t}(y,u)\,\mathrm{d}t=0.$ (7)
+We consider cylindrical test functions $\varphi\in\text{Cyl}(Y\times U\times I)$, i.e.
+of the form $\varphi(y,u,t)=\psi(\pi^{d}(y,u),t)$ where $\pi^{d}:Y\times U\to\mathbb{R}^{d}$ maps $(y,u)\to(\langle(y,u),e_{1}\rangle,\dots,\langle(y,u),e_{d}\rangle)$ where $\{e_{1},e_{2},\dots,e_{d}\}$ is any orthonormal family in $Y\times U$.
+In the finite dimensional setting, one may take $\varphi\in C_{c}^{\infty}(Y\times U)$ to be smooth and compactly supported (Ambrosio et al., 2005, Remark 8.1.1).
+In Appendix E, we prove Lemma 1, which is key in proving Theorem 4 below.
+Informally, Lemma 1 states that if (7) is satisfied for a joint distribution and triangular vector field, then the continuity equation is also satisfied for the corresponding conditional distributions and $U$ components of the vector field.
+Lemma 1 (Triangular Vector Fields Preserve Conditionals)
+Suppose $v_{t}(y,u)=(0,v_{t}^{U}(y,u))$ is triangular and that $(\gamma_{t})\subset\mathbb{P}_{p}^{\mu}(Y\times U)$ is a path of measures such that $(v_{t},\gamma_{t})$ satisfy the continuity equation in the distributional sense.
+Then, it follows that for $\mu$-almost every $y\in Y$, we have $\partial_{t}\gamma_{t}^{y}+\nabla\cdot(v_{t}^{U}(y,{-})\gamma_{t}^{y})=0$.
+Absolutely Continuous Curves.
+In this section, we state our characterization of absolutely continuous curves in $\mathbb{P}_{p}^{\mu}(Y\times U)$.
+Informally, given such a curve, Theorem 3 provides us with a triangular vector field which generates the curve, in the sense that the pair solve the continuity equation.
+Theorem 3 (Absolutely Continuous Curves in $\mathbb{P}_{p}^{\mu}(Y\times U)$)
+Let $I\subset\mathbb{R}$ be an open interval, and suppose $\gamma_{t}:I\to\mathbb{P}_{p}^{\mu}(Y\times U)$ is an absolutely continuous in the $W_{p}^{\mu}$ metric with $|\gamma^{\prime}|(t)\in L^{1}(I)$.
+Then, there exists a Borel vector field $v_{t}(y,u)$ such that
+  1. (a)
+    $v_{t}$ is triangular
+  2. (b)
+    $v_{t}\in L^{p}(\gamma_{t},Y\times U)$ and $\left\lVert v_{t}\right\rVert_{L^{p}(\gamma_{t},Y\times U)}\leq|\gamma^{\prime}|(t)$ for a.e. $t$
+  3. (c)
+    $(v_{t},\gamma_{t})$ solve the continuity equation distributionally.
+Conversely, we show in Theorem 4 that if the pair $(\gamma_{t},v_{t})$ solve the continuity equation and $v_{t}$ is triangular, then the curve $(\gamma_{t})$ is absolutely continuous and $|\gamma^{\prime}(t)|\leq\left\lVert v_{t}\right\rVert_{L^{p}(\gamma_{t},Y\times U)}$.
+The main technique of this result is to study the collection of conditional continuity equations (which is feasible by Lemma 1) and to apply the converse of Ambrosio et al.
+(2005, Theorem 8.3.1).
+In this setting, the infinite-dimensional result is obtained via a finite-dimensional approximation argument.
+Theorem 4 (Continuous Curves Generated by Triangular Vector Fields)
+Suppose that $\gamma_{t}:I\to\mathbb{P}_{p}^{\mu}(Y\times U)$ is narrowly continuous and $(v_{t})$ is a triangular vector field such that $(\gamma_{t},v_{t})$ solve the continuity equation with $\left\lVert v_{t}\right\rVert_{L^{p}(\gamma_{t},Y\times U)}\in L^{1}(I)$.
+Then, $\gamma_{t}:I\to\mathbb{P}_{p}^{\mu}(Y\times U)$ is absolutely continuous in the $W_{p}^{\mu}$ metric and $|\gamma^{\prime}|(t)\leq\left\lVert v_{t}\right\rVert_{L^{p}(\mu_{,}Y\times U)}$ for almost every $t$.
+As a corollary of Theorem 3 and Theorem 4, we obtain a conditional version of the Benamou-Brenier theorem (Benamou and Brenier, 2000).
+The proof of Theorem 5 largely follows the unconditional case (see e.g.
+Ambrosio et al.
+(2005, Chapter 8)), but we include it for the sake of completeness.
+Theorem 5 (Conditional Benamou-Brenier)
+Let $1<p<\infty$. For any $\eta,\nu\in\mathbb{P}_{p}^{\mu}(Y\times U)$, we have
+[…]
+6 COT Flow Matching
+We have thus far seen that the COT problem (3) admits a dynamical formulation by Theorem 5, where one may take the underlying vector fields to be triangular.
+We use these results to design a principled model for conditional generation based on flow matching (Lipman et al., 2022; Albergo et al., 2023b; Tong et al., 2023).
+We hereafter use the squared-distance cost (i.e.
+$p=2$).
+Figure 1: Samples from the ground-truth joint target distribution and the various models.
+Samples from COT-FM more closely match the ground-truth distribution than the baselines.
+In the final column, we plot conditional KDEs for samples drawn conditioned on the $y$ value indicated by the dashed horizontal line.
+See Appendix F for a larger figure and additional results.
+Flow Matching.
+We assume that we have access to samples $z_{0}=(y_{0},u_{0})\sim\eta(y_{0},u_{0})\in{\mathbb{P}_{p}^{\mu}(Y\times U)}$ from a source measure, and samples $z_{1}=(y_{1},u_{1})\sim\nu(y_{1},u_{1})\in\mathbb{P}_{p}^{\mu}(Y\times U)$ from a target measure.
+Let $z=(z_{0},z_{1})\sim\rho(z_{0},z_{1})\in\Pi(\eta,\nu)$ be any coupling of the source and target measure.
+We specify a collection of measures and vector fields on $Y\times U$ via
+$\gamma_{t}(y,u\mid z)=\mathcal{N}\left(y,u\mid tz_{1}+\left(1-t\right)z_{0},C\right)\qquad v_{t}(y,u\mid z)=z_{1}-z_{0}$ (8)
+where $C$ is any trace-class covariance operator (Da Prato and Zabczyk, 2014).
+As is standard in flow matching (Lipman et al., 2022; Kerrigan et al., 2024), we obtain from Equations (8) a marginal measure $\gamma_{t}(y,u)$ and vector field $v_{t}(y,u)$ satisfying the continuity equation via
+$\gamma_{t}(y,u)=\int_{(Y\times U)^{2}}\gamma_{t}(y,u\mid z)\,\mathrm{d}\rho(z)\qquad v_{t}(y,u)=\int_{(Y\times U)^{2}}v_{t}(y,u\mid z)\frac{\,\mathrm{d}\gamma_{t}(y,u\mid z)}{\,\mathrm{d}\gamma_{t}(y,u)}\,\mathrm{d}\rho(z).$ (9)
+This marginal path $(\gamma_{t})_{t=0}^{1}$ interpolates between the source measure ($t=0$) and a smoothed version of the target measure ($t=1$).
+To transform source samples from $\eta$ into target samples from $\nu$, we seek to learn the intractable vector field $v_{t}(y,u)$ with a model $v^{\theta}(t,y,u)$ by minimizing the loss [^1] Previous work has referred to this as the conditional flow matching loss (Tong et al., 2023), which is not to be confused with the notion of conditioning that we focus on in this work.
+$\mathcal{L}(\theta)=\mathbb{E}_{t,\rho(z),\gamma_{t}(y,u\mid z)}\left\lVert v^{\theta}(t,y,u)-v_{t}(y,u\mid z)\right\rVert^{2}$ (10)
+which has the same $\theta$-gradient as the MSE loss to the true vector field $u_{t}(y,u)$ (Tong et al., 2023).
+COT Flow Matching.
+In the preceding section, $\rho(z)$ may be an arbitrary coupling between $\eta$ and $\nu$.
+Motivated by Proposition 1, we will choose $\rho$ to be a COT coupling.
+Under sufficient regularity conditions (see Appendix B), this COT plan will be induced by a triangular map.
+In turn, Theorem 2 gives us that this triangular map is generated by a triangular vector field of the form (8).
+Thus, we parametrize our model $u^{\theta}$ to also be triangular.
+Moreover, we recover the optimal dynamic transport given in Theorem 5 as $\text{Tr}(C)\to 0$ by a pointwise application of (Tong et al., 2023, Proposition 3.4).
+Given a collection of samples $\{z_{0}^{i},z_{1}^{i}\}_{i=1}^{n}$ drawn from $\eta$ and $\nu$, we approximate a conditional optimal coupling $\rho$ using standard numerical techniques with the cost function $c_{\epsilon}(y_{0},u_{0},y_{1},u_{1})=|y_{1}-y_{0}|^{2}+\epsilon|u_{1}-u_{0}|^{2}$ for some $0<\epsilon\ll 1$.
+Intuitively, such a cost penalizes mass transfer along the $Y$ dimension, which is precisely the constraint sought in the COT problem (3).
+Hosseini et al.
+(2023, Prop.
+3.11) show that as $\epsilon\downarrow 0$, we recover the true optimal triangular map.
+The COT coupling can either be precomputed for small datasets or computed on each minibatch drawn during training.
+After training, we obtain a learned triangular vector field $v^{\theta}(t,y,u)$.
+Given an arbitrary fixed $y\in Y$, we may approximately sample from the target $\nu(u\mid y)$ by sampling $u_{0}\sim\eta(u_{0}\mid y)$ and numerically solving the corresponding flow equation $\partial_{t}(y,u_{t})=v^{\theta}(t,y,u_{t})$ with initial condition $(y,u_{0})$.
+Source Measure.
+Our framework is agnostic to the choice of source measure $\eta$, allowing for flexibility in the modeling process.
+The main requirement is that the $Y$-marginals of the source $\eta$ and target $\eta$ must match.
+In some scenarios, this is trivially satisfied.
+If one is interested in using a source distribution which is simply random noise, one may take $\eta(y_{0},u_{0})=\pi^{Y}_{\#}\nu(y_{0})\otimes\eta_{U}(u_{0})$ to be the product of two independent distributions where $\eta_{U}$ is arbitrary, e.g.
+Gaussian noise.
+7 Experiments
+Table 1: Distances between the ground-truth and generated joint distributions for the 2D datasets.
+Our method (COT-FM) obtains lower distances than the considered baselines.
+Average results $\pm$ one standard deviation are reported across five test sets, with the lowest average distance in bold.
+Checkerboard Moons Circles Swissroll
+$W_{2}$ (1e-2) MMD (1e-3) $W_{2}$ (1e-2) MMD (1e-3) $W_{2}$ (1e-2) MMD (1e-3) $W_{2}$ (1e-2) MMD (1e-3)
+PCP-Map $6.27{\scriptstyle\pm 0.81}$ $0.21{\scriptstyle\pm 0.13}$ $8.44{\scriptstyle\pm 1.09}$ $0.22{\scriptstyle\pm 0.10}$ $6.19{\scriptstyle\pm 0.43}$ $\mathbf{0.20}{\scriptstyle\pm 0.17}$ $5.35{\scriptstyle\pm 0.93}$ $0.16{\scriptstyle\pm 0.13}$
+COT-Flow $8.20{\scriptstyle\pm 0.49}$ $0.26{\scriptstyle\pm 0.16}$ $18.49{\scriptstyle\pm 2.22}$ $1.32{\scriptstyle\pm 0.79}$ $10.04{\scriptstyle\pm 1.69}$ $0.24{\scriptstyle\pm 0.22}$ $6.47{\scriptstyle\pm 0.69}$ $0.19{\scriptstyle\pm 0.19}$
+FM $8.81{\scriptstyle\pm 0.58}$ $0.24{\scriptstyle\pm 0.20}$ $15.55{\scriptstyle\pm 0.77}$ $1.85{\scriptstyle\pm 0.22}$ $7.03{\scriptstyle\pm 0.17}$ $0.45{\scriptstyle\pm 0.11}$ $8.18{\scriptstyle\pm 0.34}$ $0.58{\scriptstyle\pm 0.09}$
+COT-FM (Ours) $\mathbf{4.69}{\scriptstyle\pm 1.00}$ $\mathbf{0.17}{\scriptstyle\pm 0.13}$ $\mathbf{6.50}{\scriptstyle\pm 1.41}$ $\mathbf{0.13}{\scriptstyle\pm 0.10}$ $\mathbf{5.56}{\scriptstyle\pm 0.43}$ $\mathbf{0.20}{\scriptstyle\pm 0.04}$ $\mathbf{4.64}{\scriptstyle\pm 1.26}$ $\mathbf{0.15}{\scriptstyle\pm 0.19}$
+Table 2: Statistical distances between MCMC and posterior samples $u\sim{\nu(u\mid y)}$ for each method on the LV dataset. Average results $\pm$ one standard deviation reported across five test sets.
+$W_{2}$ (1e-2) MMD (1e-3)
+PCP-Map $5.04{\scriptstyle\pm 0.05}$ $2.67{\scriptstyle\pm 2.1}$
+COT-Flow $4.86{\scriptstyle\pm 1.1}$ $\mathbf{0.83}{\scriptstyle\pm 0.50}$
+FM $11.41{\scriptstyle\pm 0.26}$ $2.65{\scriptstyle\pm 0.14}$
+COT-FM (Ours) $\mathbf{4.02}{\scriptstyle\pm 0.06}$ $0.95{\scriptstyle\pm 0.03}$
+We now illustrate our methodology (COT-FM) on a variety of conditional simulation tasks.
+We compare our method against several competitive baselines, namely PCP-Map (Wang et al., 2023), COT-Flow (Wang et al., 2023), and WaMGAN (Hosseini et al., 2023).
+These baselines are chosen as they reflect current state-of-the-art approaches to learning COT maps.
+We additionally compare against flow matching (Lipman et al., 2022; Wildberger et al., 2024) without COT, i.e.
+where the coupling between the source and target measures is the independent coupling $\rho(z_{0},z_{1})=\eta\otimes\nu$.
+Overall, our method (COT-FM) typically outperforms these baselines across the diverse and challenging set of tasks we consider.
+We find that PCP-Map (Wang et al., 2023) is a strong baseline, but we emphasize that this model relies on the use of an input-convex neural network (Amos et al., 2017) and it is hence unclear how to adapt this method to e.g.
+images.
+Appendix F contains further details and results for all of our experiments.
+[^2] Code for all of our experiments is available at https://github.com/GavinKerrigan/cot_fm.
+2D Synthetic Data.
+We first consider synthetic distributions where $Y=U=\mathbb{R}$.
+Our source measure is taken to be the independent product $\eta(y,u)=\pi^{Y}_{\#}\nu\otimes\mathcal{N}(0,1)$.
+We plot ground-truth joint distributions and samples for two datasets in Figure 1.
+See Appendix F for additional results.
+Samples from our method (COT-FM) closely match those from the ground-truth distribution, whereas samples from PCP-Map and COT-Flow (Wang et al., 2023) can produce samples in regions of zero support under the ground-truth distribution.
+In Table 1, we provide a quantitative analysis, where we measure the $W_{2}$ and MMD distances between the generated and ground-truth joint distributions.
+This is motivated by Proposition 1, as triangular maps which couple the joint distributions necessarily couple the conditional distributions.
+Our method outperforms the baselines across all metrics.
+Lotka-Volterra (LV) Dynamical System.
+Here we estimate parameters of the LV model given only noisy observations of its solution.
+The LV model has parameters $u=(\alpha,\beta,\gamma,\delta)\in\mathbb{R}^{4}_{\geq 0}$ and a pair of coupled nonlinear ODEs of the form
+$\frac{\,\mathrm{d}p_{1}(t)}{\,\mathrm{d}t}=\alpha p_{1}-\beta p_{1}p_{2}\quad\frac{\,\mathrm{d}p_{2}(t)}{\,\mathrm{d}t}=-\gamma p_{2}+\delta p_{1}p_{2}$ (11)
+whose solution $p(t)=(p_{1}(t),p_{2}(t))\in\mathbb{R}^{2}_{\geq 0}$ represents the number of prey and predator species at time $t\in[0,T]$.
+Following Alfonso et al.
+(2023), we assume $p(0)=(30,1)$ and that $\log(u)\sim\mathcal{N}(m,0.5I)$ with $m=(-0.125,-3,-0.125,-3)$.
+Given parameters $u\in\mathbb{R}^{4}_{\geq 0}$, we simulate Equation (11) for $t\in\{0,2,\dots,20\}$ to obtain a solution $z(u)\in\mathbb{R}^{22}_{\geq 0}$.
+An observation $y\in\mathbb{R}^{22}_{\geq 0}$ is obtained by the addition of log-normal noise, i.e.
+$\log(y)\sim\mathcal{N}(\log(z(u),0.1I)$.
+We thus may simulate many $(y,u)$ pairs from the target measure for training.
+Figure 2: Sample KDEs on the Lotka-Volterra inverse problem. The red lines denote the true parameter values.
+As a benchmark, we follow the settings of Alfonso et al.
+(2023) and choose parameters $u=(0.83,0.041,1.08,0.04)$ to generate a single observation $y$ as described above.
+In Figure 2, we plot a histogram of $10,000$ samples from the posterior $\nu(u\mid y)$ of COT-FM.
+Since the ground-truth posterior is intractable, we compare against differential evolution Metropolis MCMC (Braak, 2006).
+Samples from our method qualitatively resemble those from MCMC, and the posterior mode is typically close to the true unknown $u$ (shown in red).
+Our method is quantitatively closest to the MCMC samples in the $W_{2}$ metric, and competitive in the MMD metric (Table 2).
+Darcy Flow Inverse Problem.
+Here we consider an infinite-dimensional Bayesian inverse problem from the 2D Darcy flow PDE.
+The setting is adapted from Hosseini et al.
+(2023).
+We opt to compare against WaMGAN (Hosseini et al., 2023), as this is currently the only other extant amortized function-space COT method, and FFM (Kerrigan et al., 2023) as a function-space flow matching ablation.
+Figure 3: Darcy flow illustration.
+A true permeability $u$ is shown, as well as the pressure field $\rho$ and its observed, noisy version $y$.
+We compare an ensemble average of posterior samples from the various methods against MCMC (pCN) (Cotter et al., 2013).
+COT-FM achieves the lowest MSE to pCN.
+The Darcy flow PDE is an elliptic equation on a smooth domain $\Omega\subseteq\mathbb{R}^{d}$ which relates a permeability field $\exp(u)$, a pressure field $\rho$, and a source term $f$ via $-\text{div}\exp(u)\nabla\rho=f$ on $\Omega$ subject to $\rho=0$ on $\partial\Omega$.
+Our goal is to recover the permeability $u$ from noisy measurements $y$ of the pressure $\rho$.
+Both the unknown $u$ and observations $y$ are functions and thus infinite-dimensional.
+To define our target measure, we specify a prior $\nu(u)=\mathcal{N}(0,C)$ with a Matérn kernel $C$ of lengthscale $\ell=1/2$ and $\nu=3/2$.
+Given $u\sim\eta(u)$, the Darcy flow PDE is solved numerically (Alnæs et al., 2015) to obtain a solution $\rho(u)$ observed at some finite but arbitrary number of points $\{x_{1},\dots,x_{n}\}\subset\mathbb{R}^{2}$.
+An observation $y(u)$ is obtained by adding Gaussian noise to each observation, i.e.
+$y(u)\sim\mathcal{N}(\rho(u),\sigma^{2}I)$ where $\sigma=2.5\times 10^{-2}$.
+We implement all models via a Fourier Neural Operator (Li et al., 2020), allowing us to work with arbitrary discretizations, as required by the functional nature of this problem.
+Table 3: Predictive performance of the generated samples on the Darcy flow inverse problem. Average result $\pm$ one standard deviation obtained on 5 test sets of 5,000 samples each.
+MSE (1e-2) CRPS (1e-2)
+WaMGAN $6.55{\scriptstyle\pm 0.07}$ $18.75{\scriptstyle\pm 0.10}$
+FFM $7.30{\scriptstyle\pm 0.07}$ $\mathbf{15.47}{\scriptstyle\pm 0.06}$
+COT-FFM (Ours) $\mathbf{5.40}{\scriptstyle\pm 0.08}$ $15.56{\scriptstyle\pm 0.08}$
+We provide an illustration in Figure 3.
+As the true posterior is intractable, we compare against preconditioned Crank-Nicolson (pCN) (Cotter et al., 2013), a function-space MCMC method.
+In Figure 3, we plot the mean posteriors obtained from the various methods.
+Qualitatively, both COT-FM and FFM are good approximations to pCN, while WaMGAN has visual artifacts.
+However, the MSE between our method and the pCN mean is lower than that of FFM.
+Table 3 provides a quantitative comparison between the methods on a test set of 5,000 samples, where we measure MSE and CRPS (Hersbach, 2000).
+We compare the ensemble mean of 10 samples against the true $u$ value as running pCN for each observation is prohibitively expensive.
+COT-FM outperforms FFM and WaMGAN in terms of MSE and is on-par with FFM in terms of CRPS.
+See Appendix F for further results.
+8 Conclusion
+We analyze conditional optimal transport from a geometric and dynamical point of view.
+Our analysis culminates in the characterization of absolutely continuous curves of measures in a conditional Wasserstein space, resulting in a conditional analog of the Benamou-Brenier Theorem.
+We use these result to build on the framework of triangular transport and flow matching to develop simulation-free methods for conditional generative models.
+Our methods are applicable across a wide class of problems, and we demonstrate our methodology on several challenging inverse problems.
+Limitations and Broader Impacts.
+A limitation COT-FM is that computing the full COT plan can be expensive for large datasets, necessitating the use of minibatch approximations potentially resulting in sub-optimal plans.
+Moreover, computing the COT plan incurs a small additional computational cost compared to standard flow matching.
+As with all generative models, a potential negative impact is the potential for disinformation through generated samples being purported as real.
+Acknowledgments and Disclosure of Funding
+This research was supported by the Hasso Plattner Institute (HPI) Research Center in Machine Learning and Data Science at the University of California, Irvine, by the National Science Foundation under award 1900644, and by the National Institutes of Health under award R01-LM013344.
+Appendix: Table of Contents
+[…]
+[…]
+Appendix A Unconditional Optimal Transport
+We provide here a brief and informal overview of optimal transport in the standard unconditional setting.
 For more details, we refer to the standard references of Villani et al.
@@
 (2005).
-Fix a cost function $c:X\times X\to\mathbb{R}\cup\{+\infty\}$.
+Let $X$ be a separable metric space and fix a cost function $c:X\times X\to\mathbb{R}\cup\{+\infty\}$.
 Suppose we have two Borel measures $\eta,\nu\in\mathbb{P}(X)$.
@@
 $W_{p}(\eta,\nu)=\min_{(\gamma_{t},v_{t})}\left\{\int_{0}^{1}\int_{X}|v_{t}(x)|^{p}\,\mathrm{d}\gamma_{t}(x)\,\mathrm{d}t\mid\gamma_{0}=\eta,\gamma_{1}=\nu,\partial_{t}\gamma_{t}+\text{div}(v_{t}\gamma_{t})=0\right\}.$ (15)
-Here, we constrain our minimization problem over the set of measures and vector fields $(\gamma_{t},v_{t})$ interpolating between $\eta$ and $\nu$, satisfying a continuity equation (see Section 5.1).
+Here, we constrain our minimization problem over the set of measures and vector fields $(\gamma_{t},v_{t})$ interpolating between $\eta$ and $\nu$, satisfying a continuity equation (see Section 5).
 In Section 4, we study a generalization of the Wasserstein distances for conditional optimal transport problems.
 In particular, Theorem 5 provides a generalization of the Benamou-Brenier theorem to the conditional setting which recovers a conditional Wasserstein distance.
-3.2 Static Conditional Optimal Transport
-In the unconditional setting, one seeks an optimal coupling between two given measures.
-In contrast, in the conditional setting, we would like to find a family of couplings that can transport a given source measure into any conditional measure of the target distribution.
-In other words, given a target measure $\nu\in\mathbb{P}(Y\times U)$ and some source measure $\eta\in\mathbb{P}(U)$, one seeks a transport map $T:Y\times U\to U$ such that
-$T_{\#}(y,{-})_{\#}\eta=\nu^{y}\qquad\forall y\in Y.$ (5)
-If such a map were available, by drawing samples $u_{0}\sim\eta$ and transforming them, one would obtain samples $T(y,u)\sim\nu^{y}$.
-Solving this transport problem for each fixed $y$ is expensive at best, or impossible when only has a single (or no) samples $(y,u)\sim\nu$ for any given $y$.
-Thus, one must leverage information across different observations $y$.
-To that end, recent work has focused on the notion of triangular mappings $T:Y\times U\to Y\times U$ (Hosseini et al., 2023; Baptista et al., 2020) (Bogachev and Ruas, 2007, Section 10.10) of the form
-$T(y,u)=\left(T_{Y}(y),T_{U}(T_{Y}(y),u)\right)$ (6)
-for some $T_{Y}:Y\to Y$ and $T_{U}:Y\times U\to U$. Triangular mappings are of interest as they allow us to obtain conditional couplings from joint couplings.
-Proposition 1 (Theorem 2.4 (Baptista et al., 2020), Prop. 2.3 (Hosseini et al., 2023)).
-Suppose $\eta,\nu\in\mathbb{P}(Y\times U)$ and $T:Y\times U\to Y\times U$ is triangular.
-If $T_{\#}\eta=\nu$, then $T_{U}(T_{Y}(y),{-})_{\#}\eta^{y}=\nu^{T_{Y}(y)}$ for $\pi^{Y}_{\#}\eta$-almost every $y$.
-In many scenarios of practical interest, the source measure $\eta$ and the target measure $\nu$ have the same $Y$-marginals.
-We will henceforth make this assumption, and use $\mu=\pi^{Y}_{\#}\eta=\pi^{Y}_{\#}\nu$ to represent this marginal.
-In this case, we may take $T_{Y}$ to be the identity mapping, so that the conclusion of Proposition 1 simplifies to $T_{U}(y,{-})_{\#}\eta^{y}=\nu^{y}$ for $\mu$-almost every $y$.
-We note that in situations where such an assumption does not hold, one may simply preprocess the source measure $\eta$ via an invertible mapping $T_{Y}$ satisfying $[T_{Y}]_{\#}[\pi^{Y}_{\#}\eta]=\pi^{Y}_{\#}\nu$ (Hosseini et al., 2023, Prop 3.2).
-Definition 2 (Conditional Wasserstein Space).
-Suppose $\mu\in\mathbb{P}(Y)$ is a given probability measure over $Y$ and $1\leq p<\infty$. The set
-$\mathbb{P}^{\mu}(Y\times U)=\left\{\gamma\in\mathbb{P}_{p}(Y\times U)\mid\pi^{Y}_{\#}\gamma=\mu\right\}$ (7)
-is the space of joint measures on $Y\times U$ having fixed $Y$-marginals $\mu$.
-The set $\mathbb{P}_{p}^{\mu}(Y\times U):=\mathbb{P}^{\mu}(Y\times U)\cap\mathbb{P}_{p}(Y\times U)$ is the conditional $p$-Wasserstein space, consisting of joint measures with finite $p$th moments and $Y$-marginals $\mu$.
-Given a source and target measures $\eta,\nu\in\mathbb{P}^{\mu}(Y\times U)$, the conditional Monge problem seeks to find a triangular mapping solving
+Appendix B Conditional Optimal Transport
+This section contains additional discussion regarding the static COT problem, supplementing Section 3.1.
+We refer to Hosseini et al.
+(2023), Baptista et al.
+(2020), and Chemseddine et al.
+(2023) for further results and details.
+Given a source and target measures $\eta,\nu\in\mathbb{P}_{p}^{\mu}(Y\times U)$, the conditional Monge problem seeks to find a triangular mapping solving
 $\inf_{T}\left\{\int_{Y\times U}c(y,u,T(y,u))\,\mathrm{d}\eta(y,u)\mid T_{\#}\eta=\nu,T:(y,u)\mapsto(y,T_{U}(y,u))\right\}.$ (16)
@@
 In other words, if $\gamma\in\mathcal{R}_{Y}$, then samples $(y_{0},u_{0},y_{1},u_{1})\sim\gamma$ have $y_{0}=y_{1}$ almost surely.
-In addition, for any $\eta,\nu\in\mathbb{P}(Y\times U)$, we define the set of $(Y)$-restricted couplings $\Pi_{Y}(\eta,\nu)$ to be the probability measures in $\mathcal{R}_{Y}$ whose marginals are $\eta$ and $\nu$, i.e.
+In addition, for any $\eta,\nu\in\mathbb{P}(Y\times U)$, we define the set of triangular couplings $\Pi_{Y}(\eta,\nu)$ to be the probability measures in $\mathcal{R}_{Y}$ whose marginals are $\eta$ and $\nu$, i.e.
 $\Pi_{Y}(\eta,\nu)=\left\{\gamma\in\mathcal{R}_{Y}\mid\pi^{1,2}_{\#}\gamma=\eta,\pi^{3,4}_{\#}\gamma=\nu\right\}.$ (18)
@@
 Assuming sufficient regularity assumptions on the conditional measures, unique solutions to the conditional Monge problem exist.
-Proposition 3 (Prop 3.3 (Hosseini et al., 2023)).
+We restate these results here for the sake of completeness.
+Proposition 3 (Prop 3.3 (Hosseini et al., 2023))
 Fix $\eta,\nu\in\mathbb{P}^{\mu}(Y\times U)$.
@@
 Moreover, $\gamma^{\star,y_{0}}(y_{1},u_{0},u_{1})=\hat{\gamma}^{\star,y_{0}}(u_{0},u_{1})\delta(y_{1}-y_{0})$ where for $\mu$-almost every $y$ the measure $\gamma^{\star,y}$ is an optimal coupling for $\eta^{y},\nu^{y}$ under the cost $c^{y}(u_{0},u_{1})=c(y,u_{0},y,u_{1})$
-Proposition 4 (Prop 3.8 (Hosseini et al., 2023)).
+Proposition 4 (Prop 3.8 (Hosseini et al., 2023))
 Fix $1<p<\infty$ and $\eta,\nu\in\mathbb{P}_{p}^{\mu}(Y\times U)$.
@@
 If $\nu^{y}$ also assign zero measure to Gaussian null sets for $\mu$-almost every $y$, then $T^{\star}$ is injective $\eta$-almost everywhere.
-4 Conditional Wasserstein Space
-We now introduce a metric on the space $\mathbb{P}_{p}^{\mu}(Y\times U)$.
-Intuitively, the conditional Wasserstein distance measures the usual Wasserstein distance between all of the conditional distributions in expectation under the fixed $Y$-marginal $\mu$.
-Definition 5 (Conditional $p$-Wasserstein Distance).
-Suppose $\eta,\nu\in\mathbb{P}_{p}^{\mu}(Y\times U)$ and $1\leq p<\infty$. The function
-$W_{p}^{\mu}(\eta,\nu)=\left(\mathbb{E}_{y\sim\mu}\left[W_{p}^{p}(\eta^{y},\nu^{y})\right]\right)^{1/p}=\left(\int_{Y}W_{p}^{p}(\eta^{y},\nu^{y})\,\mathrm{d}\mu(y)\right)^{1/p}$ (12)
-is the conditional $p$-Wasserstein distance. Here, $W_{p}^{p}:\mathbb{P}_{p}(U)\times\mathbb{P}_{p}(U)\to\mathbb{R}$ is the usual $p$-Wasserstein distance for measures on $U$.
-By an application of Jensen’s inequality we observe that $W_{p}^{\mu}(\eta,\nu)\geq\mathbb{E}_{y\sim\mu}\left[W_{p}(\eta^{y},\nu^{y})\right]$.
-For $p>1$, this inequality is strict unless $W_{p}(\eta^{y},\nu^{y})$ is $\mu$-almost surely constant.
-We emphasize that Definition 5 is purely formal at this stage.
-We do not know if this integral is well-defined, or if it actually defines a metric.
-Before proceeding, we first note that $W_{p}^{\mu}(\eta,\nu)$ may be viewed as the minimal value of the constrained Kantorovich problem in Equation (11) when one takes the cost to be the metric on the space $Y\times U$.
-Although Hosseini et al.
-(2023) studied the existence of minimizers to this problem, they did not study the metric properties of the resulting space.
-Proposition 6 (Equivalent Formulation of the $p$-Wasserstein Distance).
+Appendix C Closed-Form Conditional Wasserstein Distance for Gaussian Measures
+In this section, we provide additional details and results regarding the closed-form conditional Wasserstein distance for Gaussian distributions. See Section 4 in the main paper.
+Suppose $Y$ and $U$ are Euclidean spaces (of possibly different dimensions), and that $\eta,\nu\in\mathbb{P}_{p}^{\mu}(Y\times U)$ are Gaussians of the form
+$\eta=\mathcal{N}\left(\begin{bmatrix}m\\ m_{u}^{\eta}\end{bmatrix},\begin{bmatrix}\Sigma&\Sigma^{\eta}_{12}\\ \Sigma^{\eta}_{21}&\Sigma^{\eta}_{22}\end{bmatrix}\right)\qquad\nu=\mathcal{N}\left(\begin{bmatrix}m\\ m_{u}^{\nu}\end{bmatrix},\begin{bmatrix}\Sigma&\Sigma^{\nu}_{12}\\ \Sigma^{\nu}_{21}&\Sigma^{\nu}_{22}\end{bmatrix}\right).$ (20)
+This form is chosen to ensure that $\eta$ and $\nu$ have equal $Y$-marginals. It follows that $\mu=\pi^{Y}_{\#}\eta=\pi^{Y}_{\#}\nu=\mathcal{N}(m,\Sigma)$. Let
+$Q^{\eta}=\Sigma_{22}^{\eta}-\Sigma_{21}^{\eta}\Sigma^{-1}\Sigma_{12}^{\eta}\qquad Q^{\nu}=\Sigma_{22}^{\nu}-\Sigma_{21}^{\nu}\Sigma^{-1}\Sigma_{12}^{\nu}\qquad R=(\Sigma_{21}^{\eta}-\Sigma_{21}^{\nu})\Sigma^{-1}.$ (21)
+We have that the conditionals $\eta^{y},\nu^{y}$ are available in closed-form:
+$\eta^{y}=\mathcal{N}\left(m_{u}^{\eta}+\Sigma_{21}^{\eta}\Sigma^{-1}(y-m),Q^{\eta}\right)\qquad\nu^{y}=\mathcal{N}\left(m_{u}^{\nu}+\Sigma_{21}^{\nu}\Sigma^{-1}(y-m),Q^{\nu}\right).$ (22)
+Thus, for any fixed $y$, we use the known closed-form unconditional Wasserstein distance to obtain
+$\begin{split}W_{2}^{2}(\eta^{y},\nu^{y})=&\big{|}m_{u}^{\eta}-m_{u}^{\nu}+R(y-m)\big{|}^{2}+\text{Tr}\left(Q^{\eta}+Q^{\nu}-2\left((Q^{\eta})^{1/2}Q^{\nu}(Q^{\eta})^{1/2}\right)^{1/2}\right).\end{split}$ (23)
+We now take an expectation over $y\sim\mu=\mathcal{N}(m,\Sigma)$ to compute $W_{2}^{\mu,2}$.
+Observe that $R(y-m)\sim\mathcal{N}(0,R\Sigma R^{\sf T})$ and that $\mathbb{E}_{y\sim\mu}[|R(y-m)|^{2}]=\text{Tr}(R\Sigma R^{\sf T})$.
+Thus,
+$W_{2}^{\mu,2}(\eta,\nu)$ $=\mathbb{E}_{y\sim\mu}\left[W_{2}^{2}(\eta^{y},\nu^{y})\right]$ (24)
+$=\mathbb{E}_{y\sim\mu}\left[|m_{u}^{\eta}-m_{u}^{\nu}|^{2}+2\langle m_{u}^{\eta}-m_{u}^{\nu},R(y-m)\rangle+|R(y-m)|^{2}\right]$ (25)
+$\qquad+\text{Tr}\left(Q^{\eta}+Q^{\nu}-2\left((Q^{\eta})^{1/2}Q^{\nu}(Q^{\eta})^{1/2}\right)^{1/2}\right)$
+$=|m_{u}^{\eta}-m_{u}^{\nu}|^{2}+\text{Tr}\left(Q^{\eta}+Q^{\nu}-2\left((Q^{\eta})^{1/2}Q^{\nu}(Q^{\eta})^{1/2}\right)^{1/2}+R\Sigma R^{T}\right).$ (26)
+This form, perhaps unsurprisingly, closely resembles the unconditional Wasserstein distance between two Gaussians, except for the presence of an additional $\text{Tr}(R\Sigma R^{\sf T})$ term.
+Note that when $\eta,\nu$ have uncorrelated $Y,U$ components, we precisely recover $W_{2}^{2}(\pi^{U}_{\#}\eta,\pi^{U}_{\#}\nu)$ as one may expect.
+As a special case of interest, if $Y=U=\mathbb{R}$ and
+$\eta=\mathcal{N}(0,I)\qquad\nu=\mathcal{N}\left(0,\begin{bmatrix}1&\rho\\ \rho&1\end{bmatrix}\right)\quad|\rho|<1$ (27)
+then we obtain as a special case of Equation (24) that $W_{2}^{\mu,2}(\eta,\nu)=2(1-\sqrt{1-\rho^{2}})$. This is zero if and only if $\rho=0$, i.e. $\eta=\nu$.
+Appendix D Proofs: Section 4
+In this section, we provide detailed proofs of our claims in Section 4, regarding the metric properties of the conditional Wasserstein space.
+D.1 Metric Properties
+We first note that $W_{p}^{\mu}(\eta,\nu)$ may be viewed as the minimal value of the constrained Kantorovich problem in Equation (3) when one takes the cost to be the metric on the space $Y\times U$.
+Similar results, relating the conditional Wasserstein distance to triangular couplings, have appeared previously, but our proof is independent of these prior works (Chemseddine et al., 2023; Gigli, 2008).
+Proposition 5 (Equivalent Formulation of the Conditional Wasserstein Distance)
 Fix $\eta,\nu\in\mathbb{P}_{p}^{\mu}(Y\times U)$ and $1\leq p<\infty$. Then, $W_{p}^{\mu}(\eta,\nu)$ is well-defined, finite, and
@@
 Thus all considered $p$-Wasserstein distances on $U$ are finite. ∎
-We now show that $W_{p}^{\mu}$ is a metric on $\mathbb{P}_{p}^{\mu}(Y\times U)$.
-Proposition 7 ($W_{p}^{\mu}$ is a Metric).
-For $1\leq p<\infty$, the conditional $p$-Wasserstein distance $W_{p}^{\mu}$ is a metric on $\mathbb{P}_{p}^{\mu}(Y\times U)$.
+We now proceed to prove several metric properties of our distance.
+See 2
 Proof.
+Part (a). This is simply a restatement of Proposition 5.
+Part (b).
 Fix $\eta,\nu,\rho\in\mathbb{P}_{p}^{\mu}(Y\times U)$.
@@
 $=W_{p}^{\mu}(\eta,\rho)+W_{p}^{\mu}(\rho,\nu).$ (37)
-4.1 Metric Properties of Conditional Wasserstein Distances
-We now briefly explore some properties of the space $\mathbb{P}_{p}^{\mu}(Y\times U)$ equipped with the conditional Wasserstein distance $W_{p}^{\mu}$.
-In particular, we observe that the topology on $\mathbb{P}_{p}^{\mu}(Y\times U)$ generated by $W_{p}^{\mu}$ is stronger than that generated by $W_{p}$.
-Note that we easily see $W_{p}^{\mu}(\eta,\nu)\geq W_{p}(\eta,\nu)$ for every $\eta,\nu\in\mathbb{P}_{p}^{\mu}(Y\times U)$ by the characterization in Proposition 6.
-The converse is, however, false, and the metrics are inequivalent.
-Proposition 8 ($W_{p}^{\mu}$ is stronger than $W_{p}$).
-There does not in general exist a constant $C>0$ such that $W_{p}^{\mu}(\eta,\nu)\leq CW_{p}(\eta,\nu)$ for all $\eta,\nu\in\mathbb{P}_{p}^{\mu}(Y\times U)$.
-Proof.
+Part (c).
 We provide a counterexample.
@@
 $W_{p}^{\mu,p}(\eta_{k},\nu_{k})=k^{p}|u_{0}|^{p}\qquad W_{p}^{p}(\eta_{k},\nu_{k})=\min\{k^{p}|u_{0}|^{p},|y_{1}-y_{0}|^{p}\}.$ (39)
-Moreover, as $k\to\infty$ we have $W_{p}^{\mu}(\mu_{k},\nu_{k})\to\infty$ but $W_{p}^{p}(\nu_{k},\eta_{k})$ remains bounded. ∎
-In addition, we can show that the conditional Wasserstein distance dominates the Wasserstein distance on the $U$ marginals.
-Proposition 9 (Conditional Wasserstein Dominates $U$-Marginal Distance).
-For $1\leq p<\infty$ and $\eta,\nu\in\mathbb{P}_{p}^{\mu}(Y\times U)$, we have
-$W_{p}\left(\pi^{U}_{\#}\eta,\pi^{U}_{\#}\nu\right)\leq W_{p}^{\mu}(\eta,\nu).$ (25)
-Proof.
+Moreover, as $k\to\infty$ we have $W_{p}^{\mu}(\mu_{k},\nu_{k})\to\infty$ but $W_{p}^{p}(\nu_{k},\eta_{k})$ remains bounded. See Figure 4.
+Part (d).
+First, the unconditional distance $W_{p}(\eta,\nu)$ may be obtained via an unrestricted coupling in $\Pi(\eta,\nu)$, i.e.
+the set of all joint measures on $Y\times U$ having marginals $\eta,\nu$.
+Since $\Pi(\eta,\nu)\supseteq\Pi_{Y}(\eta,\nu)$, by part (a) we see that $W_{p}(\eta,\nu)\leq W_{p}^{\mu}(\eta,\nu)$.
 Let $\gamma^{\star}(y_{0},u_{0},y_{1},u_{1})=\gamma^{\star,y_{1}}(u_{0},u_{1})\delta(y_{1}-y_{0})\mu(y_{1})$ be an optimal $\gamma^{\star}\in\Pi_{Y}(\eta,\nu)$.
@@
 $\geq W_{p}^{p}(\pi^{U}_{\#}\eta,\pi^{U}_{\#}\nu)$ (46)
-since $\gamma\in\Pi(\pi^{U}_{\#}\eta,\pi^{U}_{\#}\nu)$ is a coupling but potentially sub-optimal. ∎
-Example: Gaussian Measures.
-We conclude with an example where the conditional $2$-Wasserstein distance may be explicitly computed.
-In particular, this example shows that the inequality in Proposition 9 may be strict.
-Suppose $Y$ and $U$ are Euclidean spaces (of possibly different dimensions), and that $\eta,\nu\in\mathbb{P}_{p}^{\mu}(Y\times U)$ are Gaussians of the form
-$\eta=\mathcal{N}\left(\begin{bmatrix}m\\ m_{u}^{\eta}\end{bmatrix},\begin{bmatrix}\Sigma&\Sigma^{\eta}_{12}\\ \Sigma^{\eta}_{21}&\Sigma^{\eta}_{22}\end{bmatrix}\right)\qquad\nu=\mathcal{N}\left(\begin{bmatrix}m\\ m_{u}^{\nu}\end{bmatrix},\begin{bmatrix}\Sigma&\Sigma^{\nu}_{12}\\ \Sigma^{\nu}_{21}&\Sigma^{\nu}_{22}\end{bmatrix}\right).$ (33)
-It follows that $\mu=\pi^{Y}_{\#}\eta=\pi^{Y}_{\#}\nu=\mathcal{N}(m,\Sigma)$. Because we may obtain $\eta^{y},\nu^{y}$ in closed form, we can compute $W_{p}^{\mu}(\eta,\nu)$ in closed form. Let
-$Q^{\eta}=\Sigma_{22}^{\eta}-\Sigma_{21}^{\eta}\Sigma^{-1}\Sigma_{12}^{\eta}\qquad Q^{\nu}=\Sigma_{22}^{\nu}-\Sigma_{21}^{\nu}\Sigma^{-1}\Sigma_{12}^{\nu}\qquad R=(\Sigma_{21}^{\eta}-\Sigma_{21}^{\nu})\Sigma^{-1}$ (34)
-Then, we obtain by a brute-force calculation that
-$W_{2}^{\mu,2}(\eta,\nu)=|m_{u}^{\eta}-m_{u}^{\nu}|^{2}+\text{Tr}\left(Q^{\eta}+Q^{\nu}-2\left((Q^{\eta})^{1/2}Q^{\nu}(Q^{\eta})^{1/2}\right)^{1/2}+R\Sigma R^{T}\right).$ (35)
-Note that when $\eta,\nu$ have uncorrelated $Y,U$ components, we precisely recover $W_{2}^{2}(\pi^{U}_{\#}\eta,\pi^{U}_{\#}\nu)$ as one may expect.
-As a special case of interest, if $Y=U=\mathbb{R}$ and
-$\eta=\mathcal{N}(0,I)\qquad\nu=\mathcal{N}\left(0,\begin{bmatrix}1&\rho\\ \rho&1\end{bmatrix}\right)\quad|\rho|<1$ (36)
-then we obtain as a special case of Equation (35) that $W_{2}^{\mu,2}(\eta,\nu)=2(1-\sqrt{1-\rho^{2}})$.
-This is zero if and only if $\rho=0$, i.e.
-$\eta=\nu$.
-However, $\pi^{U}_{\#}\eta=\pi^{U}_{\#}\nu=\mathcal{N}(0,1)$.
-4.2 Conditional Wasserstein Space as a Geodesic Space
-In this section, we show that $\mathbb{P}_{p}^{\mu}(Y\times U)$ is a geodesic space.
-In fact, we show the stronger statement that there exists a constant speed geodesic between any two measures in $\mathbb{P}_{p}^{\mu}(Y\times U)$, which is a generalization of known results (Santambrogio, 2015, Theorem 5.27).
-First, we introduce some preliminary notions.
-A curve is a continuous function $\gamma_{\bullet}:I\to\mathbb{P}_{p}^{\mu}(Y\times U)$ where $I=(a,b)\subseteq\mathbb{R}$ is any open interval of finite length.
-A curve is said to be absolutely continuous if there exists $m\in L^{1}((a,b))$ such that
-$W_{p}^{\mu}(\gamma_{s},\gamma_{t})\leq\int_{s}^{t}m(\tau)\,\mathrm{d}\tau\qquad\forall a<s\leq t<b.$ (37)
-If $(\gamma_{t})$ is an absolutely continuous curve, then its metric derivative
-$|\gamma^{\prime}|(t)=\lim_{s\to t}\frac{W_{p}^{\mu}(\gamma_{s},\gamma_{t})}{|s-t|}$ (38)
-exists for almost every $t\in(a,b)$, and, moreover, we almost surely have $|\gamma^{\prime}|(t)\leq m(t)$ pointwise for any $m$ satisfying Equation (37) (Ambrosio et al., 2005, Theorem 1.1.2).
-A curve $(\gamma_{t})$ is called a constant speed geodesic if for all $a<s\leq t<b$, we have
-$W_{p}^{\mu}(\gamma_{s},\gamma_{t})=|t-s|W_{p}^{\mu}(\gamma_{a},\gamma_{b}).$ (39)
-It is straightforward to show that every constant speed geodesic is absolutely continuous.
-Theorem 10 ($\mathbb{P}_{p}^{\mu}(Y\times U$) is a Geodesic Space).
-$\mathbb{P}_{p}^{\mu}(Y\times U)$ admits constant-speed geodesics.
-That is, for any $\eta,\nu\in\mathbb{P}_{p}^{\mu}(Y\times U)$, there exists a constant speed geodesic, contained within $\mathbb{P}_{p}^{\mu}(Y\times U)$, between $\eta$ and $\nu$.
+since $\gamma\in\Pi(\pi^{U}_{\#}\eta,\pi^{U}_{\#}\nu)$ is a coupling but potentially sub-optimal.
+$Y$$U$$y_{0}$$y_{1}$$u_{0}$$u_{k}$
+Figure 4: The counterexample in Proposition 2. The measure $\eta_{k}$ is shown in black and the measure $\nu_{k}$ is shown in white.
+D.2 Geodesics
+We now study the geodesics in the space $\mathbb{P}_{p}^{\mu}(Y\times U)$.
+See 1
 Proof.
@@
 ∎
-When an optimal restricted coupling $\gamma^{\star}\in\Pi_{Y}(\eta,\nu)$ is induced by an injective triangular map $T^{\star}$, we may recover a constant speed geodesic in $\mathbb{P}_{p}^{\mu}(Y\times U)$, generalizing the McCann interpolant (McCann, 1997) to the conditional setting.
-We refer to Proposition 4 for sufficient conditions on $\eta,\nu$ under which such a $T^{\star}$ exists.
-Informally, samples from $(y_{0},u_{0})\sim\eta$ flow in a straight path at a constant speed to their destination $T^{\star}(y_{0},u_{0})$.
-We note that this theorem, while primarily about geodesics, relies on notions discussed in Section 5.
-In fact, this may be viewed as a special case of the results in Section 5, where we may explicitly identify the vector field.
-Theorem 11 (Conditional McCann Interpolants).
-Fix $\eta,\nu\in\mathbb{P}_{p}^{\mu}(Y\times U)$.
-Suppose $T^{\star}(y,u)=(y,T^{\star}_{\mathcal{U}}(y,u))$ is a block triangular map solving the conditional Monge problem (8).
-Define the maps $T_{t}:Y\times U\to Y\times U$ for $0\leq t\leq 1$ via
-$T_{t}=(1-t)I+tT^{\star}\qquad(y,u)\mapsto(y,(1-t)u+tT^{\star}_{\mathcal{U}}(y,u))$ (49)
-and define the curve of measures $\gamma_{t}=[T_{t}]_{\#}\eta\in\mathbb{P}_{p}^{\gamma}(Y\times U)$. Then,
-  1. 1.
-    $(\gamma_{t})$ is absolutely continuous and a constant speed geodesic between $\eta,\nu$
-  2. 2.
-    The vector field $v_{t}(T_{t}^{\star}(y,u))=(0,T_{U}^{\star}(y,u)-u)$ generates the path $\gamma_{t}$, in the sense that $(\gamma_{t},v_{t})$ solve the continuity equation (61).
+See 2
 Proof.
@@
 Thus, $\left\lVert v_{t}\right\rVert_{L^{p}(\gamma_{t},Y\times U)}=|\gamma^{\prime}|(t)$ for almost every $t\in[0,1]$ by Lebesgue differentiation.
-5 Conditional Benamou-Brenier Theorem
-In this section, we prove a characterization of all absolutely continuous curves in the space $\mathbb{P}_{p}^{\mu}(Y\times U)$.
-Here we assume $1<p<\infty$.
-First, we introduce the notion of a triangular vector field on $Y\times U$, which informally has no velocity in the $Y$ component.
-Triangular vector fields serve as the dynamic analog of triangular maps (6).
-Definition 12 (Triangular Vector Fields).
-Let $I\subset\mathbb{R}$ be an open interval.
-A time-dependent Borel vector field $v:I\times Y\times U\to Y\times U$ is said to be triangular if there exists a Borel vector field $v^{U}:I\times Y\times U\to U$ such that
-$v_{t}(y,u)=\left(0,v_{t}^{U}(y,u)\right).$ (60)
-In other words, $v_{t}=\iota\left(v_{t}^{U}\right)$ where $\iota:U\to Y\times U$ is the inclusion map.
-5.1 Continuity Equation
-Let $I\subset\mathbb{R}$ be an open interval. The continuity equation
-$\partial_{t}\gamma_{t}+\text{div}(v_{t}\gamma_{t})=0\qquad\text{on }Y\times U\times I$ (61)
-describes the evolution of a measure $\gamma_{t}$ which flows along a given vector field $v_{t}$.
-This equation must be understood distributionally, i.e.
-for every $\varphi$ in an appropriate space of test functions,
-$\int_{I}\int_{Y\times U}\left(\partial_{t}\varphi(y,u,t)+\langle v_{t}(y,u),\nabla_{y,u}\varphi(y,u,t)\rangle\right)\,\mathrm{d}\gamma_{t}(y,u)\,\mathrm{d}t=0.$ (62)
-We consider cylindrical test functions $\varphi\in\text{Cyl}(Y\times U\times I)$, i.e.
-of the form $\varphi(y,u,t)=\psi(\pi^{d}(y,u),t)$ where $\pi^{d}:Y\times U\to\mathbb{R}^{d}$ maps $(y,u)\to(\langle(y,u),e_{1}\rangle,\dots,\langle(y,u),e_{d}\rangle)$ where $\{e_{1},e_{2},\dots,e_{d}\}$ is any orthonormal family in $Y\times U$.
-In the finite dimensional setting, one may take $\varphi\in C_{c}^{\infty}(Y\times U)$ to be smooth and compactly supported (Ambrosio et al., 2005, Remark 8.1.1).
-We now prove a lemma that will be critical to our later results.
-Informally, a solution to the continuity equation with a triangular vector field will result in the conditional measures almost surely satisfying the continuity equaiton as well.
-Lemma 13 (Triangular Vector Fields Preserve Conditionals).
-Suppose $v_{t}(y,u)=(0,v_{t}^{U}(y,u))$ is triangular and that $(\gamma_{t})\subset\mathbb{P}_{p}^{\mu}(Y\times U)$ is a path of measures such that $(v_{t},\gamma_{t})$ satisfy the continuity equation (61) in the distributional sense.
-Then, it follows that for $\mu$-almost every $y\in Y$, we have
-$\partial_{t}\gamma_{t}^{y}+\nabla\cdot(v_{t}^{U}(y,{-})\gamma_{t}^{y})=0.$ (63)
+Appendix E Proofs: Section 5
+In this section, we provide proofs of all claims made in Section 5.
+E.1 Continuity Equation
+We begin with a lemma that is used in the proof of Theorem 4.
+Informally, a solution to the continuity equation with a triangular vector field will result in the conditional measures almost surely satisfying the continuity equation as well.
+See 1
 Proof.
@@
 $\int_{I}\int_{U}\left(\partial_{t}\varphi(u,t)+\langle v_{t}^{U}(y,u),\nabla_{u}\varphi(u,t)\right)\,\mathrm{d}\gamma_{t}^{y}(u)\,\mathrm{d}\mu(y)\,\mathrm{d}t=0\qquad\mu\text{-almost every }y.$ (69)
-We note that having $v_{t}$ be triangular is sufficient, but certainly not necessary, for the conditional continuity equation to almost surely hold.
-For instance, the vector field in $\mathbb{R}^{d}$ that rotates $\mathcal{N}(0,I)$ about the origin preserves all conditional distributions.
-5.2 Absolutely Continuous Curves
-We will show in Theorem 14 that every absolutely continuous curve $(\gamma_{t})\subseteq\mathbb{P}_{p}^{\mu}(Y\times U)$ is generated (in the sense of satisfying the continuity equation) by a triangular vector field such that ${\left\lVert v_{t}\right\rVert_{L^{p}(\gamma_{t},Y\times U)}\leq|\gamma^{\prime}|(t)}$ for almost every $t\in I$.
-This result is obtained as a generalization of the technique in Ambrosio et al.
-(2005, Theorem 8.3.1), which is an application of variational methods for PDEs.
-Conversely, we show in Theorem 15 that if the pair $(\gamma_{t},v_{t})$ solve the continuity equation and $v_{t}$ is triangular, then the curve $(\gamma_{t})$ is absolutely continuous and $|\gamma^{\prime}(t)|\leq\left\lVert v_{t}\right\rVert_{L^{p}(\gamma_{t},Y\times U)}$.
-The main technique of this result is to study the collection of conditional continuity equations (which is feasible by Lemma 13) and to apply the converse of Ambrosio et al.
-(2005, Theorem 8.3.1).
-In this setting, the infinite-dimensional result is obtained via a finite-dimensional approximation argument.
+E.2 Absolutely Continuous Curves
+We now proceed to prove the main results of this section.
+First, we introduce some preliminary notions.
 We define the map $j_{q}:L^{q}(\gamma,Y\times U)\to L^{p}(\gamma,Y\times U)$ for $11/p+1/q=$ 1 via
@@
 which is the Fréchet differential of the convex functional $\frac{1}{q}\left\lVert w\right\rVert_{L^{q}(\gamma,Y\times U)}^{q}$. A straightforward calculation shows that this map satisfies
-$\left\lVert j_{q}(w)\right\rVert^{p}_{L^{p}(\gamma,Y\times U)}=\left\lVert w\right\rVert_{L^{q}(\gamma,Y\times U}^{q}=\int_{Y\times U}\langle j_{q}(w),w\rangle\,\mathrm{d}\gamma(y,u).$ (69)
-Theorem 14 (AC Curves in $\mathbb{P}_{p}^{\mu}(Y\times U)$).
-Let $I\subset\mathbb{R}$ be an open interval, and suppose $\gamma_{t}:I\to\mathbb{P}_{p}^{\mu}(Y\times U)$ is an absolutely continuous in the $W_{p}^{\mu}$ metric with $|\gamma^{\prime}|(t)\in L^{1}(I)$.
-Then, there exists a Borel vector field $v_{t}(x)$ such that
-  1. 1.
-    $v_{t}$ is triangular, i.e. $v_{t}(y,u)=\iota(v_{t}^{U}(y,u))$ for some Borel $v_{t}^{U}:Y\times U\to U$
-  2. 2.
-    $v_{t}\in L^{p}(\gamma_{t},Y\times U)$ for a.e.-$t$
-  3. 3.
-    $\left\lVert v_{t}\right\rVert_{L^{p}(\gamma_{t},Y\times U)}\leq|\gamma^{\prime}|(t)$ for a.e.-$t$
-  4. 4.
-    $(v_{t},\gamma_{t})$ solve the continuity equation distributionally.
+$\left\lVert j_{q}(w)\right\rVert^{p}_{L^{p}(\gamma,Y\times U)}=\left\lVert w\right\rVert_{L^{q}(\gamma,Y\times U)}^{q}=\int_{Y\times U}\langle j_{q}(w),w\rangle\,\mathrm{d}\gamma(y,u).$ (71)
+See also Ambrosio et al. (2005, Chapter 8).
+See 3
 Proof.
@@
 $\left\lVert v_{t}\right\rVert_{L^{p}(\gamma_{t},Y\times U)}\leq|\gamma^{\prime}|(t)\qquad\text{a.e.-}t.$ (87)
-Theorem 15 (Continuous Curves Generated by triangular Vector Fields).
-Suppose that $\gamma_{t}:I\to\mathbb{P}_{p}^{\mu}(Y\times U)$ is narrowly continuous and $(v_{t})$ is a triangular vector field such that $(\gamma_{t},v_{t})$ solve the continuity equation with $\left\lVert v_{t}\right\rVert_{L^{p}(\gamma_{t},Y\times U)}\in L^{1}(I)$.
-Then, $\gamma_{t}:I\to\mathbb{P}_{p}^{\mu}(Y\times U)$ is absolutely continuous in the $W_{p}^{\mu}$ metric and $|\gamma^{\prime}|(t)\leq\left\lVert v_{t}\right\rVert_{L^{p}(\mu_{,}Y\times U)}$ for almost every $t$.
+We now prove, in some sense, a converse of the previous theorem.
+See 4
 Proof.
@@
 $|\gamma^{\prime}|(t)\leq\left\lVert v_{t}\right\rVert_{L^{p}(\gamma_{t},Y\times U)}.$ (100)
-As a corollary of Theorem 14 and Theorem 15, we obtain a conditional version of the Benamou-Brenier theorem (Benamou and Brenier, 2000).
-We note that the proof of Theorem 16 largely follows the unconditional case (see e.g.
-Ambrosio et al.
-(2005, Chapter 8)), but we include it for the sake of completeness.
-Theorem 16 (Conditional Benamou-Brenier).
-Suppose $Y,U$ are separable Hilbert spaces, $1<p<\infty$, and $\mu\in\mathbb{P}(Y)$. Then, for any $\eta,\nu\in\mathbb{P}_{p}^{\mu}(Y\times U)$, we have
-[…]
+Together, Theorem 3 and Theorem 4 give us a dynamical interpretation of the conditional Wasserstein distance.
+The following result is a conditional analogue of the well-known Benamou-Brenier Theorem (Benamou and Brenier, 2000).
+Here, we note that the following proof follows the standard proof closely – the main legwork in obtaining this conditional generalization is through the previous two theorems.
+See 5
 Proof.
@@
 Thus, $W_{p}^{\mu,p}(\eta,\nu)=M$ as desired. ∎
-6 COT Flow Matching
-Figure 1: Results for the Darcy flow inverse problem. We compare the posterior mean and variance of our method (COT-FM) to pCN (Cotter et al., 2013) and WaMGAN (Hosseini et al., 2023).
-We have thus far seen that the COT problem (11) admits a dynamical formulation (99), where one may take the underlying vector fields to be triangular.
-We use these results to design a principled model for conditional generation based on flow matching (Lipman et al., 2022; Albergo et al., 2023b).
-More specifically, recent work (Tong et al., 2023; Pooladian et al., 2023a) has proposed to couple the source and target measures in flow matching via optimal transport.
-We extend these techniques to the conditional setting.
-In this section we make several simplifying assumptions.
-First, we use the squared-distance cost $|(y_{0},u_{0})-(y_{1},u_{1})|^{2}$ throughout (i.e.
-$p=2$).
-Second, we assume that $Y$ and $U$ are finite dimensional Euclidean spaces, but we emphasize that our techniques may be applied to infinite dimensional spaces under additional regularity assumptions (Kerrigan et al., 2023, 2022; Lim et al., 2023; Baldassari et al., 2024).
-Third, we assume that our source and target measures admit densities with respect to the Lebesgue measure on $Y\times U$.
-This is sufficient (Prop.
-4) for the existence of a unique COT plan which is induced by a triangular map.
-We will henceforth, in an abuse of notation, identify measures with their densities.
-Flow Matching.
-We assume that we have access to samples $z_{0}=(y_{0},u_{0})\sim\eta(y_{0},u_{0})\in\mathbb{P}_{p}^{\mu}(Y\times U)$ from a source measure, and samples $z_{1}=(y_{1},u_{1})\sim\nu(y_{1},u_{1})\in\mathbb{P}_{p}^{\mu}(Y\times U)$ from a target measure.
-Let $z=(z_{0},z_{1})\sim\rho(z_{0},z_{1})\in\Pi(\eta,\nu)$ be any coupling of the source and target measure.
-Following Tong et al.
-(2023), we specify a collection of measures and vector fields via
-$\gamma_{t}(y,u\mid z)=\mathcal{N}\left(y,u\mid tz_{1}+(1-t)z_{0},C\right)$ (102)
-$u_{t}(y,u\mid z)=z_{1}-z_{0}$ (103)
-where $C$ is any trace-class covariance operator (Da Prato and Zabczyk, 2014), e.g. $C=\sigma^{2}I$.
-As is standard in flow matching (Lipman et al., 2022; Tong et al., 2023; Kerrigan et al., 2023), we obtain from Equations (102) and (103) a marginal measure $\gamma_{t}(y,u)$ and vector field $u_{t}(y,u)$ satisfying the continuity equation via
-$\gamma_{t}(y,u)=\int_{(Y\times U)^{2}}\gamma_{t}(y,u\mid z)\,\mathrm{d}\rho(z)$ (104)
-$u_{t}(y,u)=\int_{(Y\times U)^{2}}u_{t}(y,u\mid z)\frac{\,\mathrm{d}\gamma_{t}(y,u\mid z)}{\,\mathrm{d}\gamma_{t}(y,u)}\,\mathrm{d}\rho(z).$ (105)
-To model the intractable $u_{t}(y,u)$, we minimize the loss [^1] This is often called the conditional flow matching loss (Tong et al., 2023), which is not to be confused with the notion of conditioning that we focus on in this work.
-$\mathcal{L}(\theta)=\mathbb{E}_{t,\rho(z),\gamma_{t}(y,u\mid z)}\left\lVert u^{\theta}(t,y,u)-u_{t}(y,u\mid z)\right\rVert^{2}$ (106)
-which has the same gradient (with respect to $\theta$ as the MSE loss to the true vector field of interest $u_{t}(y,u)$ (Tong et al., 2023, Theorem 3.2).
-COT Flow Matching.
-Figure 2: Further visualizations of our results on the Darcy flow inverse problem.
-The first row shows the ground-truth posterior.
-Row two shows samples obtained from out COT-FM method, and row three shows the same for WaMGAN (Hosseini et al., 2023).
-The prior and conditioning information are shown in the final two rows.
-In the preceding section, $\rho(z)$ may be an arbitrary coupling between $\eta$ and $\nu$.
-Motivated by Propositions 1 and 4, we will choose $\rho$ to be a conditional optimal coupling, so that $(y_{0},u_{0},y_{1},u_{1})\sim\rho$ has the form $\left(y,u,y,T_{U}(y,u)\right)$.
-By Theorem 16 and Theorem 11, under such a $\rho$ the vector fields in Equation (103) and Equation (105) will be triangular.
-Thus, we parametrize our model $u^{\theta}$ to also be triangular.
-If the covariance operator $C$ in Equation (102) is taken such that $\text{tr}(C)\to 0$, then we recover the optimal path of measures given by the conditional Benamou-Brenier Theorem via a pointwise application of (Tong et al., 2023, Proposition 3.4).
-Given a collection of samples $\{z_{0}^{i},z_{1}^{i}\}_{i=1}^{n}$ drawn from $\eta$ and $\nu$, we approximate a conditional optimal coupling $\rho$ using standard numerical techniques (e.g., those available in the POT package (Flamary et al., 2021)) with the cost function
-$c_{\epsilon}(y_{0},u_{0},y_{1},u_{1})=|y_{1}-y_{0}|^{2}+\epsilon|u_{1}-u_{0}|^{2}$ (107)
-for some $0<\epsilon\ll 1$.
-Intuitively, such a cost penalizes mass transfer along the $Y$ dimension, which is precisely the constraint sought in the COT problem (11).
-Hosseini et al.
-(2023, Prop.
-3.11) show that as $\epsilon\downarrow 0$, we recover the true optimal triangular map.
-When our datasets are not too large, we precompute this COT coupling prior to training our flow matching model.
-Otherwise, we compute this COT coupling for each minibatch drawn at training time.
-After training, we obtain a learned triangular vector field $u^{\theta}_{t}(y,u)$.
-Given an arbitrary fixed $y\in Y$, we may approximately sample from the target $\nu(u\mid y)$ by sampling $u_{0}\sim\eta(u_{0}\mid y)$ and numerically solving the corresponding differential equation.
-For some tasks, like domain translation, note that we are given a fixed $u_{0}$ of interest rather than sampling $u_{0}$ from the source.
-In these cases, we are effectively sampling from $\nu(u\mid y,u_{0})$.
-Source Measure.
-Our framework is agnostic to the choice of source measure $\eta$, allowing for great flexibility in the modeling process.
-The main requirement is that the $Y$-marginal of $\eta$ must match the $Y$-marginal of the target $\nu$.
-In some scenarios, this is trivially satisfied.
-For instance, if one is interested in using a source distribution which is simply random noise, one may take $\eta(y_{0},u_{0})=\pi^{Y}_{\#}\nu(y_{0})\eta_{U}(u_{0})$ to be the product of two independent distributions where $\eta_{U}(u)$ is arbitrary, e.g.
-Gaussian noise.
-Sampling from such an $\eta$ can be achieved by sampling noise from $\eta_{U}(u)$ and an independent $y$ from the target measure.
-We explore more complex choices for $\eta$ in Section 7.
-7 Experiments
-The following section is devoted to illustrate our methodology on a variety of different problems related to conditional simulation.
-First, we explore a likelihood-free inference problem in function space on a Darcy flow inverse problem.
-The dataset used in this case mirrors those in Hosseini et al.
-(2023), as we aim to directly compare with their Conditional Wasserstein Monge GAN (WaMGAN) method.
-In addition, we demonstrate our method on two unpaired image-to-image translation tasks (Nazarpour and Chen, 2017; Choi et al., 2020)
-7.1 Bayesian Inverse Problems
-In this experiment, we consider simulation-based posterior inference on the parameters characterizing a stochastic forward model, whose likelihood evaluation is intractable.
-Darcy Flow
-In this experiment the forward model corresponding to the 2D Darcy flow, an ellipitic PDE on a smooth domain $\Omega\subset\mathbb{R}^{d}$ characterized by a permeability field $\exp(u)$, a pressure field $\rho$, and a source term $f$:
-$-\text{div}\exp(u)\nabla\rho=f\quad\text{in}\quad\Omega,\qquad\rho=0\quad\text{on}\quad\partial\Omega.$ (108)
-For simplicity we denote the deterministic part of the forward model as $\mathfrak{F}:U\rightarrow Y$, and the stochasticity arises from only having access to noisy measurements $y=\mathfrak{F}(u)+\epsilon\in Y,$ where $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$.
-Although the general setting of the experiment corresponds to the one in Hosseini et al.
-(2023), we consider the more general task of performing posterior inference for $u$ when both $u$ and $y$ are observed at arbitrary resolutions, effectively posing the problem in function space.
-Hence, we consider $Y\subseteq\mathcal{Y}$, $U\subseteq\mathcal{U}$, where $\mathcal{Y},\mathcal{U}$ are infinite-dimensional Hilbert spaces.
-We remark that, although the measurements $Y$ are observed on a grid their resolution does not need to be fixed, allowing for training on observations at different resolutions.
-It follows that $\epsilon$ follows a Gaussian measure on $Y$ with a white noise kernel.
-Notice that although the conditional probability measure $\nu^{u}:=\pi^{U}_{\#}\nu$ is Gaussian, evaluating it requires simulating from Equation (108).
-Figure 3: Zero-shot super-resolution on the Darcy flow inverse problem.
-Our model (COT-FM) is trained using a Fourier Neural Operator (Li et al., 2020), allowing us to generate samples at higher resolutions than available at training time.
-We compare to WaMGAN (Hosseini et al., 2023) trained with the same architechture.
-We assign to $u_{1}$ a prior corresponding to the Gaussian measure $\mathcal{N}(0,C_{1})$, where $C_{1}$ corresponds to a symmetric, non-negative and trace-class covariance operator (Da Prato and Zabczyk, 2014).
-In order to obtain a dataset $\mathcal{D}_{1}=\{y_{1,i},u_{1,i}\}_{i=1}^{n}$ of samples on the target domain, we simulate $y_{1}$ from Equation (108) parameterizing the forward model with samples $u_{1}\sim\mathcal{N}(0,C_{1})$.
-We denote the target measure as $\nu$.
-We construct the source measure $\eta$ as the product measure $\mathcal{N}(0,C_{0})\times\mu$, where $C_{0}$ is symmetric, non-negative, trace-class and $\mu=\pi^{Y}_{\#}\eta$ corresponds to the marginal data distribution.
-The choice of picking $C_{0}\neq C_{1}$, while complicating the problem, reflects a scenario where we want to perform inference with a weakly informative prior that does not reflect the true data generating process.
-Notice that sampling from a conditional coupling $\gamma\in\Pi_{Y}(\eta,\nu)$ would still recover samples from the true posterior.
-Here, we assume that $C_{0},C_{1}$ are sufficiently regular such that the path of measures in Equation (104) is well-defined.
+Appendix F Experiment Details
+In this section, we provide additional details regarding all of our experiments, as well as additional results not contained within the main paper.
+All models can be trained on a single GPU with less than 24 GB of memory, and our experiments were parallelized over 8 such GPUs on a local server.
+We first describe our setting for the 2D and Lotka-Volterra experiments, as these share a similar setup.
+Details for the Darcy flow inverse problem are described in the corresponding section.
+Models.
+For FM and COT-FM, our model architecture is an MLP with SeLU activations (Klambauer et al., 2017).
+Time conditioning is achieved by concatenating the time variable as an input to the network.
+The covariance operator $C$ chosen in the path of measures in Equation (8) is taken to be $C=\sigma^{2}I$ where $\sigma$ is a hyperparameter.
+Our implementation of FM is adapted from the torchcfm package Tong et al.
+(2023), available under the MIT License.
+For PCP-Map and COT-Flow, we adapt the open-source implementations from Wang et al.
+(2023), available under the MIT License.
+Training and Model Selection.
+Hyperparameter tuning of the PCP-Map and COT-Flow models was performed directly using the code of Wang et al.
+(2023), essentially implementing grid-search with an early stopping procedure.
+We refer to the paper and codebase of Wang et al.
+(2023) for further details.
+For COT-FM and FM, we perform a random grid search over 100 hyperparameter settings using the grid described in Table 4.
+For all model types, we select the best model used to generate the results in the paper as the training checkpoint that resulted in the lowest $W_{2}$ error to the joint target distribution on a held-out validation set.
+For training, we use the Adam optimizer where we only tune the learning rate, leaving all other settings as their defaults in pytorch.
+Table 4: Hyperparameter grid used for random search of the FM and COT-FM models on the 2D and Lotka-Volterra datasets.
+Hyperparameter Description Values
+$\epsilon$ COT coupling strength [1e-6, 1e-4, 1e-2, 1e-1]
+$\sigma$ Variance for $C=\sigma^{2}I$ in (8) [1e-3, 1e-2, 1e-1, 5e-1]
+Batch Size Training batch size [256, 512, 1024]
+Width Layer width in MLP [256, 512, 1024, 2048]
+LR Learning rate [1e-4, 3e-4, 7e-4, 1e-3]
+Layers Number of MLP layers [4, 6, 8]
+F.1 2D Synthetic Data
+Data Generation.
+This experiment consists of four 2D synthetic datasets, where $Y=U=\mathbb{R}$.
+The datasets moons, circles, swissroll are available through scikit-learn (Pedregosa et al., 2011).
+The moons dataset is generated with noise=$0.05$ followed by standard scaling with a mean of $m=(0.5,0.25)$ and standard deviation of $\sigma=(0.75,0.25)$.
+The circles dataset is generated with factor=0.5 and noise=0.05.
+The swissroll dataset is generated with noise=0.75, followed by projection to the first two coordinates and re-scaling by a factor of $12$.
+All other unstated parameters are left as their default values.
+We use the code available from Hosseini et al.
+(2023) to generate the checkerboard dataset.
+For all datasets, we generate a training set (i.e., samples from the target distribution) of $20,000$ samples and $1,000$ held-out validation samples for model selection.
+Means and standard deviations in Table 1 are reported across five independent testing sets of 5,000 samples for the best representative of each model type.
+In COT-FM, to generate samples from the source distribution, we sample an additional $20,000$ points from the target distribution and keep only the $Y$ coordinates.
+This ensures that the source and target have equal $Y$ marginals.
+During training, standard Gaussian noise $\mathcal{N}(0,1)$ is sampled for the $U$ coordinate of these source points at each minibatch.
+We use minibatch COT couplings (Tong et al., 2023) in this experiment as computing the full COT plan was prohibitively expensive in terms of memory usage.
+However, we note that we use large batch sizes, meaning that the COT plan we find in this way should not be too far from optimal.
+All couplings are computed using the POT Python package (Flamary et al., 2021).
+Figure 5: Samples from the ground-truth joint target distribution and the various models for the 2D datasets.
+Samples from COT-FM more closely match the ground-truth distribution than the baselines.
+A common failure mode for the baselines is to generate samples from regions with zero support under the true data distributions.
+Table 1 contains a quantitative evaluation.
+Figure 6: Conditional KDEs shown for each of the methods on the 2D datasets.
+The conditioning variable $y$ is fixed at the horizontal dashed line shown in Figure 5.
+In all plots, the orange solid line indicates the CKDE of the ground-truth joint samples.
+In each column, the dashed blue line indicates the CKDE of samples generated from the respective method.
+F.2 Lotka-Volterra Dynamical System
+Data Generation.
+We adopt the settings of Alfonso et al.
+(2023) for this experiment.
+As described in the main paper, we assume $p(0)=(30,1)$ and that $\log(u)\sim\mathcal{N}(m,0.5I)$ with $m=(-0.125,-3,-0.125,-3)$.
+Given parameters $u\in\mathbb{R}^{4}_{\geq 0}$, we simulate Equation (11) for $t\in\{0,2,\dots,20\}$ to obtain a solution $z(u)\in\mathbb{R}^{22}_{\geq 0}$.
+An observation $y\in\mathbb{R}^{22}_{\geq 0}$ is obtained by the addition of log-normal noise, i.e.
+$\log(y)\sim\mathcal{N}(\log(z(u),0.1I)$.
+We thus may simulate many $(y,u)$ pairs from the target measure for training.
+We generate a training set of $10,000$ $(y,u)$ pairs using the procedure described above and a held-out validation set of $10,000$ $(y,u)$ pairs for model selection.
+Means and standard deviations in Table 2 are reported across five independent testing sets of 5,000 samples for the best representative of each model type.
+Figure 2 and Figures 9, 8, 7, 10 show $10,000$ samples from each model, as well as $10,000$ samples from the differential evolution Metropolis MCMC sampler (Braak, 2006) after a burn-in of $50,000$ samples.
+This is implemented through the PyMC Python package (Abril-Pla et al., 2023).
+For COT-FM we use the full COT couplings, i.e.
+without minibatches.
+This is available to use due to the smaller size of the training set used in this experiment.
+The COT couplings are computed in the same way as the previous section, and as described in Section 7.
+Figure 7: KDE plots of the samples on the Lotka-Volterra system, using the settings described in Section 7.
+Plots include one-dimensional KDEs on the diagonal, as well as all two-dimensional pairs.
+In all plots, samples from MCMC are drawn in orange, and samples from our method (COT-FM) are indicated in blue.
+The true unknown parameters are indicated by the red vertical line in the diagonal plots, or the black x in the off-diagonal plots.
+Figure 8: KDE plots of the samples on the Lotka-Volterra system, using the settings described in Section 7.
+Plots include one-dimensional KDEs on the diagonal, as well as all two-dimensional pairs.
+In all plots, samples from MCMC are drawn in orange, and samples from PCP-Map are indicated in blue.
+The true unknown parameters are indicated by the red vertical line in the diagonal plots, or the black x in the off-diagonal plots.
+Figure 9: KDE plots of the samples on the Lotka-Volterra system, using the settings described in Section 7.
+Plots include one-dimensional KDEs on the diagonal, as well as all two-dimensional pairs.
+In all plots, samples from MCMC are drawn in orange, and samples from COT-Flow are indicated in blue.
+The true unknown parameters are indicated by the red vertical line in the diagonal plots, or the black x in the off-diagonal plots.
+Figure 10: KDE plots of the samples on the Lotka-Volterra system, using the settings described in Section 7.
+Plots include one-dimensional KDEs on the diagonal, as well as all two-dimensional pairs.
+In all plots, samples from MCMC are drawn in orange, and samples from flow matching (FM) are indicated in blue.
+The true unknown parameters are indicated by the red vertical line in the diagonal plots, or the black x in the off-diagonal plots.
+F.3 Inverse Darcy Flow
+Dataset.
+The training and test datasets are generated following the same procedure as Hosseini et al.
+(2023): pressure fields $u$ are sampled from a Gaussian process with Matérn kernel having $\nu=3/2$ and lengthscale $\ell=1/2$, on a regular $40\times 40$ grid.
+The parameters are then exponentiated and used to simulate the permeability fields $p$ from the forward model $\mathfrak{F}$ solving the Darcy flow PDE, using FEniCS (Alnæs et al., 2015).
+Stochasticity arises from adding Gaussian noise to the permeability fields, obtaining $y=\mathfrak{F}(u)+\epsilon,\,\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$.
+For our experiments we observe $y$ on a $100\times 100$ grid, and we use $\sigma=2.5\times 10^{-2}$.
+We note that this level of noise is quite considerable, as it accounts for roughly $60\%$ of the variability in the $y$.
+Figure 11 showcases a data point for reference.
+Our source and target training sets contain $1\times 10^{4}$ samples each, and our test set comprises $5\times 10^{3}$ samples.
+We remark that although $y$ and $u$ are observed on a grid their resolution does not need to be fixed, allowing for training at different resolutions.
+Figure 11: Example of one random data point from the Darcy flow dataset.
+Figure 12: Additional results on the Darcy flow inverse problem.
+A true permeability $u$ is shown, as well as the pressure field $\rho$ and its observed, noisy version $y$.
+We compare an ensemble average of posterior samples from the various methods against MCMC (pCN) (Cotter et al., 2013).
+COT-FM achieves the lowest MSE to pCN.
+We note here that WaMGAN has clear visual artifacts despite achieving reasonable MSE and CRPS scores.
+Models.
 […]
-Figure 4: Samples from our model on a supervised image-to-image translation task, where translation is performed from Chinese MNIST to MNIST (a) and from MNIST to Chinese MNIST (b).
-This demonstrates that our method is able to use non-trivial source distributions and is able to successfully draw samples from the corresponding conditional measure.
+  • WaMGAN (Hosseini et al., 2023): using an adaptation to the FNO architecture of the original code [^3] https://github.com/TADSGroup/ConditionalOT2023, we perform a grid search as detailed in Table 5.
+We found the training procedure to be rather unstable, and for this reason we checkpoint the model every 100 epochs and report the results for the best performing model at its best checkpoint.
+We found this to be a model with learning rate $1\times 10^{-4}$, 2 full critic iterations, and monotone penalty of $1\times 10^{-3}$.
+The gradient penalty parameter did not seem to significantly affect performance on the test set, and was set to 5.
+  • FFM (Kerrigan et al., 2024): the learning rate is fixed to $5\times 10^{-4}$, and the covariance operator $C$ is set to match that of the prior, but rescaled by a factor of $\sigma=1\times 10^{-3}$.
+We use the code from the original repository [^4] https://github.com/GavinKerrigan/functional_flow_matching.
+  • COT-FFM: we set $\epsilon=1\times 10^{-5}$ in the cost function used to build the COT plan.
+The learning rate and $C$ are chosen to be the same as FFM.
+In order to build COT couplings, we take the source measure to be the product measure $\pi^{Y}_{\#}\eta\times\mathcal{N}(0,C)$.
+Approximate couplings are obtained on minibatches of size $256$.
+It should be noted that in any scenario where the source and the target $U-$marginals are identical, using the OT coupling would yield the identity mapping as the optimal vector field minimizing (10).
+Hence, the OT-CFM model (Tong et al., 2023) is inapplicable here.
+Sampling.
 The resulting amortized sampler, denoted for simplification by the mapping $(y,u_{0})\mapsto u_{1}=\tilde{T}_{U}(y,u_{0})$, will parameterize an approximate posterior measure.
-Notice that, in contrast to variational inference techniques, no distributional assumptions are made about the approximate posterior.
-In turn, integrals are obtained numerically by Monte Carlo sampling $K-$many samples from the prior, resulting in the approximation
+Notice that, in contrast to classical variational inference techniques, no distributional assumptions are made about the approximate posterior.
+In turn, integrals are obtained numerically by Monte Carlo sampling $K$ samples from the prior, resulting in the approximation
 […]
-In Figure 1, we show an example of how our method compares to pCN and WaMGAN on recovering the posterior mean and variance on a given example.
-For a fair comparison and to ensure resolution-invariance, we used the same architecture for both neural models, with the discriminator having an additional layer projecting to a scalar value.
-However, the WaMGAN method used twice as many parameters as it requires two neural networks.
-Each statistic is computed on $2\times 10^{3}$ samples from each method.
-We showcase the ability of our method to perform zero-shot super-resolution in Figure 3.
-A random selection of samples from our method and from WaMGAN, as well as the prior sample and the different conditions used for generation, are displayed in Figure 2.
-Qualitatively, our samples appear to better match the true posterior compared to WaMGAN (Hosseini et al., 2023), while being cheaper to sample than pCN (Cotter et al., 2013).
-7.2 Unpaired Image-to-Image Translation
-Figure 5: Uncurated sample from the test set of the Animal Faces HQ dataset (Choi et al., 2020) from the dogs class, translated to cats.
-These results demonstrate our method on an unsupervised image-to-image task, where we utilize a pre-trained content extractor.
-We now demonstrate our method on an image-to-image translation task.
-In particular, in this section our source measures are distributions over images.
-The objective of domain translation is to convert information from one domain to another by adjusting domain-specific attributes, denoted style, while preserving the domain-invariant ones, denoted content (Isola et al., 2017).
-Semantic attributes can be associated with human-assigned labels (supervised) or inferred by a learning algorithm that learns domain-invariant representations of the data (unsupervised).
-Pseudocode is available in Appendix A.
-Historically, the problem of domain translation was tackled through methods that heavily relied on parallel datasets—pairs of corresponding samples from the source and target domains.
-However, obtaining such paired datasets is often impractical or impossible in many real-world applications, prompting the shift towards methods that do not require paired data (unpaired domain translation) (Pang et al., 2021).
-Domain translation is possible for a learning algorithm if it satisfies cycle consistency, positing that translation of a source sample $u_{s}$ into a target sample $\hat{u}_{t}$ and then back to the source domain should recover $\hat{u}_{s}=u_{s}$, and the same shall hold by starting from a target sample.
-Notice that this property is achieved naturally by our model, and hence does not require additional terms in the loss as is the case for GAN- or VAE-based methods (Zhu et al., 2017; Isola et al., 2017; Liu et al., 2017).
-Here, the space $U$ corresponds to the space of images.
-For the space $Y$, we explore two choices.
-First, if supervised information (i.e., a class label) is available, we take $Y$ to be the space of labels.
-In this setting, if the class distribution in both the source and target distribution is balanced, then the $Y$-marginals of the source and target measures match.
-When such information is not available, we extract semantic information from our images using a pre-trained encoder, mapping images to a latent space $Y$.
-In this case, we assume the encoding distribution matches across the source and target distributions.
-Exploring techniques for balancing this distribution is an interesting avenue for future work.
-Supervised.
-We showcase in Figure 4 the ability of our method to address the task of class-conditional image translation, using the MNIST and the Chinese MNIST (Nazarpour and Chen, 2017) datasets.
-The second dataset was preprocessed in order to center the images around their content, and both are reshaped to a resolution of $32\times 32$.
-The size of the training set was small enough for us to compute the empirical COT plan on the entire dataset, and pairs were resampled at each epoch.
-We parameterize the neural network in COT-FM with a UNet (Ronneberger et al., 2015), and we compute the $Y-$distance in the COT plan by taking the $L_{2}$ distance between one-hot encoded vectors for the labels.
-The UNet has hidden dimensionality of 128, and we use three downsampling/upsampling blocks with multipliers $(1,2,2)$.
-The optimizer is Adam (Kingma and Ba, 2014) with a learning rate of $10^{-4}$, at a batch size of 128 for $3\times 10^{5}$ training steps.
-Unsupervised.
-Samples are mapped into a shared representation space by a pre-trained feature extractor $\mathfrak{F}:\mathbb{R}^{d}\rightarrow[0,1]^{e}$, obtaining content variables $y=\mathfrak{F}(u)$.
-These effectively act as pseudo-labels when building the empirical COT matrix.
-In our example, we perform translation from the dogs class to the cats class in the AnimalFacesHQ dataset (Choi et al., 2020) and we use CLIP as our feature extractor (Radford et al., 2021) as found in the official Hugging Face repository in its version clip-vit-large-patch14.
-The training set is comprised of 5067 images of cats and 4680 images of dogs, enabling the computation of the COT plan in advance.
-As with the previous experiment, samples are repaired by sampling from the COT plan at each epoch.
-Images are observed at a native resolution of $512\times 512$, but to ease the memory requirements we employ another pre-trained encoder $\mathfrak{E}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{z}$ for dimensionality reduction, effectively performing flow matching in latent space (Dao et al., 2023).
-The procedure is detailed in Algorithm 1.
-We take the encoder $\mathfrak{E}$ to be the variational autoencoder (Kingma and Welling, 2013) from Stable Diffusion v2.1, enabling training of the COT-FM network at a dimensionality of $4\times 64\times 64$.
-We use the Adam optimizer (Kingma and Ba, 2014) with a learning rate of $2\times 10^{-4}$, with a batch size of 32 for $4\times 10^{5}$ training steps.
-We show several uncurated conditional samples from our model in Figure 5.
-8 Conclusion
-We analyze conditional optimal transport from a geometric and dynamical point of view.
-Our analysis culminates in the characterization of absolutely continuous curves of measures in a conditional Wasserstein space, resulting in a conditional analog of the Benamou-Brenier Theorem.
-We use these result to build on the framework of triangular transport and flow matching to develop simulation-free methods for conditional generative models.
-Our methods are applicable across a wide class of problems, and we demonstrate our methodology on an infinite-dimensional Bayesian inverse problem as well as image translation tasks.
-Acknowledgments
-This research was supported by the Hasso Plattner Institute (HPI) Research Center in Machine Learning and Data Science at the University of California, Irvine, by the National Science Foundation under award 1900644, and by the National Institutes of Health under award R01-LM013344.
-Appendix A Pseudocode
-1:  Input: Data $\{u_{0,i},u_{1,i}\}_{i=1}^{B}\sim\gamma^{\star}_{COT}$, pretrained content extractor $\mathfrak{F}$, interpolant network $v_{\theta}$, pretrained encoder $\mathfrak{E}$ (optional).
-2:  Initialize: $\mathcal{L}=0$
-3:  $\{y_{0,i},y_{1,i}\}_{i=1}^{B}\leftarrow\{\mathfrak{F}(u_{0,i}),\mathfrak{F}(u_{1,i})\}_{i=1}^{B}$
-4:  if  $\mathfrak{E}$ is given  then
-5:     $\{u_{0,i},u_{1,i}\}_{i=1}^{B}\leftarrow\{\mathfrak{E}(u_{0,i}),\mathfrak{E}(u_{1,i})\}_{i=1}^{B}$
-6:  end if
-7:  for $i=1,\dots,B$ do
-8:     Sample $t_{i}\sim\mathcal{U}([0,1])$
-9:     Interpolate: $u_{t,i}=(1-t_{i})u_{0,i}+t_{i}u_{1,i}$
-10:     Evaluate interpolant network:
-$\hat{v}_{t_{i}}^{0}=v_{\theta}(u_{t,i},t_{i},y_{0,i}),\quad\hat{v}_{t_{i}}^{1}=v_{\theta}(u_{t,i},t_{i},y_{1,i})$
-[…]
-12:  end for
-13:  $\theta_{\text{new}}\leftarrow$ OptimizerStep$(\nabla_{\theta}\mathcal{L})$
-Algorithm 1 DCOT Flow for Unsupervised Image-to-Image Translation, Training Step
+Table 5: Hyperparameter search space for WaMGAN
+Parameter Search Space
+Learning rate $1\times 10^{-3}$, $5\times 10^{-4}$, $1\times 10^{-4}$}
+Full critic iter. {2, 5, 10}
+Monotone penalty $1\times 10^{-3}$, $5\times 10^{-2}$, $1\times 10^{-1}$}
+Gradient penalty {1, 5, 10}